### 12 Mar 2026 v1

# Forecasting and Manipulating the Forecasts of Others

Abstract
In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades. We provide the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals. Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic two-time kernels, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit. The characterization yields an explicit information wedge $\mathcal{V}^{i}_{t}$—a deterministic Volterra process—that prices the marginal value of shifting opponents’ posteriors. The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
1 Introduction
In strategic environments with private information, evaluating a change in policy—a disclosure mandate, a shift in the precision of public communication, a redesign of market transparency—requires understanding how the equilibrium responds, not merely how agents behave under existing rules. When a regulator alters the information structure, each agent’s optimal action depends on their beliefs about the state and on their beliefs about how others are filtering, and both change with the policy. The filtering environment and the strategic equilibrium are jointly determined.
This joint determination is the central obstacle. Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates. Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements. Under the new policy, signal precision changes, which changes how firms filter, which changes how each firm forecasts the pricing decisions of others, which changes equilibrium inflation dynamics. Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable. The characterization in this paper makes it explicit.
The same obstruction arises across a wide range of fields: in market microstructure, where prices aggregate dispersed information and trading itself is informative; in dynamic information design, where the mapping from signal structures to equilibrium outcomes is intractable when filtering and equilibrium interact; in contract theory, where multiple agents exert unobservable effort on a shared outcome and draw inferences from it; and in decentralized control—autonomous vehicle fleets, multi-robot systems with misaligned operators—wherever agents observe a shared state through private channels and actions reshape what others sense. Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
Townsend [28] crystallized the difficulty: in a game with private signals, each agent must forecast others’ forecasts of the state, generating an infinite hierarchy of conditional expectations. Sargent [26] showed that even in linear–Gaussian environments the hierarchy prevents finite-dimensional recursive analysis, because each agent’s action enters the state dynamics, entangling every agent’s past actions, beliefs, and signals in a way the Kalman filter cannot close.
The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games. Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning. We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by orthogonal projection onto deterministic two-time kernels with no truncation. The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the deterministic kernels governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
In the single-agent problem the conditional mean of the state is a sufficient statistic, but the state estimate is a linear functional of the noise estimate, not the reverse. In the multi-agent setting the state estimate fails to close the equilibrium because it entangles exogenous shocks with endogenous responses—precisely the Townsend hierarchy. Estimating the noise path avoids this entanglement. We call the path $\widehat{W}_{t}^{i}(\cdot):=\mathbb{E}[W_{\cdot}\mid\mathcal{F}_{t}^{i}]$ the player’s noise-state; optimal controls are affine functionals of it, just as single-agent controls are affine in the state estimate.
The main results establish that the best response to any profile of deterministic-kernel strategies is itself deterministic-kernel (closure over the full admissible $L^{2}$ class), so Nash equilibrium reduces to a deterministic fixed point in two-time kernels. The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$: a Volterra process with deterministic mean $\bar{\mathcal{V}}^{i}(t)$ and deterministic kernel $\mathcal{V}^{i}(t,r)$, which prices the marginal value to player $i$ of shifting opponents’ posteriors. The mean wedge distorts equilibrium mean actions; the kernel wedge distorts equilibrium impulse responses. Both vanish identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters. The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths $\{P^{i}(\cdot)\}$ and affects the equilibrium only through deterministic filtering sensitivities, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
(i) Candidate Opponents use Volterra controls (ii) Closure State & filters have det. kernels (iii) Deviations Fixed maps $\Rightarrow$ det. propagation (iv) Best response FOC $\Rightarrow$ control linear in $\widehat{W}^{\,i}$ (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
Figure 1: The Volterra fixed-point loop. (i) Assume opponents use Volterra controls with deterministic kernels. (ii) The state and filtering equations close on deterministic two-time kernels. (iii) Under fixed strategy maps, unobserved deviations propagate deterministically. (iv) The maximum principle produces a best response that is itself Volterra, closing the loop. (v) A fixed point of this best-response map is a Nash equilibrium.
1.1 Related Literature
The joint determination of filtering and equilibrium has been approached from several directions, each achieving tractability by restricting or removing the informational externality that makes the problem hard.
Higher-order expectations and the social value of information.
The infinite regress of “forecasting the forecasts of others” was identified by Townsend [28] and shown by Sargent [26] to resist finite-dimensional recursive analysis even in linear–Gaussian environments. Morris and Shin [20] demonstrated that the social value of public information can be negative when agents overweight common signals—a static manifestation of the belief-manipulation channel that arises dynamically through the information wedge $\mathcal{V}^{i}_{t}$. Our Proposition 3.6 provides a dynamic counterpart: Section 4 shows that nearly all welfare gains from pooling come from eliminating the strategic channel rather than improving state estimation, consistent with the Morris–Shin intuition that the coordination externality, not statistical efficiency, drives welfare losses.
Frequency-domain methods and exogenous-signal models.
A substantial literature solves dispersed-information models in the frequency domain [9, 14, 24], culminating in Huo and Takayama [12], who obtain finite-state representations for equilibrium aggregates under exogenous ARMA signals. These methods succeed because the filtering environment is invariant to changes in control laws—the spectral factorization can be solved independently of the policy. When controls enter the state dynamics the transfer function depends on the policy, breaking this separation; the present paper addresses this complementary case.
Common-information approaches.
A powerful technique in decentralized control conditions on a common information base shared by all agents to obtain a recursive reduction [21]. The approach handles asymmetric information but resolves it through a coordinator’s beliefs; the channel by which player $i$’s action shifts player $k$’s posterior, which shifts $k$’s action, which shifts $i$’s signal, is not captured.
Team theory.
The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for LQG teams. Team-theoretic solutions bypass the belief hierarchy because aligned objectives eliminate the incentive to manipulate others’ forecasts. The information wedge $\mathcal{V}^{i}_{t}$ is precisely the object that vanishes under team objectives and survives under misaligned payoffs.
Mean-field LQG games.
Mean-field games take large-population limits in which no single agent’s action measurably affects any other agent’s signal [6, 11]. Carmona and Delarue [7] provide a comprehensive two-volume treatment including LQG specifications with common noise. The bilateral informational externality is absent by construction in the mean-field limit; $\mathcal{V}^{i}_{t}$ is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
Truncation of the belief hierarchy.
Nimark [22] approximates the higher-order belief hierarchy by truncation at finite order—the closest existing method in spirit to the present paper. The deterministic kernel system derived here provides an exact benchmark against which truncation error can be assessed.
Applied continuous-time filtering games.
Kyle [15], Back [2], and Foster and Viswanathan [8] solve strategic trading problems with private information, each exploiting special structure—a single informed trader, correlated-signal symmetry—to close the belief hierarchy through model-specific conjectures. Sannikov [25] solves a continuous-time principal–agent problem by showing that the agent’s continuation value serves as a sufficient statistic for the principal’s problem, yielding a recursive characterization despite the agent’s private action. The technique relies on the bilateral structure: the principal’s belief about a single agent’s hidden action can be summarized by a one-dimensional state variable, and the agent has no beliefs about other agents to forecast. With multiple agents exerting unobservable effort on a shared outcome—the moral-hazard analogue of our setting—each agent draws inferences from the shared output about others’ effort, generating the same belief hierarchy that arises here. The noise-state framework extends naturally to this setting: each agent’s hidden effort enters the output dynamics exactly as controls enter the state in (2.1), and the information wedge prices the value of shifting the principal’s and co-agents’ posteriors about one’s effort. The present paper provides a general framework that nests these as special cases: in the Kyle–Back embedding (Section 5), the information wedge corresponds to the strategic component of price impact, and the characterization extends to multi-trader settings where guess-and-verify breaks down.
Information design.
Kamenica and Gentzkow [13] established Bayesian persuasion; Bergemann and Morris [3, 4] introduced Bayes correlated equilibrium and survey the field, identifying dynamic decentralized settings as a central open problem. Bergemann, Heumann, and Morris [5] provide precision comparative statics in static coordination games analogous to the propagation $\Delta P^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta J^{i}$ derived here. The difficulty in extending these results dynamically is that filtering and equilibrium interact; our characterization provides the equilibrium mapping that dynamic information design requires as an input.
Rational inattention.
Sims [27] introduced rational inattention; Maćkowiak and Wiederholt [17] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [18] for a survey). These models take the strategic environment as given when solving the attention problem; Corollary 3.7 and the FOC (C.9) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.
2 The Decentralized LQG Game
Informal description of the game.
The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon. Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational. Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable.
2.1 Primitives
Consider $n$ players interacting over a finite time horizon $[0,T]$ on a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$.
State dynamics.
The state $X_{t}\in\mathbb{R}^{d}$ evolves as
$dX_{t}=\bigl(A(t)X_{t}+\sum_{i=1}^{n}D_{t}^{i}\bigr)dt+\Sigma(t)\,dW_{t}^{0},\qquad X_{0}=x_{0},$ (2.1)
where $A(t)\in\mathbb{R}^{d\times d}$ is deterministic and bounded, $W^{0}$ is a $d$-dimensional Brownian motion, $\Sigma(t)\in\mathbb{R}^{d\times d}$ is deterministic and bounded, and $D_{t}^{i}\in\mathbb{R}^{d}$ is player $i$’s control.
Remark 2.1 (Control matrices and dimensions).
All results extend to $D_{t}^{i}=B_{i}(t)U_{t}^{i}$ with control $U^{i}_{t}$ by appropriately modifying the cost function and first order conditions.
Observations.
Player $i$ does not observe $X_{t}$ directly. Instead, they observe a noisy measurement process $Y^{i}_{t}\in\mathbb{R}^{d}$ given by
$dY^{i}_{t}=\sqrt{P^{i}(t)}X_{t}\,dt+dW^{i}_{t},\quad Y^{i}_{0}=0,$ (2.2)
where: [^1] The square-root parameterization is convenient for interpreting $P^{i}(t)$ as a precision matrix, but all results extend to observation channels of the form $dY^{i}_{t}=H^{i}(t)X_{t}\,dt+E^{i}(t)\,dW_{t}$ with arbitrary deterministic $H^{i}(t)\in\mathbb{R}^{d_{i}\times d}$ and $E^{i}(t)E^{i}(t)^{\top}=I$, including correlated observation noises across players and signals that weight different state coordinates asymmetrically. All kernel formulas remain verbatim under this generalization.
  • $W^{i}_{t}\in\mathbb{R}^{d}$ is a standard Brownian motion (player $i$’s observation noise),
  • $P^{i}(t)\in\mathbb{R}^{d\times d}$ is symmetric positive semidefinite, and $\sqrt{P^{i}(t)}$ denotes its unique symmetric positive semidefinite square root.
  • All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
Stacked noise and block selectors.
We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$. For each $j\in\{0,\dots,n\}$, the block selector $E^{j}\in\mathbb{R}^{d\times(n+1)d}$ extracts the $j$-th channel: $E^{j}W_{t}=W^{j}_{t}$. The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$. Player $i$ directly observes $\Pi^{i}W$ and must infer $(I-\Pi^{i})W$ from drift-based learning.
Player $i$’s information at time $t$ is their observation history, $\mathcal{F}^{i}_{t}:=\sigma(Y^{i}_{s}:s\leq t)$. An admissible control is $\mathcal{F}^{i}_{t}$-adapted with $\mathbb{E}[\int_{0}^{T}\|D^{i}_{t}\|^{2}\,dt]<\infty$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
Unresolved state kernel.
For each player $i$, define
$\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u),$
the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate. Formally, $\widetilde{X}^{i}(t,u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
Objectives.
Each player $i$ incurs the random cost
$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{i}_{XX}(t)X_{t}+2(G^{i}_{X}(t))^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{i}_{DD}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{i}_{XX}(T)X_{T}+2G^{i}_{X}(T)X_{T},\end{split}$ (2.3)
where $G^{i}_{XX}(t)$ and $G^{i}_{DD}(t)$ are symmetric positive semidefinite matrices (with $G^{i}_{DD}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls. Each player $i$ seeks to minimize
$J^{i}(D^{i};D^{-i}):=\mathbb{E}\bigl[\mathcal{C}^{i}(D^{i};D^{-i})\bigr].$ (2.4)
Opponents’ controls enter player $i$’s cost only through $X_{t}$, which they move via (2.1) and which player $i$ observes through (2.2).
2.2 Nash Equilibrium
Because each player’s drift depends on opponents’ estimates $\widehat{X}_{t}^{k}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{k}]$, optimal actions generate an infinite hierarchy of conditional expectations. The following equilibrium concept fixes opponents’ strategy maps rather than their realized actions, which is essential for the closure results in Section 3.
Definition 2.1 (Nash equilibrium in private-signal strategies).
Throughout, a player’s choice is a strategy map: for each $t\in[0,T]$, the action $D_{t}^{i}$ is a (progressively measurable) functional of the private observation history $Y^{i}_{\cdot\wedge t}$. We write $D_{t}^{i}=D_{t}^{i}[Y^{i}]$, where the bracket notation $D_{t}^{i}[\cdot]$ denotes the map from the observation path $(Y^{i}_{s})_{s\leq t}$ to the action at time $t$. We identify a strategy with its induced adapted control process, assuming $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
A profile $D^{*}=(D^{*,1},\ldots,D^{*,n})$ is a Nash equilibrium if for every player $i$ and every admissible alternative strategy map $\widetilde{D}^{i}[\cdot]$,
$J^{i}(D^{*,i};D^{*,-i})\leq J^{i}(\widetilde{D}^{i};D^{*,-i}),$
where the inequality compares the costs induced by the corresponding strategy profiles. Equivalently: in a unilateral deviation by $i$, opponents’ maps are held fixed and are applied to the deviated signal paths.
In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the Volterra class these maps are parameterized by deterministic kernels that remain unchanged off equilibrium.
The requirement that opponents’ strategy maps be held fixed under deviations—rather than their realized actions or their realized beliefs—is the formal expression of the Lucas critique in this environment: if player $i$ changes their control law, the state dynamics change, which changes every opponent’s signal process, filtering problem, and actions, which feeds back into the state.
This paper aims to establish that the Volterra class is closed under best responses and to derive a deterministic kernel system whose fixed points correspond to Nash equilibria in that class.
Noise-state.
For each player $i$ and each $t\in[0,T]$, define the noise-state as the conditional mean path of the aggregated primitive noise:
$\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}],\qquad 0\leq u\leq t.$ (2.5)
In linear–Gaussian settings, the conditional law of primitive shocks is Gaussian with deterministic conditional covariance, so $\widehat{W}_{t}^{i}(\cdot)$ is a complete (path-valued) summary of player $i$’s private information.
Definition 2.2 (Primitive-noise Volterra process).
An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Volterra process if there exist deterministic functions
$\bar{L}:[0,T]\to\mathbb{R}^{m},\qquad L:[0,T]^{2}\to\mathbb{R}^{m\times(n+1)d},$
such that, for every $t\in[0,T]$,
$L_{t}\;=\;\bar{L}(t)+\int_{0}^{t}L(t,s)\,dW_{s},\qquad\int_{0}^{t}\|L(t,s)\|^{2}\,ds<\infty.$
When convenient, we extend $L(t,s)$ by 0 for $s>t$.
Remark 2.2 (Properties of primitive-noise Volterra processes).
By construction, primitive-noise Volterra processes are $\mathbb{F}$-adapted Gaussian processes. Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$. Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the Volterra class is that the representation kernel $L(t,\cdot)$ is deterministic (for each fixed $t$).
Definition 2.3 (Noise-state Volterra control / strategy).
Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3). An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state Volterra control if there exist deterministic functions
$\bar{D}^{i}:[0,T]\to\mathbb{R}^{d},\qquad D^{i}:[0,T]^{2}\to\mathbb{R}^{d\times(n+1)d},$
such that, for every $t\in[0,T]$,
$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}(t,u)\|^{2}\,du<\infty,$
and we extend $D^{i}(t,u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
3 Main results
We work in the linear–Gaussian model of Section 2 and establish two closure results—filtering and best response—that reduce Nash equilibrium to a deterministic fixed point in two-time kernels. The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
Filtering closure.
Recall the noise-state $\widehat{W}_{t}^{i}(\cdot)$ (2.5). We record the deterministic-kernel representation of the estimated-noise increment and its innovations gain in noise-state coordinates. Two-time kernels throughout live on the causal triangle $\Delta_{T}:=\{(t,s)\in[0,T]^{2}:0\leq s\leq t\leq T\}$, extended by 0 off $\Delta_{T}$. We write $d_{u}$ for increments in the path index $u$ at fixed estimation time $t$, and $d_{t}$ for updates in $t$ at fixed $u$. The state is assumed to admit the Volterra form $X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s}$, where $X(t,s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
Theorem 3.1 (Volterra Filtering Closure).
Fix a player $i$. Assume the state admits the primitive-noise Volterra form
$X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s},$
and player $i$ observes
$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
with deterministic $P^{i}(t)\succeq 0$ and block selector $E^{i}$.
Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$.
(i) The estimated-noise path increments admit the decomposition
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Big(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Big)\,du,\qquad u<t,$ (3.1)
where the deterministic kernel $F^{i}$ is given explicitly by
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (3.2)
(ii) The mixed $(t,u)$-update satisfies
$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
The unresolved kernel $\widetilde{X}^{i}$ is determined algebraically from $F^{i}$ by (A.12); the filtering kernel $F^{i}$ is the unique solution of the forward evolution system (A.14)–(A.15) (Definition A.1).
Remark 3.1 (Regression interpretation and scale separation).
Discretizing time, each observation increment is affine in the noise increments $\Delta W_{1:k}$, so by joint Gaussianity the conditional mean $\mathbb{E}[\Delta W_{1:k}\mid\Delta Y^{i}_{1:k}]$ is a deterministic linear projection; each layer of “beliefs about beliefs” is a further deterministic composition of projections, which is why the regress closes. The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component—absolutely continuous in $u$ with density driven by $F^{i}$—captures drift-based inference about unobserved noise coordinates at scale $O(dt)$.
Best-response closure and information wedges.
When opponents use noise-state Volterra strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state Volterra; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
Theorem 3.2 (Multiplayer belief-adjoint kernels and information wedges).
Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state Volterra strategy with deterministic kernels, which pins down a deterministic forward kernel environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
Structure. Because the equilibrium control is a Volterra process $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}_{X}^{i},\bar{H}^{k,i})$, one for the kernel coefficients $(H_{X}^{i},H^{k,i})$—that should be read in parallel.
In each subsystem, the single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state. This channel introduces belief-adjoint variables $(\bar{H}^{k,i},H^{k,i})$ that track the shadow value of perturbing opponent $k$’s noise-state estimate, coupled into the physical costate through the opponents’ filtering gains.
Mean backward system. The mean physical costate $\bar{H}_{X}^{i}(t)$ and mean belief-adjoints $\bar{H}^{k,i}(t,u)$ satisfy
$\frac{d}{dt}\bar{H}_{X}^{i}(t)$ $=-\big(G^{i}_{XX}(t)\bar{X}(t)+G_{X}(t)\big)-A(t)^{\top}\bar{H}_{X}^{i}(t)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}(t),\;\text{mean information wedge}},$ (3.3)
$\frac{d}{dt}\bar{H}^{k,i}(t,u)$ $=-\big(D^{k}(t,u)\big)^{\top}\bar{H}_{X}^{i}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz,\qquad k\neq i,$ (3.4)
with terminal conditions $\bar{H}_{X}^{i}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$ and $\bar{H}^{k,i}(T,\cdot)=0$.
The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}(t)$ is the mean information wedge, absent in the single-agent problem. In the belief-adjoint equation (3.4), the forcing $D^{k}(t,u)^{\top}\bar{H}_{X}^{i}(t)$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}(t,u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock. The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
Kernel backward system. The kernel physical costate $H_{X}^{i}(t,r)$ and kernel belief-adjoints $H^{k,i}(t,u,r)$ satisfy the structurally identical system with $\bar{X}(t)\to X(t,r)$, $G_{X}(t)\to 0$, and $\bar{H}^{k,i}(t,u)\to H^{k,i}(t,u,r)$:
$\frac{d}{dt}H_{X}^{i}(t,r)$ $=-\,G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H_{X}^{i}(t,r)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz}_{=:\;\mathcal{V}^{i}(t,r),\;\text{kernel information wedge}},$ (3.5)
$\frac{d}{dt}H^{k,i}(t,u,r)$ $=-\big(D^{k}(t,u)\big)^{\top}H_{X}^{i}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz,\qquad k\neq i,$ (3.6)
with terminal conditions $H_{X}^{i}(T,r)=G^{i}_{XX}(T)X(T,r)$ and $H^{k,i}(T,\cdot,r)=0$.
The two subsystems are identical in structure because they are the Volterra mean and kernel of a single stochastic costate. We define the (full) information wedge as the Volterra process
$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r},$ (3.7)
which is the total distortion of the physical costate relative to the single-agent benchmark. The mean wedge $\bar{\mathcal{V}}^{i}(t)$ captures how belief manipulation distorts the equilibrium level (mean actions); the kernel wedge $\mathcal{V}^{i}(t,r)$ captures how it distorts the equilibrium response to shocks. Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.
Policy interpretation of the wedge.
The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation. When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3). A model that omits this channel predicts the single-agent response where the true equilibrium response includes a strategic distortion whose mean is given by $\bar{\mathcal{V}}^{i}(t)$ and whose shock-by-shock profile is given by $\mathcal{V}^{i}(t,r)$. The mean wedge $\bar{\mathcal{V}}^{i}(t)$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}(t,r)$ governs the distortion in equilibrium impulse responses. Section 4 quantifies both channels numerically.
Corollary 3.3 (Exogenous-signal reduction).
If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4)–(3.6) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
Lemma 3.4 (Why the maximum-principle stationarity condition is global here).
Fix opponents’ strategy maps $D^{-i}$ in the deterministic-kernel class (hence fixed linear operators from signal histories to actions). Then player $i$’s induced objective $D^{i}\mapsto J^{i}(D^{i};D^{-i})$ is a strictly convex functional on the admissible $L^{2}$ control space. Consequently, the spike-variation stationarity condition (B.8) is not only necessary but sufficient: any admissible control satisfying it is the unique global best response.
Proof sketch.
With opponents’ maps fixed and linear, the closed-loop drift of $X$ is affine in $D^{i}$, so $(X,D^{-i})$ depend affinely on $D^{i}$. The cost (2.4) is quadratic with $G^{i}_{DD}(t)\succ 0$, hence strictly convex in $D^{i}$. ∎
Corollary 3.5 (Noise-state Volterra best response (global closure)).
Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state Volterra strategies with deterministic kernels (Definition 2.3). If $D^{i}$ is a best response to $D^{-i}$ over the full admissible $L^{2}$ control class, then for a.e. $t\in[0,T]$,
$G^{i}_{DD}(t)D_{t}^{i}+\bar{H}_{X}^{i}(t)+\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u)=0,$
and hence
$D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\bar{H}_{X}^{i}(t)-(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u).$
In particular, every best response over the full admissible class is itself a noise-state Volterra control with deterministic coefficients.
Proof.
With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12). Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 3.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$. The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula. Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the deterministic kernels mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5). ∎
Remark 3.2.
Since every best response over the full admissible $L^{2}$ class is noise-state Volterra (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state Volterra controls is automatically a Nash equilibrium of the unrestricted game. The open question is whether equilibria outside the deterministic-kernel class exist; Section 6 discusses a program for ruling them out.
Remark 3.3 (Single-player benchmark).
When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H_{X}^{i}(t,r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
Risk-sensitive extension.
Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)). The entire Volterra framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large. The filtering closure is unchanged; separation fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel. The full statement and derivation are in Appendix B.6 (Theorem B.1).
Information cost and endogenous precision.
At the equilibrium profile, the control gap between full and partial information is
$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.8)
where $D^{i,\mathrm{full}}_{t}:=-(G^{i}_{DD}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
Proposition 3.6 (Information cost).
At the equilibrium profile,
$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{i}_{DD}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.9)
where $\Sigma^{M,i}_{t}:=\operatorname{Cov}(M^{i}_{t}\mid\mathcal{F}^{i}_{t})$ is the costate estimation error covariance (deterministic by Gaussianity). Moreover, $\Delta J^{i}$ depends on $P^{i}$ only through the filtering kernels $(\widetilde{X}^{i},F^{i})$ and is independent of player $i$’s control $D^{i}$.
The total cost decomposes as
$J^{i}(D^{i};D^{*,-i})=\underbrace{J^{i}(D^{i,\mathrm{full}};D^{*,-i})}_{\text{certainty-equivalent cost}}+\;\underbrace{\Delta J^{i}(P^{i})}_{\text{information cost}},$ (3.10)
This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{i}_{X}$ inside $\Sigma^{M,i}_{t}$. The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
$\min_{P^{i}(\cdot)\succeq 0}\quad\Delta J^{i}(P^{i})+\int_{0}^{T}c^{i}(t,P^{i}(t))\,\mathrm{d}t,$ (3.11)
where $c^{i}$ is a convex attention cost.
Corollary 3.7 (Endogenous precision closure).
If each opponent $k\neq i$ uses a deterministic precision path $P^{k}(\cdot)$, then $\Delta J^{i}$ is a deterministic functional of $P^{i}(\cdot)$, and the best-response precision is deterministic.
Proof.
Immediate from Proposition 3.6: the forward environment and backward adjoint are deterministic, so (3.11) is a deterministic variational problem. ∎
Remark 3.4 (Connection to rational inattention).
When $n=1$ the FOC for (3.11) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{i}_{X}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
Remark 3.5 (Infinite horizon and the frequency domain).
Under time-homogeneous primitives, Volterra kernels reduce to convolution kernels K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals. Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
Reference: kernel objects.
For convenience, we collect the kernel objects introduced above.
$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}:=E^{i,\top}E^{i}$ selector / projector
$\widehat{W}_{t}^{i}(u)$ noise-state $X(t,s)$ state impulse response
$F_{t}^{i}(u,s)$ filtering kernel $\widetilde{X}^{i}(t,u)$ unresolved state kernel
$D^{i}(t,u)$ policy kernel $\mathcal{D}^{i}(t,s)$ primitive-noise control
$(\bar{H}_{X}^{i},H_{X}^{i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r}$ information wedge
4 Bilateral belief manipulation in a two-player game
To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision. This strips away all structure except the core strategic friction. Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$. Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior. The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}(t)$ is the component we visualize below.
4.1 Primitives, objectives, and symmetry
Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (4.1)
where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the diffusion channel
$dY_{t}^{1}=\sqrt{p}X_{t}\,dt+dW_{t}^{1},\qquad dY_{t}^{2}=\sqrt{p}X_{t}\,dt+dW_{t}^{2},$ (4.2)
with $(W^{0},W^{1},W^{2})$ mutually independent and $p>0$ representing root precision. Player $i$ chooses $D^{i}$ adapted to $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
Running costs are symmetric tracking losses with quadratic effort penalty $r>0$:
$J^{1}(D^{1};D^{2})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}-1)^{2}+r\,(D_{t}^{1})^{2}\Big)\,dt,\quad J^{2}(D^{2};D^{1})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}+1)^{2}+r\,(D_{t}^{2})^{2}\Big)\,dt.$ (4.3)
There is no terminal cost in this worked example.
The model is invariant under the sign/label symmetry
$(X,\;Y^{1},\;Y^{2},\;D^{1},\;D^{2})\ \longmapsto\ (-X,\;Y^{2},\;Y^{1},\;-D^{2},\;-D^{1}).$
[…]
Figure 2: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision $p$.
4.2 Numerical equilibrium and kernel visualization
We now present the numerical equilibrium obtained from the Picard iteration described above. All figures below are generated from the converged kernel profile $D^{1}$, its primitive-noise representation $\mathcal{D}^{1}$, the state kernel $X$, and the filtering kernel $F^{1}$.
Computation.
Equilibrium is obtained by relaxed Picard iteration on the deterministic kernel system, discretized on a uniform grid of $N=40$ points; no Monte Carlo is used. At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H_{X}^{i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile. Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{i}_{XX}(T)=0$. The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++). At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
Convergence of the fixed point.
The outer Picard residual decays geometrically, confirming convergence of the fixed-point scheme in this symmetric benchmark; we use the iteration as a constructive computation rather than an existence proof.
Impulse-response viewpoint.
The key distinction is between $D^{1}(t,s)$ (response to a shock as inferred through the filter).
(a) Primitive-noise state kernel $X(t,s)$.
(b) Equilibrium noise-state feedback kernel $D^{1}(t,s)$.
Figure 3: Equilibrium kernel time-slices, shown channel-by-channel for $(W^{0},W^{1},W^{2})$ with curves colored by evaluation time $t$. (a) Both players jointly stabilize older fundamental shocks (decay in the $W^{0}$ panel); observation-noise channels show smaller, opposed effects mediated by the belief loop. (b) The system-noise channel dominates; observation-noise channels exhibit curvature reflecting drift-based inference.
4.3 Mean actions, separation failure, and bilateral belief manipulation
Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}(t)$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks. Under perfect information, the mean policy does not depend on signal precision—a separation property. Decentralized information breaks this: changing precision changes how fast opponents’ posteriors react to drift, so players choose actions partly to shift opponents’ beliefs, not only to move the physical state.
Figure 4: Equilibrium mean control $\bar{D}^{1}(t)$ as a function of signal precision $p$, with the perfect-information benchmark (dashed). Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
Deadweight loss.
The kernel figures in Section 4 describe the variance channel (through $\mathrm{Var}(X_{t})$) while Figure 4 isolates the mean channel. In this symmetric game the mean channel captures a purely strategic inefficiency: effort spent shifting the opponent’s posterior raises both tracking loss and effort cost relative to the perfect-information benchmark.
The information wedge as a dynamic rent density.
In the general theory (Theorem 3.2), the backward equation for $\bar{H}^{X}$ contains opponent-filter coupling terms that vanish under perfect information or in the single-player case. Here these terms aggregate into
$\bar{\mathcal{V}}^{1}(t)=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}(t,z)\cdot\bar{H}^{2}(t,z)\,dz,$
which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}(t)=-(\bar{X}(t)-1)-\bar{\mathcal{V}}^{1}(t)$. The unresolved kernel $\widetilde{X}^{2}(t,z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}(t,z)$ measures the shadow value to player 1 of that shift. Their inner product $\bar{\mathcal{V}}^{1}(t)$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying. Left: mean controls $\bar{D}^{1}(t)$ (solid) and $\bar{D}^{2}(t)$ (dashed) diverge as the precision gap widens. Center: the mean state path $\bar{X}(t)$ tilts toward the better-informed player’s target. Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}(t)$ and $\bar{\mathcal{V}}^{2}(t)$ as $p_{2}$ varies ($p_{1}=3$ fixed). Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates. Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior. The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying. Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$). Right: cooperative targets ($\theta_{1}=\theta_{2}=0$). Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}(t)$. Blue: player 1; red: player 2.
Information pooling as a Pigouvian intervention.
Figure 7 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive ($\theta_{1}=1$, $\theta_{2}=-1$) and cooperative ($\theta_{1}=\theta_{2}=0$) targets. Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition. Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
5 Market microstructure as a special case
The two-player example operates within the quadratic-cost framework of Section 2. We now show that the noise-state technique embeds the Kyle–Back model with no additional machinery, despite traders maximizing expected profits against the market maker’s price rather than minimizing a quadratic objective. The embedding works because $\lambda(t)>0$ provides the strict convexity needed for the best-response closure, the filtering closure applies verbatim to the stacked observation channel, and the backward system of Appendix D yields the optimality condition. We derive the equilibrium characterization for a single informed trader and identify the two channels through which belief manipulation operates; the extension to multiple asymmetrically informed traders follows the same backward system and is a natural application of the general theory.
5.1 Model and embedding
Adjoin a fundamental-value Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$ and projector $\Pi^{V}$. The fundamental is $V_{t}:=\sigma_{V}W_{t}^{V}$. All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dS_{t}^{i}=\sqrt{p_{i}}\,V_{t}\,dt+dW_{t}^{i}$. The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]=\sigma_{V}\int_{0}^{t}E^{V}d_{u}\widehat{W}_{t}^{0}(u)$. Trader $i$ maximizes $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$. By symmetry $W\mapsto-W$, all mean terms vanish.
Embedding.
Rescaling the order-flow channel to unit noise, trader $i$’s stacked observation is
$d\!\begin{pmatrix}Z_{t}/\sigma_{0}\\ S_{t}^{i}\end{pmatrix}=\underbrace{\begin{pmatrix}\sigma_{0}^{-1}\\[2.0pt] \sqrt{p_{i}}\end{pmatrix}}_{\sqrt{P^{i}}}\!\!\begin{pmatrix}\text{demand drift}\\ \sigma_{V}E^{V}W_{t}\end{pmatrix}\!dt\;+\;\underbrace{\begin{pmatrix}E^{0}\\ E^{i}\end{pmatrix}}_{\mathcal{E}^{i}}\!dW_{t},$ (5.1)
which is the observation equation of Section 2 with $\sqrt{P^{i}}=(\sigma_{0}^{-1},\sqrt{p_{i}})^{\top}$ and $E^{i}\to\mathcal{E}^{i}$; the market maker has $\sqrt{P^{0}}=\sigma_{0}^{-1}$, $\mathcal{E}^{0}=E^{0}$.
Kyle’s lambda.
Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma}(t,s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
$dP_{t}=\lambda(t)\,dI_{t}^{0},\qquad\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du,$ (5.2)
where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12). The execution cost $\mathbb{E}[\int D_{t}^{i}P_{t}\,dt]$ is quadratic in $D_{t}^{i}$ with coefficient $\lambda(t)$—the endogenous $G^{i}_{DD}$.
5.2 Equilibrium characterization
The optimality condition from a unit demand impulse at $t$ equates the conditional mispricing to the integrated impact cost (Appendix D):
$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}(t,u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau}_{H^{i,\mathrm{imp}}(t,u):\;\text{impact cost density}},$ (5.3)
where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}(t,u)$ yields a closed backward ODE:
$\lambda(t)\,D^{i}(t,u)=-\frac{d}{dt}H^{i,\mathrm{val}}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (5.4)
with terminal condition $D^{i}(T,u)=\lambda(T)^{-1}H^{i,\mathrm{val}}(T,u)$.
Assumption 5.1 (Persistent price informativeness).
$\lambda(t)>0$ for all $t\in[0,T]$.
This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity. The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$. It plays the role of $G^{i}_{DD}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix D.
5.3 Two channels of belief manipulation
Decompose $\delta P_{\tau}$ into $\delta P_{\tau}^{\mathrm{dir}}$ (computed with $D^{k}\equiv 0$, i.e. the market maker’s response alone) and $\delta P_{\tau}^{\mathrm{ind}}:=\delta P_{\tau}-\delta P_{\tau}^{\mathrm{dir}}$ (opponent-mediated). Since $\delta P_{\tau}^{\mathrm{dir}}=\lambda(t)$ for all $\tau\geq t$, the impact adjoint splits into:
  1. (i)
    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}(\tau,u)\,d\tau$: instantaneous impact times the integrated future policy kernel. This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
  2. (ii)
    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}(\tau,u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
Back’s model.
When $n=1$ and $p_{1}=0$ (Back’s model), the opponent channel vanishes ($\delta P_{\tau}^{\mathrm{ind}}\equiv 0$) and (5.3) reduces to Back’s integral equation for the optimal trading rate. With $n\geq 2$ and $p_{i}>0$ both channels are active: each trader tempers demand not only to limit direct price impact but to prevent opponents from inferring and front-running through shared order flow. Guess-and-verify cannot close this system; the noise-state technique handles it with no additional machinery.
6 Conclusion
A single coordinate change—conditioning on primitive Brownian shocks—reduces Nash equilibrium in continuous-time LQG games with endogenous signals to a deterministic fixed point in two-time kernels, with no truncation and no large-population limit. The information wedge $\mathcal{V}^{i}_{t}$—with mean component $\bar{\mathcal{V}}^{i}(t)$ and kernel component $\mathcal{V}^{i}(t,r)$—prices bilateral belief manipulation and vanishes when signals are exogenous. The two-player example shows that nearly all welfare gains from mandatory disclosure come from eliminating this channel; the Kyle–Back embedding shows the noise-state technique extends beyond quadratic costs and identifies a second channel—opponent inference through shared order flow—that applies immediately to asymmetric precisions and continuous private learning.
Further directions.
The main open question is whether the deterministic-kernel class is without loss of generality. A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic-kernel class and which iterated best responses are expected to contract. Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
Appendix A Filtering
Roadmap.
Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection. The pipeline proceeds in three steps. A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales. A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}(t,u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds=\Sigma_{XX}^{i}(t)$. A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
A.1 Dynamics of the noise-state
Fix a player $i$. Recall the noise-state $\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}]$, $0\leq u\leq t$, and the observation channel
$dY_{t}^{i}=\Big(\bar{C}(t)+\int_{0}^{t}C(t,s)\,dW_{s}\Big)\,dt+E^{i}\,dW_{t},\qquad Y_{0}^{i}=0,$
where $C(t,s)$ is deterministic. Since $\bar{C}$ is deterministic, we may assume without loss that $\bar{C}\equiv 0$ by replacing $Y_{t}^{i}$ with $Y_{t}^{i}-\int_{0}^{t}\bar{C}(r)\,dr$.
Proposition A.1 (Volterra closure for $\widehat{W}$).
The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a Volterra process: there exists a deterministic kernel $F_{t}^{i}(\cdot,\cdot)$ such that
$\widehat{W}_{t}^{i}(u)=\Pi^{i}W_{u}+\int_{0}^{t}\!\int_{0}^{u}F_{t}^{i}(z,s)\,dz\,dW_{s},\qquad 0\leq u\leq t.$
Proposition A.2 (Dynamics of $\widehat{W}_{t}^{i}(u)$).
There exist deterministic matrix-valued functions $B^{Y}(u,t)$ and $B(u;t,s)$ such that, for each fixed $u$,
$\widehat{W}_{t}^{i}(u)-\widehat{W}_{t_{0}}^{i}(u)=\int_{t_{0}}^{t}\!\!\int_{0}^{r}B(u;r,s)\,d\widehat{W}_{r}^{i}(s)\,dr\;+\;\int_{t_{0}}^{t}B^{Y}(u,r)\,dY_{r}^{i},\qquad t\geq t_{0}\geq u,$ (A.1)
or equivalently in differential form,
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}B(u;t,s)\,d\widehat{W}_{t}^{i}(s)\Big)\,dt+B^{Y}(u,t)\,dY_{t}^{i},$ (A.2)
with
$B^{Y}(u,t)$ $=\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,C(t,s)^{\top}\,ds\;+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}},$ (A.3)
$B(u;t,s)$ $=-B^{Y}(u,t)\,C(t,s),$ (A.4)
where $\mathsf{C}_{t}^{i}(u,s):=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional error covariance kernel.
Proof.
We use the characterizing identity for conditional expectations via exponential test martingales. Fix a deterministic vector function $\vec{r}(\cdot)$ and define $\varphi_{t}^{r}$ by $\varphi_{0}^{r}=1$ and $d\varphi_{t}^{r}=i\,\varphi_{t}^{r}\,(dY_{t}^{i})^{\top}\vec{r}(t)$. A necessary and sufficient condition characterizing $\widehat{W}_{t}^{i}(u)$ is
$\mathbb{E}\!\big[\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big]=\mathbb{E}\!\big[W_{u}\,\varphi_{t}^{r}\big]\qquad\text{for all }\vec{r}(\cdot).$ (A.5)
We compute the time derivative of each side independently, then match.
Left-hand side. Apply Itô’s product rule to $\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}$, using the ansatz (A.2) and $d\varphi_{t}^{r}=i\,\varphi_{t}^{r}(dY_{t}^{i})^{\top}\vec{r}(t)$:
$d\big(\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big)=i\,\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\,(dY_{t}^{i})^{\top}\vec{r}(t)+\varphi_{t}^{r}\,d_{t}\widehat{W}_{t}^{i}(u)+d\langle\widehat{W}^{i}(u),\varphi^{r}\rangle_{t}.$
The quadratic covariation contributes $d\langle\widehat{W}^{i}(u),\varphi^{r}\rangle_{t}=i\,\varphi_{t}^{r}\,B^{Y}(u,t)\,\vec{r}(t)\,dt$. Taking expectations and collecting the $dt$ terms gives
$\frac{d}{dt}\mathbb{E}\!\big[\widehat{W}_{t}^{i}(u)\,\varphi_{t}^{r}\big]$ $=i\,\mathbb{E}\!\bigg[\varphi_{t}^{r}\Big(\widehat{W}_{t}^{i}(u)\!\int_{0}^{t}C(t,s)\,d\widehat{W}_{t}^{i}(s)^{\top}+B^{Y}(u,t)\Big)\bigg]\vec{r}(t)$
$\quad+\mathbb{E}\!\bigg[\varphi_{t}^{r}\int_{0}^{t}\!\big(B(u;t,s)+B^{Y}(u,t)\,C(t,s)\big)\,d\widehat{W}_{t}^{i}(s)\bigg].$ (A.6)
Right-hand side. Expand $\mathbb{E}[W_{u}\,\varphi_{t}^{r}]$ using the SDE for $\varphi^{r}$ and apply the tower property in the form $\mathbb{E}[\varphi_{s}^{r}\,L]=\mathbb{E}[\varphi_{s}^{r}\,\mathbb{E}[L\mid\mathcal{F}_{s}^{i}]]$. Define the conditional second-moment kernel
$\mathcal{M}_{s}^{i}(u,z):=\mathbb{E}[W_{u}W_{z}^{\top}\mid\mathcal{F}_{s}^{i}]=\mathsf{C}_{s}^{i}(u,z)+\widehat{W}_{s}^{i}(u)\,\widehat{W}_{s}^{i}(z)^{\top}.$
Then
$\mathbb{E}\!\Big[W_{u}\!\int_{0}^{s}C(s,z)\,dW_{z}^{\top}\;\Big|\;\mathcal{F}_{s}^{i}\Big]=\int_{0}^{s}\partial_{z}\mathcal{M}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz.$
Splitting $\mathcal{M}_{s}^{i}=\mathsf{C}_{s}^{i}+\widehat{W}_{s}^{i}\,\widehat{W}_{s}^{i,\top}$ separates this into a covariance piece and a product-of-means piece:
$\int_{0}^{s}\partial_{z}\mathcal{M}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz=\int_{0}^{s}\partial_{z}\mathsf{C}_{s}^{i}(u,z)\,C(s,z)^{\top}\,dz\;+\;\widehat{W}_{s}^{i}(u)\!\int_{0}^{s}C(s,z)\,d\widehat{W}_{s}^{i}(z)^{\top}.$ (A.7)
Differentiating $\mathbb{E}[W_{u}\,\varphi_{t}^{r}]$ and collecting $dt$ terms then gives
$\frac{d}{dt}\mathbb{E}\!\big[W_{u}\,\varphi_{t}^{r}\big]$ $=i\,\mathbb{E}\!\bigg[\varphi_{t}^{r}\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,C(t,s)^{\top}\,ds\;+\;\widehat{W}_{t}^{i}(u)\!\int_{0}^{t}C(t,s)\,d\widehat{W}_{t}^{i}(s)^{\top}\;+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\Big)\bigg]\vec{r}(t).$ (A.8)
Coefficient matching. Equating (A.6) and (A.8) for all $\vec{r}(\cdot)$ and all $t$ yields two conditions. Matching the terms multiplying $\vec{r}(t)$ (after canceling the common $\widehat{W}_{t}^{i}(u)\int C\,d\widehat{W}^{i,\top}$ terms that appear on both sides via (A.7)) gives (A.3). Requiring the residual drift in (A.6) to vanish gives
$\int_{0}^{t}\big(B(u;t,s)+B^{Y}(u,t)\,C(t,s)\big)\,d\widehat{W}_{t}^{i}(s)=0,$
hence $B(u;t,s)=-B^{Y}(u,t)\,C(t,s)$, which is (A.4). ∎
A.2 Specialization to the physical observation model
We now specialize to the observation equation of Section 2,
$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ in the general framework of Proposition A.2.
Innovation process.
Define the innovation
$dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt,$
where $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$. By the innovations theorem for Gaussian linear filtering (Liptser–Shiryaev [16], Theorem 7.12), $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion.
Innovation form of the dynamics.
Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X(t,s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
Identifying $\widetilde{X}^{i}$ as the filtering gain.
For $u<t$, the indicator term in (A.9) is locally constant in $u$, so the mixed $(t,u)$-update takes the form
$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}\,dt$. The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$, this picks out $\widetilde{X}^{i}(t,u)$. Thus $\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
Remark A.1 (Conditional cross-covariance interpretation).
The identification above can be restated as
$\widetilde{X}^{i}(t,u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
Integrating in $u$ recovers the full conditional cross-covariance
$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}(t,s)\,ds\;=\;\int_{0}^{t}X(t,s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X(t,s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
Algebraic identity for $\widetilde{X}^{i}$ in terms of $F^{i}$.
Recall the induced estimated-state kernel $\widehat{X}^{\,i}(t,s):=X(t,s)\Pi^{i}+\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u)$,
$\widetilde{X}^{i}(t,u)=X(t,u)(I-\Pi^{i})-\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
This identity is algebraic: once $F^{i}$ is determined (Section A.3), $\widetilde{X}^{i}$ follows with no further fixed-point step.
Remark A.2 (State uncertainty as a contraction).
The conditional state covariance satisfies
$\Sigma_{XX}^{i}(t)=\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds,$ (A.13)
which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$.
Remark A.3 (Kalman-filter interpretation).
Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma_{XX}^{i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}(t,s)$ back through the state impulse response $X(t,s)$. In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance. In the multi-player setting the identity persists because the forward kernel environment is deterministic.
A.3 The filtering kernel $F^{i}$
We define $F^{i}$ as the primitive filtering object through its evolution equation. The unresolved kernel $\widetilde{X}^{i}$ is then a derived quantity via the algebraic identity (A.12).
Definition A.1 (Filtering kernel system).
Fix player $i$. A deterministic causal kernel $F^{i}$ on $\Delta_{T}$ solves the filtering kernel system if, with $\widetilde{X}^{i}$ defined by (A.12), it satisfies $F_{0}^{i}=0$ and
$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),\qquad u,s<t,$ (A.14)
$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u).$ (A.15)
Since (A.12) expresses $\widetilde{X}^{i}$ as a deterministic linear functional of $F^{i}$, the system (A.14)–(A.15) is a closed forward ODE for $F^{i}$ (quadratic through the substitution).
Proposition A.3 (Explicit formula).
Integrating (A.14) from $\max(u,s)$ to $t$ and adding the boundary contributions gives
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.16)
The three terms are: (i) the unresolved response at observation time $s$, projected through the observation channel; (ii) the drift component of the innovation at time $u$, revealing information about earlier increments at $s\leq u$; (iii) accumulated indirect inference from observations at times $r\in[\max(u,s),\,t]$.
Remark A.4 (Self-adjointness).
Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms in (A.16) exchange and the integral is invariant (since $P^{i}(r)$ is symmetric). Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map is an orthogonal projection.
Remark A.5 (Uniqueness).
If $F^{i}$ and $\tilde{F}^{i}$ both yield the density representation of Theorem 3.1(i), Itô isometry gives $F_{t}^{i}(u,s)=\tilde{F}_{t}^{i}(u,s)$ a.e.
Remark A.6 (Computation).
In the numerical scheme of Section 4, $\widetilde{X}^{i}$ is computed from $F^{i}$ via (A.12) at each time step, and the integral in (A.16) is accumulated forward alongside the $F^{i}$ update.
We now convert the innovations-based characterization of Subsections A.1A.2 into primitive-noise coordinates. The goal is to construct the deterministic kernel $F^{i}$ such that for every $t\in[0,T]$ and $u<t$,
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Bigg(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Bigg)\,du.$
A.3.1 Derivation
Recall the innovation form of the estimated-noise dynamics (equation (A.9)): for $u\leq t$,
$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
where the first term is the direct observation at time $u$ and the second is the accumulated drift-based revision over $(u,t]$. To pass from innovations to primitive noise, we expand each $dI_{s}^{i}$ in terms of $dW$.
Expanding the innovation.
The physical-measure decomposition of the innovation is
$dI_{t}^{i}=E^{i}\,dW_{t}\;+\;\sqrt{P^{i}(t)}\,(X_{t}-\widehat{X}_{t}^{i})\,dt.$
The first piece is the martingale component; the second is absolutely continuous. Substituting the Volterra forms and using $\widetilde{X}^{i}(t,r):=X(t,r)-\widehat{X}^{\,i}(t,r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,r)\,dW_{r}\,dt.$ (A.17)
Expanding the direct term.
Substituting (A.17) at $t=u$ into $E^{i,\top}\,dI_{u}^{i}$:
$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}(u,r)\,dW_{r}\,du.$ (A.18)
The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation. The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$ at source-noise index $s\leq u$.
Expanding the integral term.
Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
$\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}(s,r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
Term (a) contributes $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\,\widetilde{X}^{i}(s,r)\,ds\Bigg)\,dW_{r}.$
Assembling the density.
Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}(s,u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.19)
The three terms admit a direct interpretation:
  • $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel. Nonzero only when $u\leq s$ (causal in the observation direction).
  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$. Nonzero only when $s\leq u$ (causal in the noise direction).
  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
Remark A.7 (Self-adjointness and uniqueness).
Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms exchange and the integral is invariant (since $P^{i}(r)$ is symmetric). Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map $(\mathcal{P}_{t}^{i}f)(s):=\Pi^{i}f(s)+\int_{0}^{t}F_{t}^{i}(r,s)^{\top}f(r)\,dr$ is self-adjoint, as required by its interpretation as an orthogonal projection. Uniqueness follows from Itô isometry: if $\tilde{F}$ also satisfies the density representation, then $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$ a.e.
Remark A.8 (Boundary condition and evolution).
Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}(u,t)=0$ for $u<t$):
$F_{t}^{i}(u,t)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u),$
which are each other’s transposes. Differentiating (A.19) in $t$ at fixed $u,s<t$ gives the evolution
$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),$
which is manifestly self-adjoint under $(u,s)\mapsto(s,u)^{\top}$.
Remark A.9 (Computational simplification).
The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration. In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
A.3.2 Formal characterization
Definition A.2 (Filtering kernel system (definition of $F^{i}$)).
Fix player $i$. The inputs are the state kernel $X(\cdot,\cdot)$ and the unresolved kernel $\widetilde{X}^{i}$ from Subsection A.2. The output is the primitive-noise density kernel $F^{i}$ of Theorem 3.1(i).
A deterministic kernel $F^{i}$ is said to be admissible if it is causal ($F_{t}^{i}(u,s)=0$ when $\max\{u,s\}>t$) and, together with the induced kernel
$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
We define $F^{i}$ to be the unique causal solution of (A.12)–(A.20).
Proposition A.4 (Verification: primitive-noise density of the noise-state).
Let $F^{i}$ be the kernel defined by Definition A.1. Then for every $t\in[0,T]$ and $u<t$,
$d_{u}\widehat{W}_{t}^{i}(u)=\Pi^{i}\,dW_{u}+\Bigg(\int_{0}^{t}F_{t}^{i}(u,s)\,dW_{s}\Bigg)\,du.$
Moreover, $F^{i}$ is unique among deterministic causal kernels with this property.
Proof of uniqueness.
If $F^{i}$ and $\tilde{F}^{i}$ both satisfy the density representation, then $\int_{0}^{t}(F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s))\,dW_{s}=0$ a.s. for each $(t,u)$. By Itô isometry, $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$, hence $F_{t}^{i}(u,s)=\tilde{F}_{t}^{i}(u,s)$ a.e. ∎
Appendix B Control First Order Conditions
This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$. By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic kernels weighting the exogenous innovation $I^{i}$. [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
B.1 The best-response problem in innovations coordinates
Fix a baseline deterministic-kernel Volterra profile and a deviating player $i$. By the innovations exogeneity established in Corollary 3.5, $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$, and $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion invariant to $D^{i}$. Player $i$’s control
$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$
is a deterministic affine functional of $I^{i}$, and the optimization is over $(\bar{D}^{i},D^{i})$ with no self-referential signal dependence. Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{i}_{DD}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
B.2 Spike variation and the first-variation system
We use spike variations to expose the adjoint structure. Fix $t\in[0,T)$ and $v\in L^{2}(\Omega,\mathcal{F}_{t}^{i};\mathbb{R}^{d})$. The spike perturbation $D_{s}^{i,\rho,\varepsilon}:=D_{s}^{i}+\rho\,v\,\mathbf{1}_{[t,t+\varepsilon]}(s)$ induces the normalized first variation $\delta X_{s}:=\lim_{\varepsilon\downarrow 0}\frac{1}{\varepsilon}\frac{\partial}{\partial\rho}X_{s}^{\rho,\varepsilon}\big|_{\rho=0}$ with $\delta X_{t}=v$.
Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients—not a linearization but the full response. The deviator’s own innovation is unaffected ($\delta I_{s}^{i}=0$), since $\delta X_{s}\in\mathcal{F}_{t}^{i}$ implies $\delta\widehat{X}_{s}^{i}=\delta X_{s}$.
Opponents’ filter response.
For each $k\neq i$, $\delta(d_{u}\widehat{W}_{s}^{k}(u))=\delta w_{s}^{k}(u)\,du$ with $\delta w_{t}^{k}(\cdot)\equiv 0$ and
$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}(s,u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X(s,r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
B.3 Transition kernels
The coupled system (B.1)–(B.2) has bounded deterministic coefficients and admits a unique deterministic two-parameter evolution family.
Definition B.1 (Block transition kernels).
For $s\geq t$, define $\Phi^{XX}(s,t)\in\mathbb{R}^{d\times d}$ and $\Phi^{Xk}(s,t,u)$ ($k\neq i$, $u\in[0,t]$) by
$\delta X_{s}=\Phi^{XX}(s,t)\,v+\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,u)\,\eta^{k}(u)\,du,$
for initial data $\delta X_{t}=v$, $\delta w_{t}^{k}(\cdot)=\eta^{k}(\cdot)$. In the spike deviation $\eta^{k}\equiv 0$, so $\delta X_{s}=\Phi^{XX}(s,t)\,v$.
These blocks satisfy
$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,dz,$ (B.3)
$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}(t,u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,X(t,u)\,dz,$ (B.4)
with $\Phi^{XX}(t,t)=I$ and $\Phi^{Xk}(t,t,u)=0$.
B.4 First variation of costs and the adjoint kernels
Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the Volterra form of $X_{s}$, the normalized cost variation is
$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
where
$\bar{H}^{X}(t)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.6)
$H^{X}(t,u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}G_{XX}(s)\,X(s,u)\,ds+\Phi^{XX}(T,t)^{\top}G^{i}_{XX}(T)\,X(T,u).$ (B.7)
Stationarity condition (globally necessary and sufficient).
Setting $\delta J_{t}^{i}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives
$G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
with equilibrium coefficients $\bar{D}^{i}(t)=-G_{DD}^{i}(t)^{-1}\bar{H}^{X}(t)$ and $D^{i}(t,u)=-G_{DD}^{i}(t)^{-1}H^{X}(t,u)$.
Equivalently, $D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a separation property).
B.5 Closed backward system for the adjoint kernels
Define the belief-adjoint objects
$\bar{H}^{k}(t,u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.9)
$H^{k}(t,u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}G_{XX}(s)\,X(s,r)\,ds+\Phi^{Xk}(T,t,u)^{\top}G^{i}_{XX}(T)\,X(T,r).$ (B.10)
Then $(\bar{H}^{X},\{\bar{H}^{k}\},H^{X},\{H^{k}\})$ satisfy
$\frac{d}{dt}\bar{H}^{X}(t)$ $=-\bigl[G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr]-A(t)^{\top}\bar{H}^{X}(t)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.11)
$\frac{d}{dt}\bar{H}^{k}(t,u)$ $=-D^{k}(t,u)^{\top}\bar{H}^{X}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.12)
$\frac{d}{dt}H^{X}(t,r)$ $=-G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H^{X}(t,r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.13)
$\frac{d}{dt}H^{k}(t,u,r)$ $=-D^{k}(t,u)^{\top}H^{X}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.14)
with terminal conditions $\bar{H}^{X}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$,  $\bar{H}^{k}(T,\cdot)=0$,  $H^{X}(T,r)=G^{i}_{XX}(T)X(T,r)$,  $H^{k}(T,\cdot,r)=0$.
B.6 Risk-sensitive extension: exponential utility FOC
Theorem B.1 (Risk-sensitive extension).
Replace (2.4) with the entropic risk measure
$J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}\!\left[e^{\,\theta_{i}\mathcal{C}^{i}}\right],\qquad\theta_{i}>0,$ (B.15)
where $\mathcal{C}^{i}$ is the random cost (2.3). A Taylor expansion gives $J^{i}_{\theta_{i}}=\mathbb{E}[\mathcal{C}^{i}]+\tfrac{\theta_{i}}{2}\operatorname{Var}(\mathcal{C}^{i})+O(\theta_{i}^{2})$, so $\theta_{i}$ governs the degree of risk aversion (the risk-neutral case $\theta_{i}\to 0$ recovers (2.4)).
Under a deterministic-kernel Volterra profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
$\widehat{W}^{\,i,\theta}_{t}:=\bigl(I-\theta_{i}C_{t}K_{t}\bigr)^{-1}\!\left(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}\right),$
where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the Volterra profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16). The formula is valid when $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$, which holds for $\theta_{i}$ sufficiently small or $T$ sufficiently short. The filtering structure is unchanged; separation fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
We now derive this result. Replace (2.4) with $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, where $\mathcal{C}^{i}$ is the random cost (2.3). The spike variation of Section B.2 carries over; differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude gives
$\delta J^{i}_{\theta_{i},t}(v)=\mathbb{E}\!\left[Z^{i}\,\delta\mathcal{C}^{i}_{t}(v)\right],\qquad Z^{i}:=\frac{e^{\,\theta_{i}\mathcal{C}^{i}}}{\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]},$
with $\delta\mathcal{C}^{i}_{t}(v)$ identical to (B.5).
Quadratic structure of the cost.
Expanding $\mathcal{C}^{i}$ via the Volterra forms $X_{s}=\bar{X}(s)+\int_{0}^{s}X(s,u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}(s)+\int_{0}^{s}\mathcal{D}^{i}(s,u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
$\mathcal{C}^{i}=c_{0}+\int_{0}^{T}\ell(r)^{\top}dW_{r}+\tfrac{1}{2}\int_{0}^{T}\!\!\int_{0}^{T}dW_{r}^{\top}K(r,s)\,dW_{s},$
where $c_{0}$ absorbs deterministic and trace terms, and the linear and symmetric quadratic kernels are
$\ell(r)$ $:=\int_{r}^{T}\!\Big[X(t,r)^{\top}\bigl(G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\bar{D}^{i}(t)\Big]\,dt$
$\qquad+\;X(T,r)^{\top}\bigl(G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr),$ (B.16)
$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X(t,r)^{\top}G^{i}_{XX}(t)\,X(t,s)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\mathcal{D}^{i}(t,s)\Big]\,dt+X(T,r)^{\top}G^{i}_{XX}(T)\,X(T,s).$ (B.17)
Tilted conditional FOC.
Since $\mathcal{C}^{i}$ is quadratic in $W$, the weight $Z^{i}$ tilts a Gaussian measure by a quadratic form, preserving Gaussianity. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ and applying the tower property gives
$G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
which has the structure of (B.8) under the tilted measure.
Risk-adjusted noise-state.
Since $\mathcal{M}^{i}_{t}$ is linear in $W$ with deterministic kernels, the FOC reduces to computing the tilted conditional mean $\widehat{W}^{\,i,\theta}_{t}:=\mathbb{E}^{Q^{i}}[W_{r}\mid\mathcal{F}_{t}^{i}]$. Under the deterministic-kernel profile, $\mathcal{L}(W\mid\mathcal{F}_{t}^{i})$ is Gaussian with mean $m_{t}:=\widehat{W}_{t}^{i}$ and deterministic covariance $C_{t}:=\mathsf{C}_{t}^{i}$. The quadratic tilt shifts this to
$C_{t,\theta}^{-1}=C_{t}^{-1}-\theta_{i}K_{t},\qquad m_{t,\theta}=C_{t,\theta}(C_{t}^{-1}m_{t}+\theta_{i}k_{t}),$
provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short). Rearranging gives
$\widehat{W}^{\,i,\theta}_{t}=(I-\theta_{i}C_{t}K_{t})^{-1}(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}),$
and substituting into the tilted FOC yields
$\boxed{G^{i}_{DD}(t)D^{i}_{t}\;+\;\bar{H}^{X}(t)\;+\;\int_{0}^{T}H^{X}(t,r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
where
$\widehat{W}^{\,i,\theta}_{t}(r):=\bigl[(I-\theta_{i}C_{t}K_{t})^{-1}\bigl(\widehat{W}_{t}^{i}+\theta_{i}\,C_{t}k_{t}\bigr)\bigr](r).$ (B.19)
The risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ differs from the risk-neutral noise-state $\widehat{W}_{t}^{i}$ in two ways: the linear kernel $k_{t}$ (B.16) shifts the conditional mean of $W$ in the direction that reduces expected cost, while the quadratic kernel $K_{t}$ (B.17) rescales conditional precision through $(I-\theta_{i}C_{t}K_{t})^{-1}$, amplifying the weight on directions of $W$-space along which cost variance is large. As $\theta_{i}\to 0$, $\widehat{W}^{\,i,\theta}_{t}\to\widehat{W}_{t}^{i}$ and (B.18) recovers (B.8).
The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$. Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure. Since $\mathcal{C}^{i}$ is quadratic in $W$ under the Volterra profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
B.7 Information cost: proof of Proposition 3.6
Full- and partial-information controls.
At the equilibrium profile $D^{*}$, the stochastic costate is
$M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u},$
where $(\bar{H}^{i}_{X},H^{i}_{X})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
$D^{i,\mathrm{full}}_{t}$ $=-(G^{i}_{DD}(t))^{-1}M^{i}_{t},$ (B.20)
$D^{i}_{t}$ $=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
Both are evaluated at the same adjoint kernels: opponents’ maps are fixed and linear with deterministic kernels, so the backward system is identical regardless of player $i$’s information. The control gap is
$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
where $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$.
Completion of the square.
Since $D^{i,\mathrm{full}}$ minimizes the strictly convex quadratic $D^{i}\mapsto J^{i}(D^{i};D^{*,-i})$ over $L^{2}(\mathcal{F})$ (Lemma 3.4), and $\delta D^{i}\in L^{2}(\mathcal{F})$, the first variation vanishes and
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{i}_{DD}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
To verify that only $G^{i}_{DD}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$. The zero-order term vanishes by definition. All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{i}_{X})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$. The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{i}_{DD}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal). The standard LQG completion-of-the-square identifies these: using $G^{i}_{DD}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
Kernel formula for $\Sigma^{M,i}_{t}$.
Substituting the control gap into (B.22):
$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{i}_{DD}(t))^{-1}\left(\int_{0}^{t}H^{i}_{X}(t,v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{i}_{DD}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.9). It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
The conditional mean of $M^{i}_{t}$ given $\mathcal{F}^{i}_{t}$ is
$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{i}_{X}(t)+\int_{0}^{t}\mathcal{H}^{i}(t,s)\,\mathrm{d}W_{s},$
where the filtered adjoint kernel is
$\mathcal{H}^{i}(t,s):=H^{i}_{X}(t,s)\,\Pi^{i}+\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s})\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini. Then
$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\bigl(H^{i}_{X}(t,s)-\mathcal{H}^{i}(t,s)\bigr)\,\mathrm{d}W_{s}=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s},$
where $\mathcal{U}^{i}(t,s):=H^{i}_{X}(t,s)(I-\Pi^{i})-\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$. By Itô isometry,
$\Sigma^{M,i}_{t}=\mathbb{E}\!\left[\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)^{\top}\right]=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathcal{U}^{i}(t,s)^{\top}\,\mathrm{d}s.$
Separation from control.
The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$. The adjoint $H^{i}_{X}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself. Both $\mathcal{U}^{i}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof. ∎
Appendix C Well-posedness of the deterministic kernel fixed point
This appendix formulates the equilibrium as a dynamic programming problem on the kernel state space, records the a priori bounds that follow from the projection structure of the filtering kernel, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
Notation.
$C$ denotes a generic constant depending on $(n,d,L,\lambda)$; $C(M)$ when it also depends on a policy-norm bound $M$.
C.1 Standing assumptions
Assumption C.1.
There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{i}_{XX}(t)\|+\|G^{i}_{X}(t)\|+\|G^{i}_{XX}(T)\|+\|G^{i}_{X}(T)\|\leq L$ and $G^{i}_{DD}(t)\succeq\lambda\,I_{d}$.
C.2 The kernel state space and its dynamics
At time $t$ the kernel state is
$\mathcal{Z}_{t}:=\bigl(\bar{X}(t),\;X(t,\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
The unresolved kernel $\widetilde{X}^{j}$ is not a state variable; it is determined algebraically from $(\mathcal{Z}_{t})$ by
$\widetilde{X}^{j}(t,u)=X(t,u)(I-\Pi^{j})-\int_{0}^{t}X(t,z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
Under a noise-state Volterra profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
$\dot{\bar{X}}(t)$ $=A(t)\bar{X}(t)+\textstyle\sum_{k}\bar{D}^{k}(t),$ (C.3)
$\dot{X}(t,s)$ $=A(t)X(t,s)+\textstyle\sum_{k}\mathcal{D}^{k}(t,s),\qquad X(s,s)=\Sigma(s),$ (C.4)
$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}(t,u)^{\top}P^{j}(t)\,\widetilde{X}^{j}(t,s),\qquad F_{0}^{j}=0,$ (C.5)
with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}(t,u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}(t,s):=D^{k}(t,s)\Pi^{k}+\int_{0}^{t}D^{k}(t,u)\,F_{t}^{k}(u,s)\,du$.
The policy $\mathbf{D}$ enters (C.3)–(C.4) but does not enter (C.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (C.2).
C.3 The Hamilton–Jacobi–Bellman equation
Let $V^{i}(t,\mathcal{Z})$ denote the equilibrium continuation cost for player $i$ from kernel state $\mathcal{Z}$ at time $t$. Since the kernel-state dynamics are deterministic, the HJB is a first-order terminal-value problem:
$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (C.6)
where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing, $\ell^{i}$ is the running cost, and $g^{i}$ is the terminal cost, both evaluated at the kernel state.
Player $i$’s control $(\bar{D}^{i}(t),D^{i}(t,\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
$\bar{D}^{i}(t)=-(G^{i}_{DD}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}(t,u)=-(G^{i}_{DD}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}_{X}^{i}(t)$ and $D_{X}V^{i}(\cdot)=H_{X}^{i}(t,\cdot)$, recovering Corollary 3.5.
The information wedge as a shadow price.
The $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term in (C.6) captures the indirect effect of the kernel state’s evolution on player $i$’s continuation cost through opponents’ filters. Along the equilibrium path, differentiating through the chain $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies
$\bar{\mathcal{V}}^{i}(t)=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
i.e. the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel. The kernel wedge $\mathcal{V}^{i}(t,r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
Endogenous precision.
If $P^{i}(\cdot)$ is player $i$’s control subject to attention cost $\int_{0}^{T}c^{i}(t,P^{i}(t))\,dt$, then $P^{i}$ enters (C.6) only through $\dot{F}^{i}$ via (C.5), and the FOC
$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}(t,\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}(t,\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
decouples from the FOC for $D^{i}$: the left side is the marginal value of accelerating the filter, which vanishes as $\mathcal{P}_{t}^{i}\to I$ (Lemma C.3), guaranteeing an interior solution when $\partial c^{i}/\partial P^{i}>0$.
C.4 A priori bounds from projection structure
The filtering kernel $F_{t}^{j}$ parametrizes a portion of the conditional projection $(\mathcal{P}_{t}^{j}f)(u):=\Pi^{j}f(u)+\int_{0}^{t}F_{t}^{j}(u,s)\,f(s)\,ds$, an orthogonal projection on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$ (Remark A.4).
Lemma C.2 (Projection bounds).
For each $j$ and $t\geq 0$: (i) $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$; (ii) both $\mathcal{P}_{t}^{j}$ and $I-\mathcal{P}_{t}^{j}$ are $L^{2}$-contractions; (iii) $\|\mathcal{F}_{t}^{j}\|_{\mathrm{op}}\leq 2$ where $\mathcal{F}_{t}^{j}:=\mathcal{P}_{t}^{j}-\Pi^{j}$.
Proof.
Standard properties of orthogonal projections; (iii) is the triangle inequality. ∎
Lemma C.3 (Dissipation).
$\{\mathcal{P}_{t}^{j}\}_{t\geq 0}$ is non-decreasing in the positive semidefinite order, and $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}([0,t])}\leq\|X(t,\cdot)\|_{L^{2}([0,t])}$ for all $t$.
Proof.
Equation (C.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing. Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (C.2), Lemma C.2(ii) gives the bound. ∎
These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$. The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$. This prevents finite-time blowup despite the quadratic nonlinearity and is the Volterra-coordinate manifestation of Kalman variance collapse.
C.5 Well-posedness
Definition C.1.
$\mathbb{B}_{T}^{M}:=\bigl\{\mathbf{D}=\{(\bar{D}^{i},D^{i})\}_{i=1}^{n}:\;\bar{D}^{i}\in C([0,T];\mathbb{R}^{d}),\;D^{i}\in L^{\infty}(\Delta_{T}),\;\|\mathbf{D}\|\leq M\bigr\}$, where $\|\mathbf{D}\|:=\max_{i}(\|\bar{D}^{i}\|_{C}+\|D^{i}\|_{L^{\infty}})$.
Definition C.2 (Best-response operator).
$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{i}_{DD})^{-1}\bar{H}_{X}^{i,\mathbf{D}},\;-(G^{i}_{DD})^{-1}H_{X}^{i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$. Fixed points are Nash equilibria (Corollary 3.5, Remark 3.2).
The following estimates are proved by Grönwall arguments on the forward system (C.3)–(C.5) (using the a priori bounds of Lemmas C.2C.3 to control the quadratic nonlinearity in (C.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
Proposition C.4 (Bounds and Lipschitz continuity).
Under Assumption C.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
(i) The forward system has a unique global solution with $\|\bar{X}\|_{C}+\|X\|_{L^{\infty}}\leq K_{\mathrm{fwd}}$, $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}}\leq K_{\mathrm{fwd}}\sqrt{T}$, $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$.
(ii) The backward system has a unique solution with $\|\bar{H}_{X}^{i}\|_{C}+\|H_{X}^{i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}_{X}^{i},H_{X}^{i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
Here $K_{\mathrm{fwd}},K_{\mathrm{bwd}}\leq C(M)e^{C(M)T}$.
Proof sketch.
For the forward system: $(\bar{X},X)$ solve linear Volterra equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled. Since $\mathcal{D}^{k}$ involves $F^{k}$ through a Volterra integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (C.2). A Grönwall argument on the aggregate $\Psi(t):=\|\delta X\|_{L^{\infty}(\Delta_{t})}+\max_{j}\sup_{r\leq t}\|\delta\widetilde{X}^{j}(r,\cdot)\|_{L^{2}}$ gives $\Psi(T)\leq C(M)Te^{C(M)T}\|\delta\mathbf{D}\|$; the factor of $T$ arises because the policy enters through integrals of length $\leq T$.
The backward estimates follow from linearity of the adjoint system, Cauchy–Schwarz on coupling terms, and backward Grönwall. ∎
Theorem C.5 (Short-horizon equilibrium).
Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$. In particular, there exists a unique Nash equilibrium in noise-state Volterra strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
Proof.
By Proposition C.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$. Self-mapping ($\mathcal{T}:\mathbb{B}_{T}^{M_{0}}\to\mathbb{B}_{T}^{M_{0}}$) holds because at $M=0$ the belief-adjoint decouples ($D^{k}\equiv 0$ kills the forcing in (B.12)), giving finite $K_{\mathrm{bwd}}(0,T)$; the intermediate value theorem yields $M_{0}$. Since $M_{0}$ is bounded as $T\to 0$, choosing $T^{*}$ small enough makes $\mathrm{Lip}(\mathcal{T})\leq\tfrac{1}{2}$. Banach’s theorem gives existence, uniqueness, and geometric convergence. Corollary 3.5 extends to the full $L^{2}$ class. ∎
Theorem C.6 (Global existence).
Under Assumption C.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
Proof sketch.
Apply the Schauder fixed-point theorem to $\mathcal{T}$ on $\mathbb{B}_{T}^{M_{0}}$, working in the topology $\tau:=(C([0,T])\times L^{2}(\Delta_{T}))^{n}$. Self-mapping holds by the same argument as above. $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits. $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward Volterra equations (via Cauchy–Schwarz) and the linear backward system. $\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}_{X}^{i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]). ∎
Remark C.1 (Uniqueness beyond the contraction regime).
Theorem C.6 gives existence but not uniqueness for $T>T^{*}$. The dissipation of Lemma C.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here. In Section 4, Picard iteration converges to the same limit from multiple initial conditions for all parameter values tested.
Remark C.2 (Stochastic coefficients).
If $A(t,\omega)$, $P^{i}(t,\omega)$, or the cost matrices depend on a public factor process $\xi_{t}$ (e.g. a finite-state Markov chain representing regime switches or stochastic automation events), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$. The projection bounds of Lemmas C.2C.3 are algebraic properties of orthogonal projections and hold pathwise, so the forward well-posedness extends to each sample path between jumps of $\xi$. Each jump triggers a regime transition in which filtering gains, strategies, and the information wedge shift discontinuously, while $\mathcal{Z}$ is continuous across jumps. The backward ODE is replaced by a backward stochastic differential equation in the adjoint kernels, driven by jumps of $\xi$; the deterministic Grönwall arguments extend to pathwise bounds conditional on the regime path.
Appendix D Kyle–Back derivations
This appendix derives the deviation propagation system, the first-order condition (5.3), and the closed backward ODE (5.4) stated in Section 5.
D.1 Drift kernels
The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)$; for trader $k$:
$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
The unresolved kernels $\widetilde{C}^{0}$ and $\widetilde{C}^{k}$ are defined by the projection decomposition (A.12) with $C^{j}$ replacing $\sqrt{P^{j}}X$ and $\Pi_{\mathcal{E}}^{j}$ replacing $\Pi^{j}$.
D.2 Deviation propagation
A spike perturbation of trader $i$’s demand at time $t$ enters aggregate order flow, triggering a coupled forward response. For $s>t$, let $\delta D_{\Sigma,s}$ be the aggregate demand perturbation and $\delta w_{s}^{j}(u)$ the noise-state density perturbation for observer $j$. The deviation system (cf. Section B.2) is:
$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
$\partial_{s}\delta w_{s}^{0}(u)$ $=\widetilde{C}^{0}(s,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,s}-\int_{0}^{s}C^{0}(s,r)\,\delta w_{s}^{0}(r)\,dr\biggr),$ (D.3)
$\partial_{s}\delta w_{s}^{k}(u)$ $=\widetilde{C}^{k}(s,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,s}\\ 0\end{pmatrix}-\int_{0}^{s}C^{k}(s,r)\,\delta w_{s}^{k}(r)\,dr\biggr),\quad k\neq i,$ (D.4)
$\delta P_{s}$ $=\sigma_{V}\int_{0}^{s}E^{V}\delta w_{s}^{0}(u)\,du,$ (D.5)
with $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, $\delta I_{s}^{i}=0$. The demand perturbation enters each observer’s stacked observation through the order-flow row only. Equation (D.2) closes the loop: opponents’ noise-state shifts feed through their policy kernels back into aggregate demand. Since $E^{V}E^{0,\top}=0$, $\delta P_{t}=\lambda(t)$; when $n=1$, $\delta D_{\Sigma,s}=0$ for $s>t$ and $\delta P_{s}=\lambda(t)$ for all $s\geq t$.
D.3 First-order condition
The optimality condition from a unit demand impulse at $t$ is
$\mathbb{E}\!\left[V_{T}-P_{t}-\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau\;\Big|\;\mathcal{F}_{t}^{i}\right]=0.$ (D.6)
The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$). The conditional future demand in Volterra form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}(\tau,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes). Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3). $D^{i}(t,u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$. To pin down $D^{i}(t,u)$, we differentiate both sides in $t$.
D.4 Leibniz differentiation and backward ODE
Impact adjoint.
Leibniz on $H^{i,\mathrm{imp}}$ gives
$\frac{d}{dt}H^{i,\mathrm{imp}}(t,u)=-\lambda(t)\,D^{i}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (D.7)
where the boundary term $\delta P_{t}\cdot D^{i}(t,u)=-\lambda(t)D^{i}(t,u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{i}_{DD}$ and where $D^{i}(t,u)$ first appears explicitly.
Value adjoint.
In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}(t,u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$. Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
$\frac{d}{dt}H^{i,\mathrm{val}}(t,u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
Closed backward ODE.
Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}(t,u)$ yields (5.4). The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$. Convexity (from $\lambda(t)>0$) ensures sufficiency.

### 18 Mar 2026 v2

@@
 In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades.
-We provide the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals.
-Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic two-time kernels, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit.
-The characterization yields an explicit information wedge $\mathcal{V}^{i}_{t}$—a deterministic Volterra process—that prices the marginal value of shifting opponents’ posteriors.
+We provide the first exact equilibrium characterization of finite-player continuous-time linear quadratic games with endogenous signals.
+Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic impulse response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit.
+The characterization yields an explicit information wedge that prices the marginal value of shifting opponents’ posteriors.
 The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
@@
 Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements.
-Under the new policy, signal precision changes, which changes how firms filter, which changes how each firm forecasts the pricing decisions of others, which changes equilibrium inflation dynamics.
+Under the new policy, signal precision changes; firms filter differently, revise their forecasts of one another’s pricing, and equilibrium inflation dynamics shift in tandem.
 Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable.
 The characterization in this paper makes it explicit.
-The same obstruction arises across a wide range of fields: in market microstructure, where prices aggregate dispersed information and trading itself is informative; in dynamic information design, where the mapping from signal structures to equilibrium outcomes is intractable when filtering and equilibrium interact; in contract theory, where multiple agents exert unobservable effort on a shared outcome and draw inferences from it; and in decentralized control—autonomous vehicle fleets, multi-robot systems with misaligned operators—wherever agents observe a shared state through private channels and actions reshape what others sense.
+The same obstruction arises across a wide range of fields.
+In market microstructure, prices aggregate dispersed information and trading itself is informative.
+Dynamic information design faces the same difficulty: the mapping from signal structures to equilibrium outcomes becomes intractable once filtering and equilibrium interact.
+Contract theory encounters it when multiple agents exert unobservable effort on a shared outcome and draw inferences from it.
+So does decentralized control of autonomous vehicle fleets and multi-robot systems with misaligned operators, wherever agents observe a shared state through private channels and actions reshape what others sense.
 Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
@@
 Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning.
-We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by orthogonal projection onto deterministic two-time kernels with no truncation.
-The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the deterministic kernels governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
+We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by linear projection onto deterministic impulse response maps with no truncation.
+The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the maps governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
 In the single-agent problem the conditional mean of the state is a sufficient statistic, but the state estimate is a linear functional of the noise estimate, not the reverse.
@@
 We call the path $\widehat{W}_{t}^{i}(\cdot):=\mathbb{E}[W_{\cdot}\mid\mathcal{F}_{t}^{i}]$ the player’s noise-state; optimal controls are affine functionals of it, just as single-agent controls are affine in the state estimate.
-The main results establish that the best response to any profile of deterministic-kernel strategies is itself deterministic-kernel (closure over the full admissible $L^{2}$ class), so Nash equilibrium reduces to a deterministic fixed point in two-time kernels.
-The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$: a Volterra process with deterministic mean $\bar{\mathcal{V}}^{i}(t)$ and deterministic kernel $\mathcal{V}^{i}(t,r)$, which prices the marginal value to player $i$ of shifting opponents’ posteriors.
-The mean wedge distorts equilibrium mean actions; the kernel wedge distorts equilibrium impulse responses.
-Both vanish identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters.
-The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths $\{P^{i}(\cdot)\}$ and affects the equilibrium only through deterministic filtering sensitivities, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
-(i) Candidate Opponents use Volterra controls (ii) Closure State & filters have det.
-kernels (iii) Deviations Fixed maps $\Rightarrow$ det.
-propagation (iv) Best response FOC $\Rightarrow$ control linear in $\widehat{W}^{\,i}$ (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
-Figure 1: The Volterra fixed-point loop.
-(i) Assume opponents use Volterra controls with deterministic kernels.
-(ii) The state and filtering equations close on deterministic two-time kernels.
+The main results establish that the best response amongst all admissible actions to any profile of linear strategies is itself linear.
+The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$, which prices the marginal value to player $i$ of shifting opponents’ posteriors.
+It vanishes identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters.
+The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths and affects the equilibrium only through deterministic impulse response maps, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
+(i) Candidate Opponents use Noise-State Linear controls (ii) Closure State & filters have det.
+response maps (iii) Deviations Fixed maps $\Rightarrow$ det.
+propagation (iv) Best response FOC $\Rightarrow$ control noise-state linear (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
+Figure 1: The Noise-State Linear fixed-point loop.
+(i) Assume opponents use Noise-state Linear controls.
+(ii) The state and filtering equations have deterministic impulse responses.
 (iii) Under fixed strategy maps, unobserved deviations propagate deterministically.
-(iv) The maximum principle produces a best response that is itself Volterra, closing the loop.
+(iv) The maximum principle produces a best response that is itself noise-state linear, closing the loop.
 (v) A fixed point of this best-response map is a Nash equilibrium.
@@
 Team theory.
-The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for LQG teams.
-Team-theoretic solutions bypass the belief hierarchy because aligned objectives eliminate the incentive to manipulate others’ forecasts.
-The information wedge $\mathcal{V}^{i}_{t}$ is precisely the object that vanishes under team objectives and survives under misaligned payoffs.
+The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for linear quadratic gaussian (LQG) teams.
+Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure of Section 3 never arises.
+The information wedge $\mathcal{V}^{i}_{t}$ is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
 Mean-field LQG games.
@@
 Nimark [22] approximates the higher-order belief hierarchy by truncation at finite order—the closest existing method in spirit to the present paper.
-The deterministic kernel system derived here provides an exact benchmark against which truncation error can be assessed.
+The deterministic impulse-response system derived here provides an exact benchmark against which truncation error can be assessed.
 Applied continuous-time filtering games.
@@
 Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational.
-Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable.
+Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable—the same structure that underlies the workhorse models of [28, 26, 20] and the mean-field LQG literature [7], and that serves as the exact benchmark against which approximations can be assessed.
 2.1 Primitives
@@
   • All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
-Stacked noise and block selectors.
+Notation for primitive shocks.
 We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$.
@@
 An admissible control is $\mathcal{F}^{i}_{t}$-adapted with $\mathbb{E}[\int_{0}^{T}\|D^{i}_{t}\|^{2}\,dt]<\infty$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
-Unresolved state kernel.
+Unresolved impulse response.
 For each player $i$, define
-$\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u),$
+$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
 the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate.
-Formally, $\widetilde{X}^{i}(t,u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
+Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
 Objectives.
+The linear state-cost coefficient $G^{X,i}_{t}=\bar{G}^{X,i}_{t}+\int_{0}^{t}G^{X,i}_{t}(r)\,dW_{r}$ is a stochastic process with deterministic mean $\bar{G}^{X,i}_{t}\in\mathbb{R}^{d}$ and deterministic kernel $G^{X,i}_{t}(r)\in\mathbb{R}^{d\times(n+1)d}$; deterministic coefficients are the special case $G^{X,i}_{t}(r)\equiv 0$.
 Each player $i$ incurs the random cost
-$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{i}_{XX}(t)X_{t}+2(G^{i}_{X}(t))^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{i}_{DD}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{i}_{XX}(T)X_{T}+2G^{i}_{X}(T)X_{T},\end{split}$ (2.3)
-where $G^{i}_{XX}(t)$ and $G^{i}_{DD}(t)$ are symmetric positive semidefinite matrices (with $G^{i}_{DD}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls.
+$\begin{split}\mathcal{C}^{i}(D^{i};D^{-i}):=&\int_{0}^{T}\left(X_{t}^{\top}G^{XX,i}(t)X_{t}+2(G^{X,i}_{t})^{\top}X_{t}+G^{i}(t)+(D^{i}_{t})^{\top}G^{DD,i}(t)D^{i}_{t}\right)dt\\ &+X_{T}^{\top}G^{XX,i}(T)X_{T}+2(G^{X,i}_{T})^{\top}X_{T},\end{split}$ (2.3)
+where $G^{XX,i}(t)$ and $G^{DD,i}(t)$ are symmetric positive semidefinite matrices (with $G^{DD,i}(t)$ positive definite), and $D^{-i}:=(D^{j})_{j\neq i}$ denotes opponents’ controls.
 Each player $i$ seeks to minimize
@@
 Equivalently: in a unilateral deviation by $i$, opponents’ maps are held fixed and are applied to the deviated signal paths.
-In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the Volterra class these maps are parameterized by deterministic kernels that remain unchanged off equilibrium.
+In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps are parameterized by deterministic maps that remain unchanged off equilibrium.
 The requirement that opponents’ strategy maps be held fixed under deviations—rather than their realized actions or their realized beliefs—is the formal expression of the Lucas critique in this environment: if player $i$ changes their control law, the state dynamics change, which changes every opponent’s signal process, filtering problem, and actions, which feeds back into the state.
-This paper aims to establish that the Volterra class is closed under best responses and to derive a deterministic kernel system whose fixed points correspond to Nash equilibria in that class.
+This paper aims to establish that the class of noise-state linear strategies is closed under best responses and to derive a deterministic impulse-response system whose fixed points correspond to Nash equilibria in that class.
+Remark 2.2 (Subjective equilibrium and policy evaluation).
+The best-response closure (Corollary 3.5) does not require player $i$ to know opponents’ maps correctly: a player holding incorrect beliefs about opponents’ strategies still solves a well-posed best-response problem against the believed profile.
+A policy-maker can therefore let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (2.1), which depend on the actual strategy profile rather than on whether agents’ beliefs about each other are correct.
+Common knowledge of rationality is sufficient for Nash equilibrium but is not required to apply the characterization as a best-response operator.
 Noise-state.
@@
 In linear–Gaussian settings, the conditional law of primitive shocks is Gaussian with deterministic conditional covariance, so $\widehat{W}_{t}^{i}(\cdot)$ is a complete (path-valued) summary of player $i$’s private information.
-Definition 2.2 (Primitive-noise Volterra process).
-An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Volterra process if there exist deterministic functions
+Definition 2.2 ((Primitive-noise) Linear Process).
+An $L^{2}$ stochastic process $L=(L_{t})_{t\in[0,T]}$ is a (primitive-noise) Linear process if there exist deterministic functions
 $\bar{L}:[0,T]\to\mathbb{R}^{m},\qquad L:[0,T]^{2}\to\mathbb{R}^{m\times(n+1)d},$
@@
 When convenient, we extend $L(t,s)$ by 0 for $s>t$.
-Remark 2.2 (Properties of primitive-noise Volterra processes).
-By construction, primitive-noise Volterra processes are $\mathbb{F}$-adapted Gaussian processes.
+Remark 2.3 (Impulse-response notation).
+We write $X_{t}(s)$ for the deterministic impulse response, not the random variable $X_{t}$ evaluated at an argument; the shock index $s$ distinguishes the two.
+The same subscript-$t$, parenthetical-shock convention applies to all two-time response maps.
+Remark 2.4 (Properties of primitive-noise linear processes).
+By construction, primitive-noise linear processes are $\mathbb{F}$-adapted Gaussian processes.
 Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$.
-Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the Volterra class is that the representation kernel $L(t,\cdot)$ is deterministic (for each fixed $t$).
-Definition 2.3 (Noise-state Volterra control / strategy).
+Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the linear class is that the response map $L(t,\cdot)$ is deterministic (for each fixed $t$).
+Definition 2.3 (Noise-state Linear control / strategy).
 Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3).
-An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state Volterra control if there exist deterministic functions
+An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state.
+That is, if there exist deterministic functions
 $\bar{D}^{i}:[0,T]\to\mathbb{R}^{d},\qquad D^{i}:[0,T]^{2}\to\mathbb{R}^{d\times(n+1)d},$
 such that, for every $t\in[0,T]$,
-$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}(t,u)\|^{2}\,du<\infty,$
-and we extend $D^{i}(t,u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
-$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
+$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}_{t}(u)\|^{2}\,du<\infty,$
+and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
+$Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
 3 Main results
-We work in the linear–Gaussian model of Section 2 and establish two closure results—filtering and best response—that reduce Nash equilibrium to a deterministic fixed point in two-time kernels.
+We work in the linear–Gaussian model of Section 2.
+Two results drive the analysis.
+First, when opponents use noise-state linear strategies, the filtering equations inherit deterministic impulse responses (Theorem 3.1).
+Second, the best response over the full admissible class is itself noise-state linear (Theorem 3.2), so Nash equilibrium reduces to a deterministic fixed point in impulse-response maps.
 The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
@@
 We write $d_{u}$ for increments in the path index $u$ at fixed estimation time $t$, and $d_{t}$ for updates in $t$ at fixed $u$.
-The state is assumed to admit the Volterra form $X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s}$, where $X(t,s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
-Theorem 3.1 (Volterra Filtering Closure).
-Fix a player $i$. Assume the state admits the primitive-noise Volterra form
-$X_{t}=\bar{X}(t)+\int_{0}^{t}X(t,s)\,dW_{s},$
+The state is assumed to admit the linear form $X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s}$, where $X_{t}(s)\in\mathbb{R}^{d\times(n+1)d}$ is a deterministic impulse-response kernel mapping primitive shocks into the physical state.
+Theorem 3.1 (Linear Filtering Closure).
+Fix a player $i$. Assume the state admits the primitive-noise linear form
+$X_{t}=\bar{X}_{t}+\int_{0}^{t}X_{t}(s)\,dW_{s},$
 and player $i$ observes
@@
 where the deterministic kernel $F^{i}$ is given explicitly by
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (3.2)
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (3.2)
 (ii) The mixed $(t,u)$-update satisfies
-$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
+$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
 The unresolved kernel $\widetilde{X}^{i}$ is determined algebraically from $F^{i}$ by (A.12); the filtering kernel $F^{i}$ is the unique solution of the forward evolution system (A.14)–(A.15) (Definition A.1).
@@
 Best-response closure and information wedges.
-When opponents use noise-state Volterra strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state Volterra; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
+When opponents use noise-state linear strategies, a player’s best response over the full admissible $L^{2}$ class remains noise-state linear; the maximum principle yields a closed backward system of deterministic adjoint kernels including the information wedge $\mathcal{V}^{i}(t)$ that prices bilateral belief manipulation.
 Theorem 3.2 (Multiplayer belief-adjoint kernels and information wedges).
-Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state Volterra strategy with deterministic kernels, which pins down a deterministic forward kernel environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
+Fix a player $i$ and suppose each opponent $k\neq i$ uses a noise-state linear strategy with deterministic impulse-response maps, which pins down a deterministic forward environment (state impulse response $X(\cdot,\cdot)$, filtering objects $(P^{k},\widetilde{X}^{k},F^{k})$, and policy kernels $D^{k}$).
 Structure.
-Because the equilibrium control is a Volterra process $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}_{X}^{i},\bar{H}^{k,i})$, one for the kernel coefficients $(H_{X}^{i},H^{k,i})$—that should be read in parallel.
+Because the equilibrium control is a noise-state linear process $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$, one for the response-map coefficients $(H^{X,i},H^{k,i})$—that should be read in parallel.
 In each subsystem, the single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state.
 This channel introduces belief-adjoint variables $(\bar{H}^{k,i},H^{k,i})$ that track the shadow value of perturbing opponent $k$’s noise-state estimate, coupled into the physical costate through the opponents’ filtering gains.
-Mean backward system. The mean physical costate $\bar{H}_{X}^{i}(t)$ and mean belief-adjoints $\bar{H}^{k,i}(t,u)$ satisfy
-$\frac{d}{dt}\bar{H}_{X}^{i}(t)$ $=-\big(G^{i}_{XX}(t)\bar{X}(t)+G_{X}(t)\big)-A(t)^{\top}\bar{H}_{X}^{i}(t)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}(t),\;\text{mean information wedge}},$ (3.3)
-$\frac{d}{dt}\bar{H}^{k,i}(t,u)$ $=-\big(D^{k}(t,u)\big)^{\top}\bar{H}_{X}^{i}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz,\qquad k\neq i,$ (3.4)
-with terminal conditions $\bar{H}_{X}^{i}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$ and $\bar{H}^{k,i}(T,\cdot)=0$.
-The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}(t)$ is the mean information wedge, absent in the single-agent problem.
-In the belief-adjoint equation (3.4), the forcing $D^{k}(t,u)^{\top}\bar{H}_{X}^{i}(t)$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}(t,u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock.
+Mean backward system. The mean physical costate $\bar{H}^{X,i}_{t}$ and mean belief-adjoints $\bar{H}^{k,i}_{t}(u)$ satisfy
+$\frac{d}{dt}\bar{H}^{X,i}_{t}$ $=-\big(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\big)-A(t)^{\top}\bar{H}^{X,i}_{t}-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}_{t},\;\text{mean information wedge}},$ (3.3)
+$\frac{d}{dt}\bar{H}^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}\bar{H}^{X,i}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (3.4)
+with terminal conditions $\bar{H}^{X,i}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$ and $\bar{H}^{k,i}_{T}(\cdot)=0$.
+The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}_{t}$ is the mean information wedge, absent in the single-agent problem.
+In the belief-adjoint equation (3.4), the forcing $D^{k}_{t}(u)^{\top}\bar{H}^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock.
 The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
 Kernel backward system.
-The kernel physical costate $H_{X}^{i}(t,r)$ and kernel belief-adjoints $H^{k,i}(t,u,r)$ satisfy the structurally identical system with $\bar{X}(t)\to X(t,r)$, $G_{X}(t)\to 0$, and $\bar{H}^{k,i}(t,u)\to H^{k,i}(t,u,r)$:
-$\frac{d}{dt}H_{X}^{i}(t,r)$ $=-\,G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H_{X}^{i}(t,r)-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz}_{=:\;\mathcal{V}^{i}(t,r),\;\text{kernel information wedge}},$ (3.5)
-$\frac{d}{dt}H^{k,i}(t,u,r)$ $=-\big(D^{k}(t,u)\big)^{\top}H_{X}^{i}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\widetilde{X}^{k}(t,z)\,H^{k,i}(t,z,r)\,dz,\qquad k\neq i,$ (3.6)
-with terminal conditions $H_{X}^{i}(T,r)=G^{i}_{XX}(T)X(T,r)$ and $H^{k,i}(T,\cdot,r)=0$.
-The two subsystems are identical in structure because they are the Volterra mean and kernel of a single stochastic costate. We define the (full) information wedge as the Volterra process
-$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r},$ (3.7)
+The kernel costate $H^{X,i}_{t}(r)$ and kernel belief-adjoints $H^{k,i}_{t}(u,r)$ satisfy the identical system with $\bar{X}_{t}\to X(t,r)$ and $\bar{G}^{X,i}_{t}\to G^{X,i}_{t}(r)$; the kernel information wedge $\mathcal{V}^{i}_{t}(r)$ is the corresponding substitution of $\bar{\mathcal{V}}^{i}_{t}$.
+We define the (full) information wedge as the linear process
+$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r},$ (3.5)
 which is the total distortion of the physical costate relative to the single-agent benchmark.
-The mean wedge $\bar{\mathcal{V}}^{i}(t)$ captures how belief manipulation distorts the equilibrium level (mean actions); the kernel wedge $\mathcal{V}^{i}(t,r)$ captures how it distorts the equilibrium response to shocks.
+The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ captures how belief manipulation distorts the equilibrium level (mean actions); the response-map wedge $\mathcal{V}^{i}_{t}(r)$ captures how it distorts the equilibrium response to shocks.
 Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.
@@
 When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3).
-A model that omits this channel predicts the single-agent response where the true equilibrium response includes a strategic distortion whose mean is given by $\bar{\mathcal{V}}^{i}(t)$ and whose shock-by-shock profile is given by $\mathcal{V}^{i}(t,r)$.
-The mean wedge $\bar{\mathcal{V}}^{i}(t)$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}(t,r)$ governs the distortion in equilibrium impulse responses.
+Any model that omits this channel misattributes the strategic distortion to the physical response: the mean error is exactly $\bar{\mathcal{V}}^{i}_{t}$, and the shock-by-shock error is $\mathcal{V}^{i}_{t}(r)$.
+The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}_{t}(r)$ governs the distortion in equilibrium impulse responses.
 Section 4 quantifies both channels numerically.
 Corollary 3.3 (Exogenous-signal reduction).
-If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4)–(3.6) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
-Lemma 3.4 (Why the maximum-principle stationarity condition is global here).
-Fix opponents’ strategy maps $D^{-i}$ in the deterministic-kernel class (hence fixed linear operators from signal histories to actions).
+If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (3.4) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
+Lemma 3.4 (Strict convexity of the best-response problem).
+Fix opponents’ strategy maps $D^{-i}$ in the deterministic impulse-response class (hence fixed linear operators from signal histories to actions).
 Then player $i$’s induced objective $D^{i}\mapsto J^{i}(D^{i};D^{-i})$ is a strictly convex functional on the admissible $L^{2}$ control space.
@@
 With opponents’ maps fixed and linear, the closed-loop drift of $X$ is affine in $D^{i}$, so $(X,D^{-i})$ depend affinely on $D^{i}$.
-The cost (2.4) is quadratic with $G^{i}_{DD}(t)\succ 0$, hence strictly convex in $D^{i}$.
+The cost (2.4) is quadratic with $G^{DD,i}(t)\succ 0$, hence strictly convex in $D^{i}$.
 ∎
-Corollary 3.5 (Noise-state Volterra best response (global closure)).
-Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state Volterra strategies with deterministic kernels (Definition 2.3).
+Corollary 3.5 (Noise-state linear best response (global closure)).
+Under the conditions of Theorem 3.2, fix an opponents’ profile $D^{-i}$ of noise-state linear strategies (Definition 2.3).
 If $D^{i}$ is a best response to $D^{-i}$ over the full admissible $L^{2}$ control class, then for a.e.
 $t\in[0,T]$,
-$G^{i}_{DD}(t)D_{t}^{i}+\bar{H}_{X}^{i}(t)+\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u)=0,$
+$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0,$
 and hence
-$D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\bar{H}_{X}^{i}(t)-(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H_{X}^{i}(t,u)\,d\widehat{W}_{t}^{i}(u).$
-In particular, every best response over the full admissible class is itself a noise-state Volterra control with deterministic coefficients.
+$D_{t}^{i}=-(G^{DD,i}(t))^{-1}\bar{H}^{X,i}_{t}-(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u).$
+In particular, every best response over the full admissible class is itself a noise-state linear control.
 Proof.
@@
 The induced objective is strictly convex (Lemma 3.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula.
-Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the deterministic kernels mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5).
+Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5).
 ∎
 Remark 3.2.
-Since every best response over the full admissible $L^{2}$ class is noise-state Volterra (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state Volterra controls is automatically a Nash equilibrium of the unrestricted game.
-The open question is whether equilibria outside the deterministic-kernel class exist; Section 6 discusses a program for ruling them out.
-Remark 3.3 (Single-player benchmark).
-When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H_{X}^{i}(t,r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
+Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state linear controls is automatically a Nash equilibrium of the unrestricted game.
+The open question is whether equilibria outside the deterministic impulse-response class exist; Section 6 discusses a program for ruling them out.
+When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H^{X,i}_{t}(r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
 Risk-sensitive extension.
 Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)).
-The entire Volterra framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large.
-The filtering closure is unchanged; separation fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel.
+The entire framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large.
+The filtering closure is unchanged; certainty equivalence fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel.
 The full statement and derivation are in Appendix B.6 (Theorem B.1).
@@
 At the equilibrium profile, the control gap between full and partial information is
-$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.8)
-where $D^{i,\mathrm{full}}_{t}:=-(G^{i}_{DD}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
+$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.6)
+where $D^{i,\mathrm{full}}_{t}:=-(G^{DD,i}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
 Proposition 3.6 (Information cost).
 At the equilibrium profile,
-$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{i}_{DD}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.9)
+$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{DD,i}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.7)
 where $\Sigma^{M,i}_{t}:=\operatorname{Cov}(M^{i}_{t}\mid\mathcal{F}^{i}_{t})$ is the costate estimation error covariance (deterministic by Gaussianity).
@@
 $J^{i}(D^{i};D^{*,-i})=\underbrace{J^{i}(D^{i,\mathrm{full}};D^{*,-i})}_{\text{certainty-equivalent cost}}+\;\underbrace{\Delta J^{i}(P^{i})}_{\text{information cost}},$ (3.8)
-This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{i}_{X}$ inside $\Sigma^{M,i}_{t}$.
+This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{X,i}$ inside $\Sigma^{M,i}_{t}$.
 The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
@@
 Remark 3.3 (Connection to rational inattention).
-When $n=1$ the FOC for (3.11) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{i}_{X}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
+When $n=1$ the FOC for (3.9) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{X,i}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
 Remark 3.4 (Infinite horizon and the frequency domain).
-Under time-homogeneous primitives, Volterra kernels reduce to convolution kernels K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals.
+Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals.
 Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
-Reference: kernel objects.
+Reference: impulse-response objects.
 For convenience, we collect the kernel objects introduced above.
 $W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}:=E^{i,\top}E^{i}$ selector / projector
-$\widehat{W}_{t}^{i}(u)$ noise-state $X(t,s)$ state impulse response
-$F_{t}^{i}(u,s)$ filtering kernel $\widetilde{X}^{i}(t,u)$ unresolved state kernel
-$D^{i}(t,u)$ policy kernel $\mathcal{D}^{i}(t,s)$ primitive-noise control
-$(\bar{H}_{X}^{i},H_{X}^{i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
-$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}(t)+\int_{0}^{t}\mathcal{V}^{i}(t,r)\,dW_{r}$ information wedge
+$\widehat{W}_{t}^{i}(u)$ noise-state $X_{t}(s)$ state impulse response
+$F_{t}^{i}(u,s)$ filtering impulse response $\widetilde{X}^{i}_{t}(u)$ unresolved state response
+$D^{i}_{t}(u)$ policy impulse response $\mathcal{D}^{i}_{t}(s)$ primitive-noise control
+$(\bar{H}^{X,i},H^{X,i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
+$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r}$ information wedge
 4 Bilateral belief manipulation in a two-player game
@@
 Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior.
-The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}(t)$ is the component we visualize below.
+The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the component we visualize below.
 4.1 Primitives, objectives, and symmetry
@@
 Computation.
-Equilibrium is obtained by relaxed Picard iteration on the deterministic kernel system, discretized on a uniform grid of $N=40$ points; no Monte Carlo is used.
-At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H_{X}^{i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile.
-Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{i}_{XX}(T)=0$.
+Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system, discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed.
+At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H^{X,i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile.
+Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{XX,i}(T)=0$.
 The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++).
@@
 Impulse-response viewpoint.
-The key distinction is between $D^{1}(t,s)$ (response to a shock as inferred through the filter).
-(a) Primitive-noise state kernel $X(t,s)$.
-(b) Equilibrium noise-state feedback kernel $D^{1}(t,s)$.
+The key distinction is between $D^{1}_{t}(s)$ (response to a shock as inferred through the filter).
+(a) Primitive-noise state kernel $X_{t}(s)$.
+(b) Equilibrium noise-state feedback kernel $D^{1}_{t}(s)$.
 Figure 3: Equilibrium kernel time-slices, shown channel-by-channel for $(W^{0},W^{1},W^{2})$ with curves colored by evaluation time $t$.
@@
 4.3 Mean actions, separation failure, and bilateral belief manipulation
-Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}(t)$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks.
-Under perfect information, the mean policy does not depend on signal precision—a separation property.
-Decentralized information breaks this: changing precision changes how fast opponents’ posteriors react to drift, so players choose actions partly to shift opponents’ beliefs, not only to move the physical state.
-Figure 4: Equilibrium mean control $\bar{D}^{1}(t)$ as a function of signal precision $p$, with the perfect-information benchmark (dashed).
+Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks.
+Under perfect information, the mean policy does not depend on signal precision—a certainty-equivalence property.
+Decentralized information breaks this: changing precision changes the speed at which opponents update, giving each player a reason to act on posteriors rather than just on the physical state.
+Figure 4: Equilibrium mean control $\bar{D}^{1}_{t}$ as a function of signal precision $p$, with the perfect-information benchmark (dashed).
 Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
-Deadweight loss.
-The kernel figures in Section 4 describe the variance channel (through $\mathrm{Var}(X_{t})$) while Figure 4 isolates the mean channel.
-In this symmetric game the mean channel captures a purely strategic inefficiency: effort spent shifting the opponent’s posterior raises both tracking loss and effort cost relative to the perfect-information benchmark.
 The information wedge as a dynamic rent density.
@@
 Here these terms aggregate into
-$\bar{\mathcal{V}}^{1}(t)=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}(t,z)\cdot\bar{H}^{2}(t,z)\,dz,$
-which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}(t)=-(\bar{X}(t)-1)-\bar{\mathcal{V}}^{1}(t)$.
-The unresolved kernel $\widetilde{X}^{2}(t,z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}(t,z)$ measures the shadow value to player 1 of that shift.
-Their inner product $\bar{\mathcal{V}}^{1}(t)$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
+$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$
+which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$.
+The unresolved kernel $\widetilde{X}^{2}_{t}(z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}_{t}(z)$ measures the shadow value to player 1 of that shift.
+Their inner product $\bar{\mathcal{V}}^{1}_{t}$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
 Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying.
-Left: mean controls $\bar{D}^{1}(t)$ (solid) and $\bar{D}^{2}(t)$ (dashed) diverge as the precision gap widens.
-Center: the mean state path $\bar{X}(t)$ tilts toward the better-informed player’s target.
+Left: mean controls $\bar{D}^{1}_{t}$ (solid) and $\bar{D}^{2}_{t}$ (dashed) diverge as the precision gap widens.
+Center: the mean state path $\bar{X}_{t}$ tilts toward the better-informed player’s target.
 Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
-Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}(t)$ and $\bar{\mathcal{V}}^{2}(t)$ as $p_{2}$ varies ($p_{1}=3$ fixed).
+Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}_{t}$ and $\bar{\mathcal{V}}^{2}_{t}$ as $p_{2}$ varies ($p_{1}=3$ fixed).
 Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates.
@@
 Right: cooperative targets ($\theta_{1}=\theta_{2}=0$).
-Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}(t)$.
+Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$.
 Blue: player 1; red: player 2.
@@
 Kyle’s lambda.
-Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma}(t,s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
+Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma,t}(s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
 $dP_{t}=\lambda(t)\,dI_{t}^{0},\qquad\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du,$ (5.2)
 where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12).
-The execution cost $\mathbb{E}[\int D_{t}^{i}P_{t}\,dt]$ is quadratic in $D_{t}^{i}$ with coefficient $\lambda(t)$—the endogenous $G^{i}_{DD}$.
+Remark 5.1 (Quadratic structure of the profit).
+Fix opponents’ strategy maps.
+Since the market maker’s filtering gains are deterministic, the price decomposes as $P_{t}=P_{t}^{-i}+\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds$, where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ and $\ell(t,s)$ is a deterministic kernel.
+Symmetrically, the future execution cost of a unit impulse at time $t$ is $\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau$, where $\delta P_{\tau}$ propagates forward through the deviation system with deterministic coefficients.
+The profit $\pi^{i}=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$ is therefore concave quadratic in $D^{i}$, when $\lambda(t)>0$ provides strict concavity.
 5.2 Equilibrium characterization
 The optimality condition from a unit demand impulse at $t$ equates the conditional mispricing to the integrated impact cost (Appendix D):
-$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}(t,u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau}_{H^{i,\mathrm{imp}}(t,u):\;\text{impact cost density}},$ (5.3)
-where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}(t,u)$ yields a closed backward ODE:
-$\lambda(t)\,D^{i}(t,u)=-\frac{d}{dt}H^{i,\mathrm{val}}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (5.4)
-with terminal condition $D^{i}(T,u)=\lambda(T)^{-1}H^{i,\mathrm{val}}(T,u)$.
+$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau}_{H^{i,\mathrm{imp}}_{t}(u):\;\text{impact cost density}},$ (5.3)
+where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}_{t}(u)$ yields a closed backward ODE:
+$\lambda(t)\,D^{i}_{t}(u)=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (5.4)
+with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
 Assumption 5.1 (Persistent price informativeness).
@@
 The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$.
-It plays the role of $G^{i}_{DD}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
+It plays the role of $G^{DD,i}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
 The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix D.
@@
   1. (i)
-    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}(\tau,u)\,d\tau$: instantaneous impact times the integrated future policy kernel.
+    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}_{\tau}(u)\,d\tau$: instantaneous impact times the integrated future policy kernel.
 This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
   2. (ii)
-    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}(\tau,u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
+    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}_{\tau}(u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
 Back’s model.
@@
 6 Conclusion
-A single coordinate change—conditioning on primitive Brownian shocks—reduces Nash equilibrium in continuous-time LQG games with endogenous signals to a deterministic fixed point in two-time kernels, with no truncation and no large-population limit.
-The information wedge $\mathcal{V}^{i}_{t}$—with mean component $\bar{\mathcal{V}}^{i}(t)$ and kernel component $\mathcal{V}^{i}(t,r)$—prices bilateral belief manipulation and vanishes when signals are exogenous.
-The two-player example shows that nearly all welfare gains from mandatory disclosure come from eliminating this channel; the Kyle–Back embedding shows the noise-state technique extends beyond quadratic costs and identifies a second channel—opponent inference through shared order flow—that applies immediately to asymmetric precisions and continuous private learning.
+Conditioning on primitive Brownian shocks collapses the Townsend belief hierarchy onto deterministic impulse-response maps, yielding the first exact equilibrium characterization of finite-player continuous-time LQG games with endogenous signals.
+The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous.
+The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs, identifying a second channel through opponent inference via shared order flow that applies to asymmetric precisions and continuous private learning.
 Further directions.
-The main open question is whether the deterministic-kernel class is without loss of generality.
-A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic-kernel class and which iterated best responses are expected to contract.
+The main open question is whether the deterministic impulse-response class is without loss of generality.
+A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class and which iterated best responses are expected to contract.
 Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
@@
 A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales.
-A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}(t,u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds=\Sigma_{XX}^{i}(t)$.
+A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds=\Sigma^{XX,i}(t)$.
 A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
@@
 where $C(t,s)$ is deterministic. Since $\bar{C}$ is deterministic, we may assume without loss that $\bar{C}\equiv 0$ by replacing $Y_{t}^{i}$ with $Y_{t}^{i}-\int_{0}^{t}\bar{C}(r)\,dr$.
-Proposition A.1 (Volterra closure for $\widehat{W}$).
-The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a Volterra process: there exists a deterministic kernel $F_{t}^{i}(\cdot,\cdot)$ such that
+Proposition A.1 (Linear closure for $\widehat{W}$).
+The map $u\mapsto\widehat{W}_{t}^{i}(u)$ is a linear process: there exists a deterministic indirect-response map $F_{t}^{i}(\cdot,\cdot)$ such that
 $\widehat{W}_{t}^{i}(u)=\Pi^{i}W_{u}+\int_{0}^{t}\!\int_{0}^{u}F_{t}^{i}(z,s)\,dz\,dW_{s},\qquad 0\leq u\leq t.$
@@
 $dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
-which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ in the general framework of Proposition A.2.
+which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ in the general framework of Proposition A.2.
 Innovation process.
@@
 Innovation form of the dynamics.
-Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X(t,s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
-$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X(t,s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
+Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
+$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
 Identifying $\widetilde{X}^{i}$ as the filtering gain.
 For $u<t$, the indicator term in (A.9) is locally constant in $u$, so the mixed $(t,u)$-update takes the form
-$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
-To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}\,dt$.
-The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$, this picks out $\widetilde{X}^{i}(t,u)$.
-Thus $\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
+$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
+To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$.
+The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$.
+Thus $\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
 Remark A.1 (Conditional cross-covariance interpretation).
 The identification above can be restated as
-$\widetilde{X}^{i}(t,u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
+$\widetilde{X}^{i}_{t}(u)\;=\;\frac{\operatorname{Cov}\!\big(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i}\big)}{du}.$
 Integrating in $u$ recovers the full conditional cross-covariance
-$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}(t,s)\,ds\;=\;\int_{0}^{t}X(t,s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
-where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X(t,s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
+$\operatorname{Cov}\!\big(X_{t},\,W_{u}\mid\mathcal{F}_{t}^{i}\big)\;=\;\int_{0}^{u}\widetilde{X}^{i}_{t}(s)\,ds\;=\;\int_{0}^{t}X_{t}(s)\,\partial_{s}\mathsf{C}_{t}^{i}(s,u)\,ds,$ (A.11)
+where the second equality follows from $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}X_{t}(s)\,d\widetilde{W}_{t}^{i}(s)$ and Itô isometry against $\widetilde{W}_{t}^{i}(u)$.
 Algebraic identity for $\widetilde{X}^{i}$ in terms of $F^{i}$.
-Recall the induced estimated-state kernel $\widehat{X}^{\,i}(t,s):=X(t,s)\Pi^{i}+\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}(t,u):=X(t,u)-\widehat{X}^{\,i}(t,u)$,
-$\widetilde{X}^{i}(t,u)=X(t,u)(I-\Pi^{i})-\int_{0}^{t}X(t,z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
+Recall the induced estimated-state kernel $\widehat{X}^{i}_{t}(s):=X_{t}(s)\Pi^{i}+\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,s)\,dz$. Since $\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u)$,
+$\widetilde{X}^{i}_{t}(u)=X_{t}(u)(I-\Pi^{i})-\int_{0}^{t}X_{t}(z)\,F_{t}^{i}(z,u)\,dz.$ (A.12)
 This identity is algebraic: once $F^{i}$ is determined (Section A.3), $\widetilde{X}^{i}$ follows with no further fixed-point step.
@@
 The conditional state covariance satisfies
-$\Sigma_{XX}^{i}(t)=\int_{0}^{t}X(t,s)\,\widetilde{X}^{i}(t,s)^{\top}\,ds,$ (A.13)
-which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}(t,s)\,dW_{s}$.
+$\Sigma^{XX,i}(t)=\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds,$ (A.13)
+which follows from Itô isometry applied to $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$.
 Remark A.3 (Kalman-filter interpretation).
-Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma_{XX}^{i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}(t,s)$ back through the state impulse response $X(t,s)$.
+Identity (A.13) says that the $d\times d$ state uncertainty $\Sigma^{XX,i}(t)$ is obtained by “projecting” the unresolved kernel $\widetilde{X}^{i}_{t}(s)$ back through the state impulse response $X_{t}(s)$.
 In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance.
@@
 Fix player $i$. A deterministic causal kernel $F^{i}$ on $\Delta_{T}$ solves the filtering kernel system if, with $\widetilde{X}^{i}$ defined by (A.12), it satisfies $F_{0}^{i}=0$ and
-$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),\qquad u,s<t,$ (A.14)
-$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u).$ (A.15)
+$\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),\qquad u,s<t,$ (A.14)
+$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u).$ (A.15)
 Since (A.12) expresses $\widetilde{X}^{i}$ as a deterministic linear functional of $F^{i}$, the system (A.14)–(A.15) is a closed forward ODE for $F^{i}$ (quadratic through the substitution).
@@
 Integrating (A.14) from $\max(u,s)$ to $t$ and adding the boundary contributions gives
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.16)
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
 The three terms are: (i) the unresolved response at observation time $s$, projected through the observation channel; (ii) the drift component of the innovation at time $u$, revealing information about earlier increments at $s\leq u$; (iii) accumulated indirect inference from observations at times $r\in[\max(u,s),\,t]$.
@@
 Recall the innovation form of the estimated-noise dynamics (equation (A.9)): for $u\leq t$,
-$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
+$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
 where the first term is the direct observation at time $u$ and the second is the accumulated drift-based revision over $(u,t]$.
@@
 The first piece is the martingale component; the second is absolutely continuous.
-Substituting the Volterra forms and using $\widetilde{X}^{i}(t,r):=X(t,r)-\widehat{X}^{\,i}(t,r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
-$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}(t,r)\,dW_{r}\,dt.$ (A.17)
+Substituting the linear forms and using $\widetilde{X}^{i}_{t}(r):=X_{t}(r)-\widehat{X}^{i}_{t}(r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
+$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(r)\,dW_{r}\,dt.$ (A.17)
 Expanding the direct term.
 Substituting (A.17) at $t=u$ into $E^{i,\top}\,dI_{u}^{i}$:
-$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}(u,r)\,dW_{r}\,du.$ (A.18)
+$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
 The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation.
-The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$ at source-noise index $s\leq u$.
+The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
 Expanding the integral term.
-Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
-$\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
-$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}(s,r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
-Term (a) contributes $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
-$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}(s,u)^{\top}P^{i}(s)\,\widetilde{X}^{i}(s,r)\,ds\Bigg)\,dW_{r}.$
+Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
+$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
+$\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}_{s}(r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
+Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
+$\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\,\widetilde{X}^{i}_{s}(r)\,ds\Bigg)\,dW_{r}.$
 Assembling the density.
-Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}(s,u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr.$ (A.19)
+Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}_{s}(u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
 The three terms admit a direct interpretation:
-  • $\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel.
+  • $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel.
 Nonzero only when $u\leq s$ (causal in the observation direction).
-  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$.
+  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$.
 Nonzero only when $s\leq u$ (causal in the noise direction).
-  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
+  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
 Remark A.7 (Self-adjointness and uniqueness).
@@
 Remark A.8 (Boundary condition and evolution).
-Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}(u,t)=0$ for $u<t$):
-$F_{t}^{i}(u,t)=\widetilde{X}^{i}(t,u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}(t,u),$
+Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}_{u}(t)=0$ for $u<t$):
+$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u),$
 which are each other’s transposes. Differentiating (A.19) in $t$ at fixed $u,s<t$ gives the evolution
-$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}(t,u)^{\top}P^{i}(t)\,\widetilde{X}^{i}(t,s),$
+$\partial_{t}F_{t}^{i}(u,s)=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),$
 which is manifestly self-adjoint under $(u,s)\mapsto(s,u)^{\top}$.
@@
 The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration.
-In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
+In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
 A.3.2 Formal characterization
@@
 A deterministic kernel $F^{i}$ is said to be admissible if it is causal ($F_{t}^{i}(u,s)=0$ when $\max\{u,s\}>t$) and, together with the induced kernel
-$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}(s,u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}(u,s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}(r,u)^{\top}P^{i}(r)\,\widetilde{X}^{i}(r,s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
+$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
 We define $F^{i}$ to be the unique causal solution of (A.12)–(A.20).
@@
 This appendix derives the stationarity condition (B.8) and the closed backward system for the adjoint kernels $(\bar{H}^{X},H^{X},\{\bar{H}^{k},H^{k}\}_{k\neq i})$.
-By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic kernels weighting the exogenous innovation $I^{i}$.
+By the innovations exogeneity established in Corollary 3.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$.
 [^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
 B.1 The best-response problem in innovations coordinates
-Fix a baseline deterministic-kernel Volterra profile and a deviating player $i$.
+Fix a baseline noise-state linear profile and a deviating player $i$.
 By the innovations exogeneity established in Corollary 3.5, $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$, and $I^{i}$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion invariant to $D^{i}$.
 Player $i$’s control
-$D_{t}^{i}=\bar{D}^{i}(t)+\int_{0}^{t}D^{i}(t,u)\,d_{u}\widehat{W}_{t}^{i}(u)$
+$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$
 is a deterministic affine functional of $I^{i}$, and the optimization is over $(\bar{D}^{i},D^{i})$ with no self-referential signal dependence.
-Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{i}_{DD}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
+Since the state dynamics are linear in $D^{i}$ and the cost is quadratic with $G^{DD,i}(t)\succ 0$, the induced objective is strictly convex: the stationarity condition derived below is the unique global optimum.
 B.2 Spike variation and the first-variation system
@@
 For each $k\neq i$, $\delta(d_{u}\widehat{W}_{s}^{k}(u))=\delta w_{s}^{k}(u)\,du$ with $\delta w_{t}^{k}(\cdot)\equiv 0$ and
-$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}(s,u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X(s,r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
-The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
-$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
+$\partial_{s}\delta w_{s}^{k}(u)=\widetilde{X}^{k}_{s}(u)^{\top}P^{k}(s)\Bigl(\delta X_{s}-\int_{0}^{s}X_{s}(r)\,\delta w_{s}^{k}(r)\,dr\Bigr),\qquad 0\leq u\leq s.$ (B.1)
+The opponent control variation is $\delta D_{s}^{k}=\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du$, giving the coupled state system
+$\frac{d}{ds}\delta X_{s}=A(s)\,\delta X_{s}+\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,\qquad\delta X_{t}=v.$ (B.2)
 B.3 Transition kernels
@@
 These blocks satisfy
-$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,dz,$ (B.3)
-$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}(t,u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}(t,z)^{\top}P^{k}(t)\,X(t,u)\,dz,$ (B.4)
+$\frac{\partial}{\partial t}\Phi^{XX}(s,t)$ $=-\Phi^{XX}(s,t)A(t)-\sum_{k\neq i}\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,dz,$ (B.3)
+$\frac{\partial}{\partial t}\Phi^{Xk}(s,t,u)$ $=-\Phi^{XX}(s,t)D^{k}_{t}(u)+\int_{0}^{t}\Phi^{Xk}(s,t,z)\,\widetilde{X}^{k}_{t}(z)^{\top}P^{k}(t)\,X_{t}(u)\,dz,$ (B.4)
 with $\Phi^{XX}(t,t)=I$ and $\Phi^{Xk}(t,t,u)=0$.
 B.4 First variation of costs and the adjoint kernels
-Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the Volterra form of $X_{s}$, the normalized cost variation is
-$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
+Fix player $i$ and suppress player superscripts on weight matrices. Using $\delta X_{s}=\Phi^{XX}(s,t)v$ and the linear form of $X_{s}$, the normalized cost variation is
+$\delta J_{t}^{i}(v)=2\,\mathbb{E}\!\left[\left.v^{\top}\Bigl(G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)\Bigr)\ \right|\mathcal{F}_{t}^{i}\right],$ (B.5)
 where
-$\bar{H}^{X}(t)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.6)
-$H^{X}(t,u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}G_{XX}(s)\,X(s,u)\,ds+\Phi^{XX}(T,t)^{\top}G^{i}_{XX}(T)\,X(T,u).$ (B.7)
+$\bar{H}^{X}_{t}$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.6)
+$H^{X}_{t}(u)$ $:=\int_{t}^{T}\Phi^{XX}(s,t)^{\top}\bigl[G^{XX}(s)\,X_{s}(u)+G^{X,i}_{s}(u)\bigr]\,ds+\Phi^{XX}(T,t)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(u)+G^{X,i}_{T}(u)\bigr].$ (B.7)
 Stationarity condition (globally necessary and sufficient).
 Setting $\delta J_{t}^{i}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives
-$G_{DD}^{i}(t)D_{t}^{i}+\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
-with equilibrium coefficients $\bar{D}^{i}(t)=-G_{DD}^{i}(t)^{-1}\bar{H}^{X}(t)$ and $D^{i}(t,u)=-G_{DD}^{i}(t)^{-1}H^{X}(t,u)$.
-Equivalently, $D_{t}^{i}=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[\bar{H}^{X}(t)+\int_{0}^{T}H^{X}(t,u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a separation property).
+$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0\qquad\text{a.s.,}$ (B.8)
+with equilibrium coefficients $\bar{D}^{i}_{t}=-G^{DD,i}(t)^{-1}\bar{H}^{X}_{t}$ and $D^{i}_{t}(u)=-G^{DD,i}(t)^{-1}H^{X}_{t}(u)$.
+Equivalently, $D_{t}^{i}=-(G^{DD,i}(t))^{-1}\mathbb{E}[\bar{H}^{X}_{t}+\int_{0}^{T}H^{X}_{t}(u)\,dW_{u}\mid\mathcal{F}_{t}^{i}]$: the control is the conditional expectation of the full-information action (a certainty-equivalence property).
 B.5 Closed backward system for the adjoint kernels
 Define the belief-adjoint objects
-$\bar{H}^{k}(t,u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G_{XX}(s)\bar{X}(s)+G_{X}(s)\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr],$ (B.9)
-$H^{k}(t,u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}G_{XX}(s)\,X(s,r)\,ds+\Phi^{Xk}(T,t,u)^{\top}G^{i}_{XX}(T)\,X(T,r).$ (B.10)
+$\bar{H}^{k}_{t}(u)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\bar{X}_{s}+\bar{G}^{X,i}_{s}\bigr]\,ds+\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr],$ (B.9)
+$H^{k}_{t}(u,r)$ $:=\int_{t}^{T}\Phi^{Xk}(s,t,u)^{\top}\bigl[G^{XX}(s)\,X_{s}(r)+G^{X,i}_{s}(r)\bigr]\,ds$
+$\qquad+\;\Phi^{Xk}(T,t,u)^{\top}\bigl[G^{XX,i}(T)\,X_{T}(r)+G^{X,i}_{T}(r)\bigr].$ (B.10)
 Then $(\bar{H}^{X},\{\bar{H}^{k}\},H^{X},\{H^{k}\})$ satisfy
-$\frac{d}{dt}\bar{H}^{X}(t)$ $=-\bigl[G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr]-A(t)^{\top}\bar{H}^{X}(t)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.11)
-$\frac{d}{dt}\bar{H}^{k}(t,u)$ $=-D^{k}(t,u)^{\top}\bar{H}^{X}(t)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k}(t,z)\,dz,$ (B.12)
-$\frac{d}{dt}H^{X}(t,r)$ $=-G^{i}_{XX}(t)X(t,r)-A(t)^{\top}H^{X}(t,r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.13)
-$\frac{d}{dt}H^{k}(t,u,r)$ $=-D^{k}(t,u)^{\top}H^{X}(t,r)+\int_{0}^{t}X(t,u)^{\top}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,H^{k}(t,z,r)\,dz,$ (B.14)
-with terminal conditions $\bar{H}^{X}(T)=G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)$,  $\bar{H}^{k}(T,\cdot)=0$,  $H^{X}(T,r)=G^{i}_{XX}(T)X(T,r)$,  $H^{k}(T,\cdot,r)=0$.
+$\frac{d}{dt}\bar{H}^{X}_{t}$ $=-\bigl[G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr]-A(t)^{\top}\bar{H}^{X}_{t}-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.11)
+$\frac{d}{dt}\bar{H}^{k}_{t}(u)$ $=-D^{k}_{t}(u)^{\top}\bar{H}^{X}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k}_{t}(z)\,dz,$ (B.12)
+$\frac{d}{dt}H^{X}_{t}(r)$ $=-\bigl[G^{XX,i}(t)X_{t}(r)+G^{X,i}_{t}(r)\bigr]-A(t)^{\top}H^{X}_{t}(r)-\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.13)
+$\frac{d}{dt}H^{k}_{t}(u,r)$ $=-D^{k}_{t}(u)^{\top}H^{X}_{t}(r)+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,H^{k}_{t}(z,r)\,dz,$ (B.14)
+with terminal conditions $\bar{H}^{X}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$,  $\bar{H}^{k}_{T}(\cdot)=0$,  $H^{X}_{T}(r)=G^{XX,i}(T)X_{T}(r)+G^{X,i}_{T}(r)$,  $H^{k}_{T}(\cdot,r)=0$.
 B.6 Risk-sensitive extension: exponential utility FOC
@@
 A Taylor expansion gives $J^{i}_{\theta_{i}}=\mathbb{E}[\mathcal{C}^{i}]+\tfrac{\theta_{i}}{2}\operatorname{Var}(\mathcal{C}^{i})+O(\theta_{i}^{2})$, so $\theta_{i}$ governs the degree of risk aversion (the risk-neutral case $\theta_{i}\to 0$ recovers (2.4)).
-Under a deterministic-kernel Volterra profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
+Under a linear profile, the best-response FOC is identical to Corollary 3.5 with $\widehat{W}_{t}^{i}$ replaced by the risk-adjusted noise-state
 $\widehat{W}^{\,i,\theta}_{t}:=\bigl(I-\theta_{i}C_{t}K_{t}\bigr)^{-1}\!\left(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}\right),$
-where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the Volterra profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16).
+where $C_{t}:=\mathsf{C}_{t}^{i}(u,s)=\operatorname{Cov}(W_{u},W_{s}\mid\mathcal{F}_{t}^{i})$ is the deterministic conditional covariance kernel of $W$ given $\mathcal{F}_{t}^{i}$ (deterministic by Gaussianity of the conditional law under the linear profile), $K_{t}$ is the symmetric quadratic kernel of $\mathcal{C}^{i}$ in the Wiener-chaos representation (B.17), and $k_{t}$ is the corresponding linear kernel (B.16).
 The formula is valid when $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$, which holds for $\theta_{i}$ sufficiently small or $T$ sufficiently short.
-The filtering structure is unchanged; separation fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
+The filtering structure is unchanged; certainty equivalence fails because the quadratic cost kernel rescales conditional precision via $(I-\theta_{i}C_{t}K_{t})^{-1}$ while $k_{t}$ shifts the conditional mean.
 We now derive this result.
@@
 Quadratic structure of the cost.
-Expanding $\mathcal{C}^{i}$ via the Volterra forms $X_{s}=\bar{X}(s)+\int_{0}^{s}X(s,u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}(s)+\int_{0}^{s}\mathcal{D}^{i}(s,u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
+Expanding $\mathcal{C}^{i}$ via the linear forms $X_{s}=\bar{X}_{s}+\int_{0}^{s}X_{s}(u)\,dW_{u}$ and $D^{i}_{s}=\bar{D}^{i}_{s}+\int_{0}^{s}\mathcal{D}^{i}_{s}(u)\,dW_{u}$ and applying stochastic Fubini yields the Wiener-chaos representation
 $\mathcal{C}^{i}=c_{0}+\int_{0}^{T}\ell(r)^{\top}dW_{r}+\tfrac{1}{2}\int_{0}^{T}\!\!\int_{0}^{T}dW_{r}^{\top}K(r,s)\,dW_{s},$
 where $c_{0}$ absorbs deterministic and trace terms, and the linear and symmetric quadratic kernels are
-$\ell(r)$ $:=\int_{r}^{T}\!\Big[X(t,r)^{\top}\bigl(G^{i}_{XX}(t)\bar{X}(t)+G^{i}_{X}(t)\bigr)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\bar{D}^{i}(t)\Big]\,dt$
-$\qquad+\;X(T,r)^{\top}\bigl(G^{i}_{XX}(T)\bar{X}(T)+G^{i}_{X}(T)\bigr),$ (B.16)
-$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X(t,r)^{\top}G^{i}_{XX}(t)\,X(t,s)+\mathcal{D}^{i}(t,r)^{\top}G^{i}_{DD}(t)\,\mathcal{D}^{i}(t,s)\Big]\,dt+X(T,r)^{\top}G^{i}_{XX}(T)\,X(T,s).$ (B.17)
+$\ell(r)$ $:=\int_{r}^{T}\!\Big[X_{t}(r)^{\top}\bigl(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\bigr)+G^{X,i}_{t}(r)^{\top}\bar{X}_{t}$
+$\qquad\qquad+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\bar{D}^{i}_{t}\Big]\,dt$
+$\qquad+\;X_{T}(r)^{\top}\bigl(G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}\bigr)+G^{X,i}_{T}(r)^{\top}\bar{X}_{T},$ (B.16)
+$K(r,s)$ $:=\int_{\max(r,s)}^{T}\!\Big[X_{t}(r)^{\top}G^{XX,i}(t)\,X_{t}(s)+X_{t}(r)^{\top}G^{X,i}_{t}(s)$
+$\qquad\qquad+G^{X,i}_{t}(r)^{\top}X_{t}(s)+\mathcal{D}^{i}_{t}(r)^{\top}G^{DD,i}(t)\,\mathcal{D}^{i}_{t}(s)\Big]\,dt$
+$\qquad+\;X_{T}(r)^{\top}G^{XX,i}(T)\,X_{T}(s)+X_{T}(r)^{\top}G^{X,i}_{T}(s)+G^{X,i}_{T}(r)^{\top}X_{T}(s).$ (B.17)
 Tilted conditional FOC.
@@
 Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ and applying the tower property gives
-$G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
+$G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}\!\left[\mathcal{M}^{i}_{t}\,\Big|\,\mathcal{F}_{t}^{i}\right]=0,\qquad\mathbb{E}^{Q^{i}}[\,\cdot\mid\mathcal{F}_{t}^{i}]:=\frac{\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]}{\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]},$
 which has the structure of (B.8) under the tilted measure.
@@
 $C_{t,\theta}^{-1}=C_{t}^{-1}-\theta_{i}K_{t},\qquad m_{t,\theta}=C_{t,\theta}(C_{t}^{-1}m_{t}+\theta_{i}k_{t}),$
-provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short). Rearranging gives
+provided $I-\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\succ 0$ (satisfied for $|\theta_{i}|$ small or $T$ short).
+Here $C_{t}$ and $K_{t}$ act as integral operators on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$, so $(I-\theta_{i}C_{t}K_{t})^{-1}$ is the resolvent (Neumann series) $(I-\theta_{i}C_{t}K_{t})^{-1}f=f+\theta_{i}C_{t}K_{t}f+\theta_{i}^{2}(C_{t}K_{t})^{2}f+\cdots$, convergent precisely when $\|\theta_{i}C_{t}^{1/2}K_{t}C_{t}^{1/2}\|_{\mathrm{op}}<1$.
+Rearranging gives
 $\widehat{W}^{\,i,\theta}_{t}=(I-\theta_{i}C_{t}K_{t})^{-1}(\widehat{W}_{t}^{i}+\theta_{i}C_{t}k_{t}),$
 and substituting into the tilted FOC yields
-$\boxed{G^{i}_{DD}(t)D^{i}_{t}\;+\;\bar{H}^{X}(t)\;+\;\int_{0}^{T}H^{X}(t,r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
+$\boxed{G^{DD,i}(t)D^{i}_{t}\;+\;\bar{H}^{X}_{t}\;+\;\int_{0}^{T}H^{X}_{t}(r)\,d_{r}\,\widehat{W}^{\,i,\theta}_{t}(r)\;=\;0,}$ (B.18)
 where
@@
 The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$.
-Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{i}_{DD}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure.
-Since $\mathcal{C}^{i}$ is quadratic in $W$ under the Volterra profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
+Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure.
+Since $\mathcal{C}^{i}$ is quadratic in $W$ under the linear profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
 B.7 Information cost: proof of Proposition 3.6
@@
 At the equilibrium profile $D^{*}$, the stochastic costate is
-$M^{i}_{t}:=\bar{H}^{i}_{X}(t)+\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}W_{u},$
-where $(\bar{H}^{i}_{X},H^{i}_{X})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
-$D^{i,\mathrm{full}}_{t}$ $=-(G^{i}_{DD}(t))^{-1}M^{i}_{t},$ (B.20)
-$D^{i}_{t}$ $=-(G^{i}_{DD}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
+$M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u},$
+where $(\bar{H}^{X,i},H^{X,i})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
+$D^{i,\mathrm{full}}_{t}$ $=-(G^{DD,i}(t))^{-1}M^{i}_{t},$ (B.20)
+$D^{i}_{t}$ $=-(G^{DD,i}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
 Both are evaluated at the same adjoint kernels: opponents’ maps are fixed and linear with deterministic kernels, so the backward system is identical regardless of player $i$’s information.
 The control gap is
-$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{i}_{DD}(t))^{-1}\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
+$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
 where $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$.
@@
 Since $D^{i,\mathrm{full}}$ minimizes the strictly convex quadratic $D^{i}\mapsto J^{i}(D^{i};D^{*,-i})$ over $L^{2}(\mathcal{F})$ (Lemma 3.4), and $\delta D^{i}\in L^{2}(\mathcal{F})$, the first variation vanishes and
-$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{i}_{DD}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
-To verify that only $G^{i}_{DD}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$.
+$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{DD,i}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
+To verify that only $G^{DD,i}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$.
 The zero-order term vanishes by definition.
-All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{i}_{X})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$.
-The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{i}_{DD}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal).
-The standard LQG completion-of-the-square identifies these: using $G^{i}_{DD}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
+All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{X,i})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$.
+The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{DD,i}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal).
+The standard LQG completion-of-the-square identifies these: using $G^{DD,i}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
 Kernel formula for $\Sigma^{M,i}_{t}$.
 Substituting the control gap into (B.22):
-$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{i}_{DD}(t))^{-1}\left(\int_{0}^{t}H^{i}_{X}(t,v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
-Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{i}_{DD}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.9).
+$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{DD,i}(t))^{-1}\left(\int_{0}^{t}H^{X,i}_{t}(v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
+Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{DD,i}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.7).
 It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
 The conditional mean of $M^{i}_{t}$ given $\mathcal{F}^{i}_{t}$ is
-$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{i}_{X}(t)+\int_{0}^{t}\mathcal{H}^{i}(t,s)\,\mathrm{d}W_{s},$
-where the filtered adjoint kernel is
-$\mathcal{H}^{i}(t,s):=H^{i}_{X}(t,s)\,\Pi^{i}+\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
-This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s})\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{i}_{X}(t,u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini.
+$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}\widehat{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
+where the resolved adjoint kernel is
+$\widehat{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)\,\Pi^{i}+\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
+This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+\bigl(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s}\bigr)\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini.
 Then
-$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\bigl(H^{i}_{X}(t,s)-\mathcal{H}^{i}(t,s)\bigr)\,\mathrm{d}W_{s}=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s},$
-where $\mathcal{U}^{i}(t,s):=H^{i}_{X}(t,s)(I-\Pi^{i})-\int_{s}^{t}H^{i}_{X}(t,u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$. By Itô isometry,
-$\Sigma^{M,i}_{t}=\mathbb{E}\!\left[\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)\left(\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathrm{d}W_{s}\right)^{\top}\right]=\int_{0}^{t}\mathcal{U}^{i}(t,s)\,\mathcal{U}^{i}(t,s)^{\top}\,\mathrm{d}s.$
+$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
+where the unresolved adjoint kernel $\widetilde{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)-\widehat{H}^{X,i}_{t}(s)=H^{X,i}_{t}(s)(I-\Pi^{i})-\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$ is the analogue of $\widetilde{X}^{i}$ for the backward system.
+By Itô isometry,
+$\Sigma^{M,i}_{t}=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\widetilde{H}^{X,i}_{t}(s)^{\top}\,\mathrm{d}s.$
 Separation from control.
 The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$.
-The adjoint $H^{i}_{X}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself.
-Both $\mathcal{U}^{i}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof.
+The adjoint $H^{X,i}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself.
+Both $\widetilde{H}^{i}_{X}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof.
 ∎
-Appendix C Well-posedness of the deterministic kernel fixed point
-This appendix formulates the equilibrium as a dynamic programming problem on the kernel state space, records the a priori bounds that follow from the projection structure of the filtering kernel, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
+Appendix C Well-posedness of the deterministic impulse-response fixed point
+This appendix formulates the equilibrium as a dynamic programming problem on the impulse-response state space, records the a priori bounds that follow from the projection structure of the filtering map, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
 Notation.
@@
 Assumption C.1.
-There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{i}_{XX}(t)\|+\|G^{i}_{X}(t)\|+\|G^{i}_{XX}(T)\|+\|G^{i}_{X}(T)\|\leq L$ and $G^{i}_{DD}(t)\succeq\lambda\,I_{d}$.
+There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{XX,i}(t)\|+\|\bar{G}^{X,i}_{t}\|+\|G^{X,i}(t,\cdot)\|_{L^{\infty}}+\|G^{XX,i}(T)\|+\|\bar{G}^{X,i}_{T}\|+\|G^{X,i}(T,\cdot)\|_{L^{\infty}}\leq L$ and $G^{DD,i}(t)\succeq\lambda\,I_{d}$.
 C.2 The kernel state space and its dynamics
 At time $t$ the kernel state is
-$\mathcal{Z}_{t}:=\bigl(\bar{X}(t),\;X(t,\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
+$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
 The unresolved kernel $\widetilde{X}^{j}$ is not a state variable; it is determined algebraically from $(\mathcal{Z}_{t})$ by
-$\widetilde{X}^{j}(t,u)=X(t,u)(I-\Pi^{j})-\int_{0}^{t}X(t,z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
-Under a noise-state Volterra profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
-$\dot{\bar{X}}(t)$ $=A(t)\bar{X}(t)+\textstyle\sum_{k}\bar{D}^{k}(t),$ (C.3)
-$\dot{X}(t,s)$ $=A(t)X(t,s)+\textstyle\sum_{k}\mathcal{D}^{k}(t,s),\qquad X(s,s)=\Sigma(s),$ (C.4)
-$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}(t,u)^{\top}P^{j}(t)\,\widetilde{X}^{j}(t,s),\qquad F_{0}^{j}=0,$ (C.5)
-with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}(t,u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}(t,s):=D^{k}(t,s)\Pi^{k}+\int_{0}^{t}D^{k}(t,u)\,F_{t}^{k}(u,s)\,du$.
+$\widetilde{X}^{j}_{t}(u)=X_{t}(u)(I-\Pi^{j})-\int_{0}^{t}X_{t}(z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
+Under a noise-state linear profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
+$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (C.3)
+$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (C.4)
+$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (C.5)
+with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
 The policy $\mathbf{D}$ enters (C.3)–(C.4) but does not enter (C.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (C.2).
@@
 where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing, $\ell^{i}$ is the running cost, and $g^{i}$ is the terminal cost, both evaluated at the kernel state.
-Player $i$’s control $(\bar{D}^{i}(t),D^{i}(t,\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
-$\bar{D}^{i}(t)=-(G^{i}_{DD}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}(t,u)=-(G^{i}_{DD}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
-Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}_{X}^{i}(t)$ and $D_{X}V^{i}(\cdot)=H_{X}^{i}(t,\cdot)$, recovering Corollary 3.5.
+Player $i$’s control $(\bar{D}^{i}_{t},D^{i}_{t}(\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
+$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
+Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}^{X,i}_{t}$ and $D_{X}V^{i}(\cdot)=H^{X,i}_{t}(\cdot)$, recovering Corollary 3.5.
 The information wedge as a shadow price.
@@
 Along the equilibrium path, differentiating through the chain $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies
-$\bar{\mathcal{V}}^{i}(t)=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}(t,z)\,\bar{H}^{k,i}(t,z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
+$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
 i.e.
 the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel.
-The kernel wedge $\mathcal{V}^{i}(t,r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
+The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
 Endogenous precision.
 If $P^{i}(\cdot)$ is player $i$’s control subject to attention cost $\int_{0}^{T}c^{i}(t,P^{i}(t))\,dt$, then $P^{i}$ enters (C.6) only through $\dot{F}^{i}$ via (C.5), and the FOC
-$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}(t,\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}(t,\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
+$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}_{t}(\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}_{t}(\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
 decouples from the FOC for $D^{i}$: the left side is the marginal value of accelerating the filter, which vanishes as $\mathcal{P}_{t}^{i}\to I$ (Lemma C.3), guaranteeing an interior solution when $\partial c^{i}/\partial P^{i}>0$.
@@
 The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$.
-This prevents finite-time blowup despite the quadratic nonlinearity and is the Volterra-coordinate manifestation of Kalman variance collapse.
+This prevents finite-time blowup despite the quadratic nonlinearity and is the primitive noise-linear manifestation of Kalman variance collapse.
 C.5 Well-posedness
@@
 Definition C.2 (Best-response operator).
-$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{i}_{DD})^{-1}\bar{H}_{X}^{i,\mathbf{D}},\;-(G^{i}_{DD})^{-1}H_{X}^{i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$.
+$\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{DD,i})^{-1}\bar{H}^{X,i,\mathbf{D}},\;-(G^{DD,i})^{-1}H^{X,i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 3.2 at the forward environment induced by $\mathbf{D}$.
 Fixed points are Nash equilibria (Corollary 3.5, Remark 3.2).
@@
 (i) The forward system has a unique global solution with $\|\bar{X}\|_{C}+\|X\|_{L^{\infty}}\leq K_{\mathrm{fwd}}$, $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}}\leq K_{\mathrm{fwd}}\sqrt{T}$, $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$.
-(ii) The backward system has a unique solution with $\|\bar{H}_{X}^{i}\|_{C}+\|H_{X}^{i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
-(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}_{X}^{i},H_{X}^{i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
+(ii) The backward system has a unique solution with $\|\bar{H}^{X,i}\|_{C}+\|H^{X,i}\|_{L^{\infty}}\leq K_{\mathrm{bwd}}$, and time derivatives bounded by $C(M)K_{\mathrm{bwd}}$.
+(iii) Both maps $\mathbf{D}\mapsto(\bar{X},X,\widetilde{X}^{j})$ and $\mathbf{D}\mapsto(\bar{H}^{X,i},H^{X,i})$ are Lipschitz with constants $C_{\mathrm{fwd}}(T,M)$ and $C_{\mathrm{bwd}}(T,M)$ respectively, both bounded by $C(M)\,T\,e^{C(M)T}$.
 Here $K_{\mathrm{fwd}},K_{\mathrm{bwd}}\leq C(M)e^{C(M)T}$.
 Proof sketch.
-For the forward system: $(\bar{X},X)$ solve linear Volterra equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled.
-Since $\mathcal{D}^{k}$ involves $F^{k}$ through a Volterra integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
+For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled.
+Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
 For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (C.2).
@@
 Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$.
-In particular, there exists a unique Nash equilibrium in noise-state Volterra strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
+In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
 Proof.
@@
 $\tau$-closedness of $\mathbb{B}_{T}^{M_{0}}$: weak-* lower semicontinuity of the $L^{\infty}$ norm preserves the bound under $L^{2}$ limits.
-$\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward Volterra equations (via Cauchy–Schwarz) and the linear backward system.
-$\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}_{X}^{i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]).
+$\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system.
+$\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]).
 ∎
@@
 D.1 Drift kernels
-The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)$; for trader $k$:
-$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma}(t,s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
+The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)$; for trader $k$:
+$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
 The unresolved kernels $\widetilde{C}^{0}$ and $\widetilde{C}^{k}$ are defined by the projection decomposition (A.12) with $C^{j}$ replacing $\sqrt{P^{j}}X$ and $\Pi_{\mathcal{E}}^{j}$ replacing $\Pi^{j}$.
@@
 Section B.2) is:
-$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}(s,u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
+$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
 $\partial_{s}\delta w_{s}^{0}(u)$ $=\widetilde{C}^{0}(s,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,s}-\int_{0}^{s}C^{0}(s,r)\,\delta w_{s}^{0}(r)\,dr\biggr),$ (D.3)
@@
 The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$).
-The conditional future demand in Volterra form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}(\tau,u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes).
+The conditional future demand in noise-state linear form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}_{\tau}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes).
 Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3).
-$D^{i}(t,u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$.
-To pin down $D^{i}(t,u)$, we differentiate both sides in $t$.
+$D^{i}_{t}(u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$.
+To pin down $D^{i}_{t}(u)$, we differentiate both sides in $t$.
 D.4 Leibniz differentiation and backward ODE
@@
 Leibniz on $H^{i,\mathrm{imp}}$ gives
-$\frac{d}{dt}H^{i,\mathrm{imp}}(t,u)=-\lambda(t)\,D^{i}(t,u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}(\tau,u)\,d\tau,$ (D.7)
-where the boundary term $\delta P_{t}\cdot D^{i}(t,u)=-\lambda(t)D^{i}(t,u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{i}_{DD}$ and where $D^{i}(t,u)$ first appears explicitly.
+$\frac{d}{dt}H^{i,\mathrm{imp}}_{t}(u)=-\lambda(t)\,D^{i}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (D.7)
+where the boundary term $\delta P_{t}\cdot D^{i}_{t}(u)=-\lambda(t)D^{i}_{t}(u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{DD,i}$ and where $D^{i}_{t}(u)$ first appears explicitly.
 Value adjoint.
-In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}(t,u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$.
+In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}_{t}(u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$.
 Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
-$\frac{d}{dt}H^{i,\mathrm{val}}(t,u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
+$\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
 Closed backward ODE.
-Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}(t,u)$ yields (5.4).
+Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}_{t}(u)$ yields (5.4).
 The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$.

### 23 Mar 2026 v3

@@
 Abstract
-In strategic environments with private information, evaluating a change in policy requires predicting how the equilibrium responds—but when actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs [28] that has resisted exact analysis for four decades.
-We provide the first exact equilibrium characterization of finite-player continuous-time linear quadratic games with endogenous signals.
-Conditioning on primitive Brownian shocks rather than the physical state—a dynamic analogue of Harsanyi’s common-prior construction—collapses the belief hierarchy onto deterministic impulse response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit.
-The characterization yields an explicit information wedge that prices the marginal value of shifting opponents’ posteriors.
-The wedge vanishes precisely when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters, and provides a closed-form mapping from information primitives to equilibrium outcomes.
+When actions reshape opponents’ signals, each agent’s optimal response depends on an infinite hierarchy of beliefs about beliefs (Townsend, 1983) that has resisted exact analysis for four decades.
+We provide the first exact equilibrium characterization of finite-player continuous-time linear–quadratic games with endogenous signals.
+Estimating primitive shocks rather than the state collapses the belief hierarchy onto deterministic impulse-response maps, reducing Nash equilibrium to a deterministic fixed point with no truncation and no large-population limit.
+The characterization yields an information wedge pricing the marginal value of shifting opponents’ posteriors; it vanishes precisely when signals are exogenous.
+In a two-player benchmark, nearly all welfare gains from pooling information come from eliminating bilateral belief manipulation rather than from improving state estimation.
+This strategic channel also governs optimal precision allocation: a planner should concentrate information on the efficient player, since arming the inefficient player maximizes the arms race while starving them of information collapses it.
 1 Introduction
-In strategic environments with private information, evaluating a change in policy—a disclosure mandate, a shift in the precision of public communication, a redesign of market transparency—requires understanding how the equilibrium responds, not merely how agents behave under existing rules.
-When a regulator alters the information structure, each agent’s optimal action depends on their beliefs about the state and on their beliefs about how others are filtering, and both change with the policy.
-The filtering environment and the strategic equilibrium are jointly determined.
-This joint determination is the central obstacle.
-Consider a central bank shifting from quarterly to real-time disclosure of its output gap estimates.
-Each firm’s pricing decision depends on its belief about the gap and on its belief about how other firms filter the bank’s announcements.
-Under the new policy, signal precision changes; firms filter differently, revise their forecasts of one another’s pricing, and equilibrium inflation dynamics shift in tandem.
-Existing tools cannot trace this chain because the Townsend hierarchy makes the mapping from signal precision to equilibrium outcomes intractable.
-The characterization in this paper makes it explicit.
-The same obstruction arises across a wide range of fields.
-In market microstructure, prices aggregate dispersed information and trading itself is informative.
-Dynamic information design faces the same difficulty: the mapping from signal structures to equilibrium outcomes becomes intractable once filtering and equilibrium interact.
-Contract theory encounters it when multiple agents exert unobservable effort on a shared outcome and draw inferences from it.
-So does decentralized control of autonomous vehicle fleets and multi-robot systems with misaligned operators, wherever agents observe a shared state through private channels and actions reshape what others sense.
-Existing approaches to the decentralized control problem typically assume cooperative teams or a common information base; the framework here requires neither.
-Townsend [28] crystallized the difficulty: in a game with private signals, each agent must forecast others’ forecasts of the state, generating an infinite hierarchy of conditional expectations.
-Sargent [26] showed that even in linear–Gaussian environments the hierarchy prevents finite-dimensional recursive analysis, because each agent’s action enters the state dynamics, entangling every agent’s past actions, beliefs, and signals in a way the Kalman filter cannot close.
-The conceptual move parallels Harsanyi’s [10] resolution of the belief hierarchy in incomplete-information games.
-Harsanyi replaced an infinite regress of beliefs about beliefs with a common prior over a type space from which all higher-order beliefs follow by conditioning.
-We replace the Townsend regress with the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$, from which all higher-order conditional expectations follow by linear projection onto deterministic impulse response maps with no truncation.
-The analogy is imperfect in a revealing way: Harsanyi’s types are drawn once and the information partition is fixed, whereas here noise unfolds continuously and actions reshape signals, so the closure requires not just a common prior but a fixed point—the maps governing how projections evolve must be consistent with the strategy profile they induce (Figure 1).
-In the single-agent problem the conditional mean of the state is a sufficient statistic, but the state estimate is a linear functional of the noise estimate, not the reverse.
-In the multi-agent setting the state estimate fails to close the equilibrium because it entangles exogenous shocks with endogenous responses—precisely the Townsend hierarchy.
-Estimating the noise path avoids this entanglement.
-We call the path $\widehat{W}_{t}^{i}(\cdot):=\mathbb{E}[W_{\cdot}\mid\mathcal{F}_{t}^{i}]$ the player’s noise-state; optimal controls are affine functionals of it, just as single-agent controls are affine in the state estimate.
-The main results establish that the best response amongst all admissible actions to any profile of linear strategies is itself linear.
-The equilibrium characterization produces an explicit information wedge $\mathcal{V}^{i}_{t}$, which prices the marginal value to player $i$ of shifting opponents’ posteriors.
-It vanishes identically when signals are exogenous to controls, formally delineating the boundary where strategic belief manipulation matters.
-The propagation $\Delta P^{i}\to\Delta\widetilde{X}^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta D^{i}$ provides an explicit mapping from information primitives to equilibrium outcomes that a model with fixed filtering cannot deliver: because information enters through precision paths and affects the equilibrium only through deterministic impulse response maps, a principal can trace how disclosure precision propagates into equilibrium payoffs without simulating higher-order beliefs.
-(i) Candidate Opponents use Noise-State Linear controls (ii) Closure State & filters have det.
-response maps (iii) Deviations Fixed maps $\Rightarrow$ det.
-propagation (iv) Best response FOC $\Rightarrow$ control noise-state linear (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
-Figure 1: The Noise-State Linear fixed-point loop.
-(i) Assume opponents use Noise-state Linear controls.
-(ii) The state and filtering equations have deterministic impulse responses.
-(iii) Under fixed strategy maps, unobserved deviations propagate deterministically.
-(iv) The maximum principle produces a best response that is itself noise-state linear, closing the loop.
-(v) A fixed point of this best-response map is a Nash equilibrium.
+Dynamic games with dispersed private information pervade economics: firms infer competitors’ pricing from market outcomes, traders learn from the order flow their own trades generate, central banks know their announcements reshape private agents’ signals.
+In each setting, actions feed back into the information environment, generating an infinite hierarchy of beliefs that Townsend [32] identified and Sargent [29] showed resists finite-dimensional analysis even in linear–Gaussian environments.
+Bergemann and Morris [6] identify dynamic decentralized information design as a central open problem.
+Four decades after Townsend, this paper resolves the hierarchy: estimating primitive shocks rather than the state collapses the infinite regress onto deterministic impulse-response maps, reducing Nash equilibrium to a deterministic fixed point.
+No truncation or large-population limit is needed: a researcher who specifies payoffs, a state equation, and a signal structure gets back equilibrium impulse-response functions; the belief hierarchy is handled internally.
+The conceptual move parallels Harsanyi’s [12] resolution of the belief hierarchy in incomplete-information games: just as a common prior over a type space closes the static regress, the common probability space of primitive shocks closes the dynamic one, with all higher-order conditional expectations following by linear projection onto deterministic impulse-response maps.
+The difference is that Harsanyi’s types are drawn once while noise unfolds continuously and actions reshape signals, so the closure requires a fixed point: the maps governing belief updating must be consistent with the strategy profile they induce (Section 4).
+The resolution produces a new object—the information wedge $\mathcal{V}^{i}_{t}$—that prices the marginal value of shifting opponents’ posteriors.
+If player $i$ increases effort by a small amount, the state drifts, opponent $k$’s signal changes, $k$ revises their posterior and adjusts their action, and the wedge is the present value of that entire chain.
+It vanishes precisely when signals do not depend on actions, delineating the boundary where strategic belief manipulation matters.
+The wedge organizes four findings, ordered by their economic content.
+Welfare decomposition.
+In a two-player game with opposing targets, the simplest setting exhibiting the full bilateral externality, nearly all welfare gains from pooling information come from eliminating the channel by which agents manipulate each other’s posteriors, not from improving state estimation (Section 2).
+The estimation benefit of pooling is identical whether agents cooperate or compete; the order-of-magnitude difference isolates the strategic externality.
+Information starvation.
+Allowing asymmetric effort costs, optimal precision allocation under competition is sharply asymmetric and in the opposite direction from what strategic intuition suggests.
+One might expect a planner to balance precision to limit manipulation—concentrating information arms the efficient player, so strategic restraint should push toward equality.
+Instead, concentrating information on the efficient player collapses the arms race, while shifting it toward the inefficient player maximizes bilateral waste (Section 2.5).
+Separation failure.
+Certainty equivalence fails: with multiple agents and endogenous signals, improving one player’s information changes their policy rule itself, not merely their state estimate, because the optimal action depends on what opponents have and have not learned.
+In any policy evaluation exercise—central bank transparency, disclosure regulation, market structure reform—the information structure cannot be treated as a parameter that shifts beliefs while leaving decision rules fixed; the wedge couples the two, and the equilibrium must be recomputed.
+The obstruction arises wherever actions enter the state and signals are not fully revealing—the regime Grossman and Stiglitz [10] showed is generic when information acquisition is costly.
+Exogenous-signal reduction.
+Any model that treats signals as exogenous, including mean-field limits, frequency-domain solutions with fixed spectral densities, and team-theoretic benchmarks, is implicitly assuming away the strategic channel that drives the first three findings.
+The wedge makes this precise: when controls do not enter the state dynamics, $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state (Corollary 4.3).
+Section 2 previews the economic consequences in a two-player game; Section 4 develops the general theory.
 1.1 Related Literature
-The joint determination of filtering and equilibrium has been approached from several directions, each achieving tractability by restricting or removing the informational externality that makes the problem hard.
+The joint determination of filtering and equilibrium has been approached from several directions; each achieves tractability by restricting or removing the informational externality that the present paper resolves exactly.
 Higher-order expectations and the social value of information.
 The infinite regress of “forecasting the forecasts of others” was identified by Townsend [32] and shown by Sargent [29] to resist finite-dimensional recursive analysis even in linear–Gaussian environments.
-Morris and Shin [20] demonstrated that the social value of public information can be negative when agents overweight common signals—a static manifestation of the belief-manipulation channel that arises dynamically through the information wedge $\mathcal{V}^{i}_{t}$.
-Our Proposition 3.6 provides a dynamic counterpart: Section 4 shows that nearly all welfare gains from pooling come from eliminating the strategic channel rather than improving state estimation, consistent with the Morris–Shin intuition that the coordination externality, not statistical efficiency, drives welfare losses.
+Morris and Shin [23] demonstrated that the social value of public information can be negative when agents overweight common signals; Angeletos and Pavan [1] characterized the efficient use of information in static coordination games, showing that welfare losses depend on the gap between equilibrium and socially optimal degrees of coordination.
+Woodford [34] showed that imperfect common knowledge generates persistent real effects of monetary policy through higher-order expectations, with the infinite hierarchy truncated by assuming exogenous private signals.
+Section 2 provides a dynamic counterpart: the information wedge $\mathcal{V}^{i}_{t}$ prices the strategic channel through which actions reshape beliefs, and nearly all welfare gains from pooling come from eliminating this channel rather than from improving state estimation.
 Frequency-domain methods and exogenous-signal models.
@@
 These methods succeed because the filtering environment is invariant to changes in control laws—the spectral factorization can be solved independently of the policy.
-When controls enter the state dynamics the transfer function depends on the policy, breaking this separation; the present paper addresses this complementary case.
+When controls enter the state dynamics, the signal process depends on the policy, the endogenous-information case that Huo and Takayama [14] show cannot admit a finite-state representation in general.
+The noise-state technique resolves this case exactly: the belief hierarchy collapses onto a deterministic fixed point with no finite-state approximation required.
 Common-information approaches.
-A powerful technique in decentralized control conditions on a common information base shared by all agents to obtain a recursive reduction [21].
+The common-information approach conditions on a shared information base to obtain a recursive reduction [24].
 The approach handles asymmetric information but resolves it through a coordinator’s beliefs; the channel by which player $i$’s action shifts player $k$’s posterior, which shifts $k$’s action, which shifts $i$’s signal, is not captured.
 Team theory.
-The decentralized decision problem with shared objectives originates with Marschak and Radner [19]; Radner [23] established certainty equivalence and linear optimal policies for linear quadratic gaussian (LQG) teams.
-Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure of Section 3 never arises.
-The information wedge $\mathcal{V}^{i}_{t}$ is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
+The decentralized decision problem with shared objectives originates with Marschak and Radner [22]; Radner [26] established certainty equivalence and linear optimal policies for Linear–Quadratic–Gaussian (LQG) teams.
+Team-theoretic solutions bypass the belief hierarchy because a social planner with aligned objectives solves a single-agent control problem—the fixed-point structure never arises.
+The information wedge $\mathcal{V}^{i}_{t}$ displayed in Section 2 is an equilibrium object: it prices the value of shifting an opponent’s posterior, which requires distinct objectives to be well-defined.
 Mean-field LQG games.
@@
 Carmona and Delarue [8] provide a comprehensive two-volume treatment including LQG specifications with common noise.
-The bilateral informational externality is absent by construction in the mean-field limit; $\mathcal{V}^{i}_{t}$ is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
+The bilateral informational externality is absent by construction in the mean-field limit; the information wedge $\mathcal{V}^{i}_{t}$ of Section 2 is a finite-$n$ object that vanishes in any limit where individual actions become negligible.
 Truncation of the belief hierarchy.
@@
 Applied continuous-time filtering games.
-Kyle [15], Back [2], and Foster and Viswanathan [8] solve strategic trading problems with private information, each exploiting special structure—a single informed trader, correlated-signal symmetry—to close the belief hierarchy through model-specific conjectures.
+Kyle [18], Back [3], and Foster and Viswanathan [9] solve strategic trading problems with private information, each exploiting special structure (a single informed trader, correlated-signal symmetry) to close the belief hierarchy through model-specific conjectures.
 Sannikov [28] solves a continuous-time principal–agent problem by showing that the agent’s continuation value serves as a sufficient statistic for the principal’s problem, yielding a recursive characterization despite the agent’s private action.
-The technique relies on the bilateral structure: the principal’s belief about a single agent’s hidden action can be summarized by a one-dimensional state variable, and the agent has no beliefs about other agents to forecast.
-With multiple agents exerting unobservable effort on a shared outcome—the moral-hazard analogue of our setting—each agent draws inferences from the shared output about others’ effort, generating the same belief hierarchy that arises here.
-The noise-state framework extends naturally to this setting: each agent’s hidden effort enters the output dynamics exactly as controls enter the state in (2.1), and the information wedge prices the value of shifting the principal’s and co-agents’ posteriors about one’s effort.
+The technique relies on the bilateral structure: the principal observes output directly, so the agent’s deviation changes her payoff but not her posterior, and the wedge $\mathcal{V}^{i}_{t}$ is absent.
+With multiple agents exerting unobservable effort on a shared outcome, each agent infers others’ effort from the shared output, the noise-state path replaces the scalar continuation value, and the belief hierarchy reappears.
 The present paper provides a general framework that nests these as special cases: in the Kyle–Back embedding (Section 5), the information wedge corresponds to the strategic component of price impact, and the characterization extends to multi-trader settings where guess-and-verify breaks down.
+Rational inattention.
+Sims [31] introduced rational inattention; Maćkowiak and Wiederholt [21] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [20] for a survey).
+These models take the strategic environment as given when solving the attention problem; the first-order conditions derived here (Section D.3) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.
 Information design.
 Kamenica and Gentzkow [15] established Bayesian persuasion; Bergemann and Morris [5, 6] introduced Bayes correlated equilibrium and survey the field, identifying dynamic decentralized settings as a central open problem.
-Bergemann, Heumann, and Morris [5] provide precision comparative statics in static coordination games analogous to the propagation $\Delta P^{i}\to\Delta\mathcal{V}^{i}_{t}\to\Delta J^{i}$ derived here.
-The difficulty in extending these results dynamically is that filtering and equilibrium interact; our characterization provides the equilibrium mapping that dynamic information design requires as an input.
-Rational inattention.
-Sims [27] introduced rational inattention; Maćkowiak and Wiederholt [17] showed that firms optimally attend to idiosyncratic over aggregate conditions, generating sluggish price adjustment (see Maćkowiak, Mat́ejka, and Wiederholt [18] for a survey).
-These models take the strategic environment as given when solving the attention problem; Corollary 3.7 and the FOC (C.9) extend the framework to settings where the marginal value of precision depends on the equilibrium through the information wedge.
+Vives [33] develops the interplay between information acquisition and competition in financial markets that the Kyle–Back embedding of Section 5 operationalizes dynamically.
+Bergemann, Heumann, and Morris [4] provide precision comparative statics in static coordination games where the effect of signal precision on payoffs admits direct analysis; dynamically, the information wedge couples the value of precision to the equilibrium, so that filtering and control no longer separate.
+The equilibrium characterization derived here (Section 4) provides the mapping from signal structures to equilibrium outcomes that dynamic information design requires as an input.
+2 Bilateral belief manipulation in a two-player game
+Before developing the general framework, we display its consequences in the simplest game exhibiting the full bilateral externality: two symmetric players with scalar state, opposing targets, and equal signal precision.
+Neither player observes the state directly; each receives a private diffusion signal whose precision $p>0$ is common knowledge.
+Each player’s control enters the state dynamics and reshapes the opponent’s signal, generating the bilateral informational externality of Townsend [32].
+The general theory (Sections 34) collapses this hierarchy onto deterministic impulse-response maps and produces an object, the information wedge $\mathcal{V}^{i}_{t}$, that prices the marginal value of shifting the opponent’s posterior.
+2.1 Primitives and objectives
+Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
+$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (2.1)
+where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the private diffusion channel
+$dY_{t}^{1}=\sqrt{p}\,X_{t}\,dt+dW_{t}^{1},\qquad dY_{t}^{2}=\sqrt{p}\,X_{t}\,dt+dW_{t}^{2},$ (2.2)
+with $(W^{0},W^{1},W^{2})$ mutually independent. Player $i$ chooses $D^{i}$ adapted to their private observation history $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
+Running costs are symmetric tracking losses with quadratic effort penalty $r>0$:
+$J^{1}(D^{1};D^{2})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}-1)^{2}+r\,(D_{t}^{1})^{2}\Big)\,dt,\quad J^{2}(D^{2};D^{1})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}+1)^{2}+r\,(D_{t}^{2})^{2}\Big)\,dt.$ (2.3)
+Player 1 wants the state near +1+1; player 2 wants it near 1-1.
+There is no terminal cost in this worked example.
+The model is invariant under the sign/label symmetry $(X,Y^{1},Y^{2},D^{1},D^{2})\mapsto(-X,Y^{2},Y^{1},-D^{2},-D^{1})$.
+[…]
+Figure 1: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision pp.
+2.2 What the general theory delivers
+The equilibrium characterization developed in Sections 34 produces three objects that organize the analysis of this game; we describe them informally here and derive them in full generality below.
+Noise-state and linear strategies.
+All randomness originates in the three primitive Brownian motions $W=(W^{0},W^{1},W^{2})$.
+Player $i$’s noise-state $\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}]$ ($0\leq u\leq t$) is the running conditional mean of the entire primitive noise path.
+In equilibrium, each player’s control takes the form
+$D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),$ (2.4)
+where $\bar{D}^{i}_{t}$ is a deterministic mean action and $D^{i}_{t}(u)$ is a deterministic impulse-response kernel governing how player $i$’s action at time $t$ responds to news about the $u$-th shock.
+The key structural result (Corollary 4.5) is that the best response to any such strategy profile, over all admissible controls, is itself of this form—so equilibrium reduces to a deterministic fixed point in the maps $(\bar{D}^{i},D^{i})$.
+Unresolved impulse response.
+The state’s impulse response to the $u$-th shock decomposes into what player $i$ has learned and what remains hidden: $X_{t}(u)=\widehat{X}^{i}_{t}(u)+\widetilde{X}^{i}_{t}(u)$.
+The unresolved component $\widetilde{X}^{i}_{t}(u)$ is the portion of the $u$-th shock’s effect on the state that player $i$ has not yet resolved—large when the shock is recent or precision low, shrinking as observations accumulate.
+It serves as the Kalman gain in the noise-state update (Theorem 4.1).
+Information wedge.
+In the single-agent problem, the backward equation for the costate (the shadow value of the state) is standard. With two agents and endogenous signals, a new term appears: the information wedge
+$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$ (2.5)
+where $\widetilde{X}^{2}_{t}(z)$ is the unresolved component of the state’s response to the $z$-th shock from player 2’s perspective and $\bar{H}^{2}_{t}(z)$ is the mean belief adjoint—the shadow value to player 1 of having shifted player 2’s estimate of the $z$-th shock (Theorem 4.2).
+The wedge enters the mean costate equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$, distorting mean actions relative to the single-agent benchmark.
+2.3 Computation
+Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system (Section 4), discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed.
+Parameters: $T=1$, $x_{0}=0$, $r=0.1$, no terminal cost.
+The Picard residual reaches $10^{-5}$ in $19$ iterations ( ${\approx}\,10$ ms in C++), decaying geometrically.
+[^1] An interactive browser visualization is available at https://sbabichenko.com/lqg.
+At longer horizons the binding constraint is numerical resolution (the filtering kernel develops near-singularities) rather than economic.
+2.4 Mean actions, separation failure, and the welfare cost of belief manipulation
+Since the stochastic integral in (2.4) has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks.
+Separation failure.
+Under perfect information, the mean policy is independent of signal precision.
+Decentralized information breaks this: precision changes the speed of opponents’ updating, creating an incentive to act on posteriors rather than on the state.
+The mechanism is the information wedge (2.5): it couples the costate to the unresolved kernel $\widetilde{X}^{2}_{t}(z)$ through the belief adjoint, introducing a dependence of the equilibrium policy on precision that is absent in the single-agent problem.
+More precisely, in the single-agent case the policy kernel factors as $D^{i}_{t}(u)=S(t)\cdot X_{t}(u)$ for a scalar Riccati gain $S(t)$ independent of signal precision.
+With two agents, the wedge prevents this factorization: changing $p$ changes the unresolved kernel, which changes the belief adjoint, which changes the policy kernel independently of the state estimate (Proposition 4.6).
+Figure 2: Equilibrium mean control $\bar{D}^{1}_{t}$ as a function of signal precision $p$, with the perfect-information benchmark (dashed).
+Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
+The wedge across precision asymmetries.
+Figures 34 display the information wedge and its consequences when player 1’s precision is fixed at $p_{1}=3$ and player 2’s precision $p_{2}$ varies.
+Figure 3: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying.
+Left: mean controls $\bar{D}^{1}_{t}$ (solid) and $\bar{D}^{2}_{t}$ (dashed) diverge as the precision gap widens.
+Center: the mean state path $\bar{X}_{t}$ tilts toward the better-informed player’s target.
+Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
+Figure 4: Mean information wedges $\bar{\mathcal{V}}^{1}_{t}$ and $\bar{\mathcal{V}}^{2}_{t}$ as $p_{2}$ varies ($p_{1}=3$ fixed).
+Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates.
+Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior.
+The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
+Information pooling as a Pigouvian intervention.
+Figure 5 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive (opposing targets $\pm 1$) and cooperative (common target 0) specifications.
+Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition.
+Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
+Remark 2.1 (Impulse responses do not depend on targets).
+The equilibrium impulse-response maps—how players respond to shocks—do not depend on the targets $\pm 1$.
+All filtering objects, unresolved components, and policy response maps are identical whether players cooperate or compete; only the deterministic tug-of-war over mean actions sees the strategic structure, through the information wedge.
+Figure 5: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying.
+Left: opposing targets ($\pm 1$).
+Right: common target (0).
+Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$.
+Blue: player 1; red: player 2.
+2.5 Optimal precision allocation
+Information pooling is a blunt instrument: it merges all signals into a common channel, eliminating some of the bilateral externality but forecloses targeted allocation.
+Given a fixed precision budget, how should a planner distribute it between the two players?
+Setup.
+Retain the state dynamics (2.1) and observation channels (2.2) with opposing targets ($\pm 1$), but allow asymmetric precisions $(p_{1},p_{2})$ subject to $p_{1}^{2}+p_{2}^{2}=\bar{P}$ and asymmetric effort costs $(r_{1},r_{2})$ with $r_{1}<r_{2}$, so that player 1 is the more efficient mover of the state.
+The quadratic budget reflects diminishing returns to concentration.
+A planner chooses the split to minimize aggregate equilibrium cost; for each candidate allocation we solve the full equilibrium via Picard iteration.
+Parameters: $\bar{P}=20$, $T=1$, with the fundamental volatility in (2.1) reduced to $\sigma=0.5$.
+Figure 6: Equilibrium quantities across the precision split, opposing targets ($\pm 1$), $p_{1}^{2}+p_{2}^{2}=\bar{P}=20$, $\sigma=0.5$, $T=1$.
+Top: asymmetric effort costs ($r_{1}=0.05$, $r_{2}=0.2$); bottom: symmetric ($r_{1}=r_{2}=0.1$).
+Left: individual mean control energies; center: total destructive effort (equilibrium solid, full-information dashed); right: mean controls and mean state (equilibrium solid, certainty-equivalent dotted).
+What one might expect.
+Under cooperation the answer is straightforward: give precision to the efficient player, perhaps sharing some to ease the inefficient player’s burden.
+Under competition, concentrating precision arms the efficient player to pull the state toward their target, so one might expect the planner to compensate by balancing precision to limit manipulation.
+The presumed tradeoff is productive efficiency versus strategic restraint.
+Information starvation.
+The tradeoff is real but the strategic-restraint side is far weaker than expected (Figure 6, top row, $r_{1}=0.05$, $r_{2}=0.2$).
+Concentrating all precision on the efficient player does shift the mean state further from zero, but the effect on total effort dominates: destructive effort drops by nearly half.
+The productive-efficiency and strategic-restraint channels point in the same direction: the presumed tradeoff does not exist.
+The mechanism is information starvation: denying the inefficient player signal access collapses their ability to exert mean control (top left panel: player 2’s effort falls to near zero).
+With player 2 largely neutralized, player 1 has less reason to manipulate as well, and both players’ effort falls.
+Conversely, giving all precision to the inefficient player produces the worst outcome: the cheap mover pushes hard against a well-informed opponent who retaliates sharply, maximizing bilateral waste.
+The planner’s optimal response is not to balance information but to create an information monopoly for the efficient player.
+The cost of one-sided dominance is a displacement in the mean state; the benefit is the collapse of bilateral waste.
+The collapse of bilateral waste dominates.
+Remark 2.2 (Commitment through incapacity).
+The mechanism parallels Schelling’s [30] observation that in a game of chicken, a driver who visibly removes the steering wheel forces the opponent to swerve.
+Here precision is common knowledge and enters the equilibrium through deterministic kernels, so denying the inefficient player signal access is a credible and observable disarmament: both players benefit not from chosen restraint but from one side’s visible inability to fight.
+Separation failure in a picture.
+The right panels display separation failure directly: certainty-equivalent mean controls (dotted) are invariant to the precision split, while equilibrium mean controls (solid) slope with the allocation through the belief adjoints.
+Under symmetric effort costs (bottom row) the pattern is symmetric around the equal split, confirming that the asymmetry above is driven by cost differences interacting with the information channel.
 3 The Decentralized LQG Game
 Informal description of the game.
-The class of games we study has three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon.
-Because actions enter the state, they simultaneously serve the physical objective and reshape what opponents observe, making each agent’s problem jointly strategic and informational.
-Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable—the same structure that underlies the workhorse models of [28, 26, 20] and the mean-field LQG literature [7], and that serves as the exact benchmark against which approximations can be assessed.
-2.1 Primitives
+The games we study have three features: (i) multiple agents share a common evolving state but observe it through private noisy channels; (ii) each agent’s action feeds back into the state dynamics and thereby into every other agent’s signal process; and (iii) agents minimize individual quadratic objectives over a finite horizon.
+Features (i)–(ii) generate the infinite belief hierarchy identified by Townsend; feature (iii) provides the linear–Gaussian structure that makes it tractable.
+3.1 State dynamics and objectives
 Consider $n$ players interacting over a finite time horizon $[0,T]$ on a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$.
@@
 All results extend to $D_{t}^{i}=B_{i}(t)U_{t}^{i}$ with control $U^{i}_{t}$ by appropriately modifying the cost function and first order conditions.
-Observations.
-Player $i$ does not observe $X_{t}$ directly. Instead, they observe a noisy measurement process $Y^{i}_{t}\in\mathbb{R}^{d}$ given by
-$dY^{i}_{t}=\sqrt{P^{i}(t)}X_{t}\,dt+dW^{i}_{t},\quad Y^{i}_{0}=0,$ (2.2)
-where: [^1] The square-root parameterization is convenient for interpreting $P^{i}(t)$ as a precision matrix, but all results extend to observation channels of the form $dY^{i}_{t}=H^{i}(t)X_{t}\,dt+E^{i}(t)\,dW_{t}$ with arbitrary deterministic $H^{i}(t)\in\mathbb{R}^{d_{i}\times d}$ and $E^{i}(t)E^{i}(t)^{\top}=I$, including correlated observation noises across players and signals that weight different state coordinates asymmetrically.
-All kernel formulas remain verbatim under this generalization.
-  • $W^{i}_{t}\in\mathbb{R}^{d}$ is a standard Brownian motion (player $i$’s observation noise),
-  • $P^{i}(t)\in\mathbb{R}^{d\times d}$ is symmetric positive semidefinite, and $\sqrt{P^{i}(t)}$ denotes its unique symmetric positive semidefinite square root.
-  • All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
-Notation for primitive shocks.
-We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$.
-For each $j\in\{0,\dots,n\}$, the block selector $E^{j}\in\mathbb{R}^{d\times(n+1)d}$ extracts the $j$-th channel: $E^{j}W_{t}=W^{j}_{t}$.
-The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$.
-Player $i$ directly observes $\Pi^{i}W$ and must infer $(I-\Pi^{i})W$ from drift-based learning.
-Player $i$’s information at time $t$ is their observation history, $\mathcal{F}^{i}_{t}:=\sigma(Y^{i}_{s}:s\leq t)$.
-An admissible control is $\mathcal{F}^{i}_{t}$-adapted with $\mathbb{E}[\int_{0}^{T}\|D^{i}_{t}\|^{2}\,dt]<\infty$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
-Unresolved impulse response.
-For each player $i$, define
-$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
-the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate.
-Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 3.1), with its algebraic representation in terms of $F^{i}$ given in (A.12).
 Objectives.
@@
 Opponents’ controls enter player $i$’s cost only through $X_{t}$, which they move via (3.1) and which player $i$ observes through (3.4).
+3.2 Information structure
+Observations.
+Player $i$ does not observe $X_{t}$ directly.
+Each player monitors the evolving state through private information sources: analyst reports, proprietary sensors, research output, market data feeds.
+Some sources arrive without cost (publicly posted prices, mandatory disclosures, direct observation of one’s own action) while others must be actively acquired or processed.
+The sources may be discrete (a finite set of news feeds) or continuous (a flow of independent measurements whose aggregate is naturally modeled by a Brownian sheet [17]).
+In this paper we take each player’s information-processing capacity as exogenous: the menu of sources and the attention allocated to each are fixed and common knowledge.
+The observation process summarizing the player’s information about the state then takes the form
+$dY^{i}_{t}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},\quad Y^{i}_{0}=0,$ (3.4)
+where $\Gamma^{i}(t)\in\mathbb{R}^{d\times d}$ is a deterministic observation gain and $E^{i}\in\mathbb{R}^{d\times(n+1)d}$ is the block selector extracting player $i$’s noise channel ($W^{i}_{t}:=E^{i}W_{t}$).
+All Brownian motions $\{W^{0},W^{1},\ldots,W^{n}\}$ are mutually independent.
+The precision matrix $P^{i}(t):=\Gamma^{i}(t)^{\top}\Gamma^{i}(t)$ measures the Fisher information rate about the state per unit time; the gain $\Gamma^{i}$ is the primitive, and $P^{i}$ is derived.
+The block-selector form is without loss: replacing $E^{i}$ by a general invertible noise loading $\Sigma^{i}(t)$ amounts to substituting $\Gamma^{i,\top}(\Sigma^{i}\Sigma^{i,\top})^{-1}$ for $\Gamma^{i,\top}$ throughout.
+Notation for primitive shocks.
+We collect all primitive noises into a single vector $W_{t}:=(W^{0}_{t},\dots,W^{n}_{t})^{\top}\in\mathbb{R}^{(n+1)d}$.
+The block projector $\Pi^{i}:=E^{i,\top}E^{i}$ is the identity on the $i$-th block and zero elsewhere: $\Pi^{i}v=(0,\dots,0,v^{i},0,\dots,0)$.
+The block-selector structure ensures that each player’s observation noise is an independent component of $W$, giving the noise-state its direct/indirect decomposition (Theorem 4.1): player $i$ (almost) directly observes $\Pi^{i}W$ and infers $(I-\Pi^{i})W$ from drift-based learning.
+Player $i$’s information at time $t$ is their observation history, $\mathcal{F}^{i}_{t}:=\sigma(Y^{i}_{s}:s\leq t)$.
+An admissible control is $\mathcal{F}^{i}_{t}$-adapted with $\mathbb{E}[\int_{0}^{T}\|D^{i}_{t}\|^{2}\,dt]<\infty$, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$.
+Unresolved impulse response.
+For each player $i$, define
+$\widetilde{X}^{i}_{t}(u):=X_{t}(u)-\widehat{X}^{i}_{t}(u),$
+the component of the state’s impulse response to the $u$-th shock that player $i$ has not yet resolved at time tt—large when the shock is recent or the observation channel weak, and shrinking as observations accumulate.
+Formally, $\widetilde{X}^{i}_{t}(u)=\operatorname{Cov}(X_{t},\,dW_{u}\mid\mathcal{F}_{t}^{i})/du$; it serves as the filtering gain in the noise-state update (Theorem 4.1), with its explicit formula in (A.12).
+Throughout Sections 45 the gain $\Gamma^{i}(\cdot)$ is fixed; Section 6 discusses the extension to endogenous precision.
 3.3 Nash Equilibrium
@@
 We write $D_{t}^{i}=D_{t}^{i}[Y^{i}]$, where the bracket notation $D_{t}^{i}[\cdot]$ denotes the map from the observation path $(Y^{i}_{s})_{s\leq t}$ to the action at time $t$.
-We identify a strategy with its induced adapted control process, assuming $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
+We require $\mathbb{E}\int_{0}^{T}\|D_{t}^{i}\|^{2}\,dt<\infty$.
 A profile $D^{*}=(D^{*,1},\ldots,D^{*,n})$ is a Nash equilibrium if for every player $i$ and every admissible alternative strategy map $\widetilde{D}^{i}[\cdot]$,
@@
 Equivalently: in a unilateral deviation by $i$, opponents’ maps are held fixed and are applied to the deviated signal paths.
-In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps are parameterized by deterministic maps that remain unchanged off equilibrium.
-The requirement that opponents’ strategy maps be held fixed under deviations—rather than their realized actions or their realized beliefs—is the formal expression of the Lucas critique in this environment: if player $i$ changes their control law, the state dynamics change, which changes every opponent’s signal process, filtering problem, and actions, which feeds back into the state.
+In a unilateral deviation by $i$, opponents’ strategy maps $Y^{k}_{\cdot\wedge t}\mapsto D_{t}^{k}[Y^{k}]$ are held fixed and applied to the perturbed observation histories; in the noise-state linear class these maps have deterministic coefficients that do not change when the state path changes.
+Holding opponents’ strategy maps fixed under deviations (rather than their realized actions or beliefs) is substantive here because signals are endogenous: changing one player’s control law changes the state dynamics, which changes every opponent’s signal, filtering problem, and realized action, even though their decision rules have not changed.
 This paper aims to establish that the class of noise-state linear strategies is closed under best responses and to derive a deterministic impulse-response system whose fixed points correspond to Nash equilibria in that class.
 Remark 3.2 (Subjective equilibrium and policy evaluation).
-The best-response closure (Corollary 3.5) does not require player $i$ to know opponents’ maps correctly: a player holding incorrect beliefs about opponents’ strategies still solves a well-posed best-response problem against the believed profile.
-A policy-maker can therefore let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (2.1), which depend on the actual strategy profile rather than on whether agents’ beliefs about each other are correct.
-Common knowledge of rationality is sufficient for Nash equilibrium but is not required to apply the characterization as a best-response operator.
+The best-response closure (Corollary 4.5) does not require player $i$ to know opponents’ maps correctly.
+A policy-maker can let each agent optimize under its own beliefs, then evaluate welfare through the realized state dynamics (3.1); common knowledge of rationality is sufficient for Nash equilibrium but not required to apply the characterization as a best-response operator.
 Noise-state.
-For each player $i$ and each $t\in[0,T]$, define the noise-state as the conditional mean path of the aggregated primitive noise:
+For each player $i$ and each $t\in[0,T]$, the noise-state encountered informally in Section 2.2 is defined as the conditional mean path of the aggregated primitive noise:
 $\widehat{W}_{t}^{i}(u):=\mathbb{E}[W_{u}\mid\mathcal{F}_{t}^{i}],\qquad 0\leq u\leq t.$ (3.5)
@@
 Moreover, by Itô isometry the square-integrability condition above is equivalent to $L_{t}\in L^{2}$ for each $t$.
-Finally, the martingale representation theorem implies that any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable admits a representation of the form $\int_{0}^{t}Z_{s}\,dW_{s}$ with a stochastic integrand $Z$; the defining restriction of the linear class is that the response map $L(t,\cdot)$ is deterministic (for each fixed $t$).
+By the martingale representation theorem, any $L^{2}$ $\mathcal{F}_{t}$-measurable random variable has a stochastic integral representation against $W$; the defining restriction of the linear class is that the integrand $L(t,\cdot)$ is deterministic.
+Equivalently, Definition 3.2 is the continuous-time analogue of an MA($\infty$) representation: $L(t,s)$ plays the role of the moving-average coefficient at lag tst-s, with the Itô integral replacing the infinite sum against i.i.d.
+innovations.
 Definition 3.3 (Noise-state Linear control / strategy).
-Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (Section 3).
+Fix a player $i$ and recall their noise-state path $\widehat{W}_{t}^{i}(\cdot)$ (3.5).
 An admissible control $D^{i}=(D_{t}^{i})_{t\in[0,T]}$ is a noise-state linear control if it is affine in the noise state.
@@
 $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u),\qquad\int_{0}^{t}\|D^{i}_{t}(u)\|^{2}\,du<\infty,$
-and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, writing $\widehat{W}^{\,i}[Y^{i}]$ to emphasize dependence on the observation path, this defines the strategy map
+and we extend $D^{i}_{t}(u)$ by 0 for $u>t$. Equivalently, this defines the strategy map
 $Y^{i}_{\cdot\wedge t}\longmapsto D_{t}^{i}[Y^{i}]:=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{\,i}[Y^{i}](u).$
 4 Main results
-We work in the linear–Gaussian model of Section 2.
-Two results drive the analysis.
-First, when opponents use noise-state linear strategies, the filtering equations inherit deterministic impulse responses (Theorem 3.1).
-Second, the best response over the full admissible class is itself noise-state linear (Theorem 3.2), so Nash equilibrium reduces to a deterministic fixed point in impulse-response maps.
-The filtering closure (Theorem 3.1) shows that noise-state dynamics have deterministic kernels; the best-response closure (Theorem 3.2) derives the adjoint system and the information wedge; Proposition 3.6 then decomposes equilibrium cost into a certainty-equivalent term and an information cost linked to endogenous precision.
+This section derives the objects previewed in Section 2 in the general $n$-player model of Section 3.
+The strategy is analogous to Harsanyi’s [12]: just as a common prior over a type space closes the belief hierarchy in static incomplete-information games, the common probability space of primitive shocks $W=(W^{0},\ldots,W^{n})$ closes it here, with all higher-order conditional expectations following by linear projection onto deterministic impulse-response maps.
+The difference is that Harsanyi’s types are drawn once while noise unfolds continuously and actions reshape signals, so the closure requires a fixed point: the maps governing belief updating must be consistent with the strategy profile they induce (Figure 7).
+(i) Candidate Opponents use Noise-State Linear controls (ii) Closure State & filters have det.
+response maps (iii) Deviations Fixed maps $\Rightarrow$ det.
+propagation (iv) Best response FOC $\Rightarrow$ control noise-state linear (v) Fixed point BR map closes $\Longleftrightarrow$ Nash eq.
+Figure 7: The Noise-State Linear fixed-point loop.
+(i) Assume opponents use Noise-state Linear controls.
+(ii) The state and filtering equations have deterministic impulse responses.
+(iii) Under fixed strategy maps, unobserved deviations propagate deterministically.
+(iv) The maximum principle produces a best response that is itself noise-state linear, closing the loop.
+(v) A fixed point of this best-response map is a Nash equilibrium.
 Filtering closure.
@@
 and player $i$ observes
-$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
+$dY_{t}^{i}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},$
 with deterministic $P^{i}(t)\succeq 0$ and block selector $E^{i}$.
-Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$.
+Let $\Pi^{i}:=E^{i,\top}E^{i}$ and $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$, and define the innovation $dI_{t}^{i}:=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$.
 (i) The estimated-noise path increments admit the decomposition
@@
 where the deterministic kernel $F^{i}$ is given explicitly by
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (3.2)
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (4.2)
 (ii) The mixed $(t,u)$-update satisfies
-$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du.$
+$d_{t}\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,dI_{t}^{i}\,du.$
 The unresolved kernel $\widetilde{X}^{i}$ is determined algebraically from $F^{i}$ by (A.12); the filtering kernel $F^{i}$ is the unique solution of the forward evolution system (A.14)–(A.15) (Definition A.1).
@@
 Discretizing time, each observation increment is affine in the noise increments $\Delta W_{1:k}$, so by joint Gaussianity the conditional mean $\mathbb{E}[\Delta W_{1:k}\mid\Delta Y^{i}_{1:k}]$ is a deterministic linear projection; each layer of “beliefs about beliefs” is a further deterministic composition of projections, which is why the regress closes.
-The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component—absolutely continuous in $u$ with density driven by $F^{i}$—captures drift-based inference about unobserved noise coordinates at scale $O(dt)$.
+The decomposition in part (i) reflects a scale separation: the direct component $\Pi^{i}\,dW_{u}$ is $O(1)$ (player $i$ observes their own noise contemporaneously), while the indirect component captures drift-based inference about unobserved shocks at scale $O(dt)$.
 Best-response closure and information wedges.
@@
 Structure.
-Because the equilibrium control is a noise-state linear process $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,dW_{u}$ with deterministic mean and kernel, the backward system splits into two structurally identical subsystems—one for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$, one for the response-map coefficients $(H^{X,i},H^{k,i})$—that should be read in parallel.
-In each subsystem, the single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state.
-This channel introduces belief-adjoint variables $(\bar{H}^{k,i},H^{k,i})$ that track the shadow value of perturbing opponent $k$’s noise-state estimate, coupled into the physical costate through the opponents’ filtering gains.
-Mean backward system. The mean physical costate $\bar{H}^{X,i}_{t}$ and mean belief-adjoints $\bar{H}^{k,i}_{t}(u)$ satisfy
-$\frac{d}{dt}\bar{H}^{X,i}_{t}$ $=-\big(G^{XX,i}(t)\bar{X}_{t}+\bar{G}^{X,i}_{t}\big)-A(t)^{\top}\bar{H}^{X,i}_{t}-\underbrace{\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz}_{=:\;\bar{\mathcal{V}}^{i}_{t},\;\text{mean information wedge}},$ (3.3)
-$\frac{d}{dt}\bar{H}^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}\bar{H}^{X,i}_{t}+\int_{0}^{t}X_{t}(u)^{\top}P^{k}(t)\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (3.4)
-with terminal conditions $\bar{H}^{X,i}_{T}=G^{XX,i}(T)\bar{X}_{T}+\bar{G}^{X,i}_{T}$ and $\bar{H}^{k,i}_{T}(\cdot)=0$.
-The first two terms in (3.3) are the standard single-agent costate equation; the third term $\bar{\mathcal{V}}^{i}_{t}$ is the mean information wedge, absent in the single-agent problem.
-In the belief-adjoint equation (3.4), the forcing $D^{k}_{t}(u)^{\top}\bar{H}^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock.
-The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$.
-Kernel backward system.
-The kernel costate $H^{X,i}_{t}(r)$ and kernel belief-adjoints $H^{k,i}_{t}(u,r)$ satisfy the identical system with $\bar{X}_{t}\to X(t,r)$ and $\bar{G}^{X,i}_{t}\to G^{X,i}_{t}(r)$; the kernel information wedge $\mathcal{V}^{i}_{t}(r)$ is the corresponding substitution of $\bar{\mathcal{V}}^{i}_{t}$.
-We define the (full) information wedge as the linear process
-$\mathcal{V}^{i}_{t}:=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r},$ (3.5)
-which is the total distortion of the physical costate relative to the single-agent benchmark.
-The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ captures how belief manipulation distorts the equilibrium level (mean actions); the response-map wedge $\mathcal{V}^{i}_{t}(r)$ captures how it distorts the equilibrium response to shocks.
-Both components—and hence $\mathcal{V}^{i}_{t}$ itself—vanish when signals are exogenous.
-Policy interpretation of the wedge.
-The information wedge $\mathcal{V}^{i}_{t}$ vanishes identically when signals are exogenous to controls, recovering the standard LQG costate equation.
-When controls enter the state dynamics, a change in signal precision $P^{k}(t)$ also operates through the belief-manipulation channel: it alters how sharply opponents’ posteriors respond to drift perturbations (through $\widetilde{X}^{k}$), which alters the shadow value of shifting those posteriors (through $\bar{H}^{k,i}$), which feeds back into the equilibrium control through (3.3).
-Any model that omits this channel misattributes the strategic distortion to the physical response: the mean error is exactly $\bar{\mathcal{V}}^{i}_{t}$, and the shock-by-shock error is $\mathcal{V}^{i}_{t}(r)$.
-The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ governs the distortion in equilibrium mean actions (Section 4.3); the kernel wedge $\mathcal{V}^{i}_{t}(r)$ governs the distortion in equilibrium impulse responses.
-Section 4 quantifies both channels numerically.
+All objects below are linear processes (Definition 3.2): deterministic mean plus a stochastic integral against $W$ with deterministic kernel.
+The single-agent costate acquires an additional coupling term: player $i$’s action shifts the state, which shifts opponent $k$’s signal, posterior, and action, feeding back into the state.
+This channel introduces belief-adjoint processes $H^{k,i}_{t}(u)$ that track the shadow value of perturbing opponent $k$’s noise-state estimate.
+Backward system. The physical costate $H^{X,i}_{t}$ and belief-adjoints $H^{k,i}_{t}(u)$ are linear processes satisfying
+$\frac{d}{dt}H^{X,i}_{t}$ $=-\big(G^{XX,i}(t)X_{t}+G^{X,i}_{t}\big)-A(t)^{\top}H^{X,i}_{t}-\underbrace{\sum_{k\neq i}P^{k}(t)\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz}_{=:\;\mathcal{V}^{i}_{t},\;\text{information wedge}},$ (4.3)
+$\frac{d}{dt}H^{k,i}_{t}(u)$ $=-\big(D^{k}_{t}(u)\big)^{\top}H^{X,i}_{t}+X_{t}(u)^{\top}P^{k}(t)\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz,\qquad k\neq i,$ (4.4)
+[…]
+The first two terms in (4.3) are the standard single-agent costate equation; the third term $\mathcal{V}^{i}_{t}$ is the information wedge, absent in the single-agent problem.
+In the belief-adjoint equation (4.4), the forcing $D^{k}_{t}(u)^{\top}H^{X,i}_{t}$ transmits the physical costate into the belief channel through opponent $k$’s policy kernel: the larger $D^{k}_{t}(u)$, the more opponent $k$’s action at time $t$ responds to the $u$-th shock, and the more valuable it is for player $i$ to have shifted $k$’s estimate of that shock.
+The integral term is a self-coupling through which belief perturbations propagate forward via the filtering gain $\widetilde{X}^{k}$; the same inner product $\int_{0}^{t}\widetilde{X}^{k}_{t}(z)\,H^{k,i}_{t}(z)\,dz$ that defines the wedge in (4.3) drives the self-coupling in (4.4), so the belief-adjoint feeds back into itself through the channel it prices.
+Since all coefficients in (4.3)–(4.4) are deterministic, taking expectations recovers a closed system for the mean coefficients $(\bar{H}^{X,i},\bar{H}^{k,i})$; matching the coefficient of $dW_{r}$ recovers a closed system for the response-map coefficients $(H^{X,i}(\cdot,r),H^{k,i}(\cdot,\cdot,r))$.
+The two subsystems decouple and are solved separately in the numerical scheme of Section 2.
+The mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the general-$n$ counterpart of the object displayed in Figures 45; the kernel wedge $\mathcal{V}^{i}_{t}(r)$ distorts equilibrium impulse responses.
+Both vanish when signals are exogenous.
 Corollary 4.3 (Exogenous-signal reduction).
 If controls do not enter the state dynamics, then $\mathcal{V}^{i}_{t}\equiv 0$ for all $i$ and $t$, the belief-adjoint equations (4.4) decouple, and the equilibrium reduces to $n$ independent single-agent LQG problems against a common exogenous state.
+Remark 4.2 (Mean-field approximation error).
+Mean-field models zero both wedge components by construction: in the continuum limit no single agent’s action measurably affects any other agent’s signal.
+Eliminating both components in a finite-$n$ model requires either exogenous signals (Corollary 4.3) or $n\to\infty$ with vanishing individual impact—the first is ruled out when information is costly [10], and concentration of informed trading pushes real markets toward finite $n$.
+In the Kyle–Back embedding (Section 5), the sign symmetry $W\mapsto-W$ zeros the mean wedge while the kernel wedge remains active; this symmetry breaks in any market with asymmetric positions.
 Lemma 4.4 (Strict convexity of the best-response problem).
@@
 $t\in[0,T]$,
-$G^{DD,i}(t)D_{t}^{i}+\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)=0,$
-and hence
-$D_{t}^{i}=-(G^{DD,i}(t))^{-1}\bar{H}^{X,i}_{t}-(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u).$
-In particular, every best response over the full admissible class is itself a noise-state linear control.
+$G^{DD,i}(t)D_{t}^{i}+\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}]=0,\qquad\text{i.e.,}\qquad D_{t}^{i}=-(G^{DD,i}(t))^{-1}\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}].$
+Since $H^{X,i}_{t}$ is a linear process and $\mathbb{E}[H^{X,i}_{t}\mid\mathcal{F}_{t}^{i}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,d\widehat{W}_{t}^{i}(u)$, every best response over the full admissible class is itself a noise-state linear control.
 Proof.
-With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [16], Theorem 7.12).
+With opponents’ maps fixed and linear, the innovation $dI_{t}^{i}:=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$ is a standard $(\mathcal{F}^{i},\mathbb{P})$-Brownian motion whose law does not depend on $D^{i}$ (Liptser–Shiryaev [19], Theorem 7.12).
 Since $\widehat{W}_{t}^{i}(\cdot)$ is a deterministic linear functional of $I^{i}_{[0,t]}$ (Theorem 4.1 in innovations coordinates), player $i$’s control is a deterministic affine functional of the exogenous process $I^{i}$.
 The induced objective is strictly convex (Lemma 4.4), so the stationarity condition (B.8) is necessary and sufficient, yielding the stated formula.
-Note that the exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem C.5).
+The exogeneity of $I^{i}$ is a property of the best-response problem: the response-maps mediating between $I^{i}$ and the noise-state depend on the full strategy profile through the forward closure, so the equilibrium remains a genuine fixed point (Theorem D.5).
 ∎
 Remark 4.3.
-Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 3.5), any Nash equilibrium of the restricted game in which players are confined to noise-state linear controls is automatically a Nash equilibrium of the unrestricted game.
-The open question is whether equilibria outside the deterministic impulse-response class exist; Section 6 discusses a program for ruling them out.
+Since every best response over the full admissible $L^{2}$ class is noise-state linear (Corollary 4.5), any Nash equilibrium in noise-state linear controls is automatically a Nash equilibrium of the unrestricted game; whether equilibria outside this class exist is discussed in Section 6.
+Remark 4.4 (Single-player reduction).
 When $n=1$, $\mathcal{V}^{i}_{t}\equiv 0$ and the ansatz $H^{X,i}_{t}(r)=S(t)X(t,r)$ reduces the backward system to the standard matrix Riccati equation for $S$, recovering classical separated LQG control.
-Risk-sensitive extension.
-Replace (2.4) with the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$, which penalizes cost variance in addition to cost mean ($\theta_{i}>0$ governs risk aversion; $\theta_{i}\to 0$ recovers (2.4)).
-The entire framework carries over: the best-response FOC is identical to Corollary 3.5 with the noise-state $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state $\widehat{W}^{\,i,\theta}_{t}$ that tilts the conditional mean of $W$ toward directions that reduce expected cost and rescales conditional precision along directions where cost variance is large.
-The filtering closure is unchanged; certainty equivalence fails because the quadratic structure of the cost reshapes the posterior without altering the observation channel.
-The full statement and derivation are in Appendix B.6 (Theorem B.1).
-Information cost and endogenous precision.
-At the equilibrium profile, the control gap between full and partial information is
-$D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$ (3.6)
-where $D^{i,\mathrm{full}}_{t}:=-(G^{DD,i}(t))^{-1}M^{i}_{t}$ is the best response under full observation of $W$, $M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u}$ is the stochastic costate, and $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$ is the noise-state estimation error.
-Proposition 3.6 (Information cost).
-At the equilibrium profile,
-$\Delta J^{i}:=J^{i}(D^{i};D^{*,-i})-J^{i}(D^{i,\mathrm{full}};D^{*,-i})=\int_{0}^{T}\mathrm{tr}\!\left[G^{DD,i}(t)^{-1}\,\Sigma^{M,i}_{t}\right]\mathrm{d}t,$ (3.7)
-where $\Sigma^{M,i}_{t}:=\operatorname{Cov}(M^{i}_{t}\mid\mathcal{F}^{i}_{t})$ is the costate estimation error covariance (deterministic by Gaussianity).
-Moreover, $\Delta J^{i}$ depends on $P^{i}$ only through the filtering kernels $(\widetilde{X}^{i},F^{i})$ and is independent of player $i$’s control $D^{i}$.
-The total cost decomposes as
-$J^{i}(D^{i};D^{*,-i})=\underbrace{J^{i}(D^{i,\mathrm{full}};D^{*,-i})}_{\text{certainty-equivalent cost}}+\;\underbrace{\Delta J^{i}(P^{i})}_{\text{information cost}},$ (3.8)
-This decomposition underlies the welfare comparisons in Section 4: the information wedge enters through $H^{X,i}$ inside $\Sigma^{M,i}_{t}$.
-The joint optimization over $(D^{i},P^{i})$ separates: $D^{i}$ is determined by Corollary 3.5 for any given $P^{i}$, and $P^{i}$ solves
-$\min_{P^{i}(\cdot)\succeq 0}\quad\Delta J^{i}(P^{i})+\int_{0}^{T}c^{i}(t,P^{i}(t))\,\mathrm{d}t,$ (3.9)
-where $c^{i}$ is a convex attention cost.
-Corollary 3.7 (Endogenous precision closure).
-If each opponent $k\neq i$ uses a deterministic precision path $P^{k}(\cdot)$, then $\Delta J^{i}$ is a deterministic functional of $P^{i}(\cdot)$, and the best-response precision is deterministic.
+Proposition 4.6 (Failure of separation).
+For $n\geq 2$ with endogenous signals, the equilibrium policy kernel $D^{i}_{t}(u)$ does not factor as $S(t)\cdot X_{t}(u)$ for any matrix-valued function $S$, unless the information wedge vanishes (Corollary 4.3).
+A change in signal precision alters the policy kernel itself, not merely the estimate to which it is applied.
 Proof.
-Immediate from Proposition 3.6: the forward environment and backward adjoint are deterministic, so (3.9) is a deterministic variational problem. ∎
-Remark 3.3 (Connection to rational inattention).
-When $n=1$ the FOC for (3.9) reduces to the continuous-time analogue of Sims [27]; with $n>1$ the information wedge enters through $H^{X,i}$, so the marginal value of precision depends on the strategic environment—a more responsive opponent raises the return to filtering because better inference enables more effective belief manipulation.
+With $n\geq 2$ the information wedge couples the costate to belief adjoints $H^{k,i}_{t}(u)$, which depend on $\widetilde{X}^{k}_{t}(u)$—the portion of the state response opponent $k$ has not yet learned.
+This introduces $(t,u)$-dependence that cannot factor through $X_{t}(u)$.
+Perturbing $P^{i}$ changes $\widetilde{X}^{k}$ through the equilibrium, changing $H^{k,i}$ and hence the policy kernel independently of the state estimate.
+The factorization holds if and only if $\mathcal{V}^{i}_{t}\equiv 0$.
+∎
+Remark 4.5 (Risk-sensitive extension).
+The framework extends to the entropic risk measure $J^{i}_{\theta_{i}}:=\theta_{i}^{-1}\log\mathbb{E}[e^{\,\theta_{i}\mathcal{C}^{i}}]$: the best-response FOC is identical to Corollary 4.5 with $\widehat{W}_{t}^{i}$ replaced by a risk-adjusted noise-state that tilts conditional means and rescales conditional precision through the quadratic cost kernel (Appendix B.6, Theorem B.1).
 Remark 4.6 (Infinite horizon and the frequency domain).
-Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [12, 14, 24] to endogenous signals.
+Under time-homogeneous primitives, response maps become time-invariant K(t,u)=k(tu)K(t,u)=k(t-u) and the fixed point becomes one in transfer functions, extending [14, 16, 27] to endogenous signals, where the transfer function of the equilibrium filter is itself an equilibrium object.
 Rational transfer functions correspond to finite-dimensional Markov sufficient statistics, so the pole structure of equilibrium kernels diagnoses whether an exact state-space reduction exists.
@@
 For convenience, we collect the kernel objects introduced above.
-$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}:=E^{i,\top}E^{i}$ selector / projector
+$W=(W^{0},\dots,W^{n})$ primitive noise $E^{i},\;\Pi^{i}$ selector / projector
 $\widehat{W}_{t}^{i}(u)$ noise-state $X_{t}(s)$ state impulse response
@@
 $D^{i}_{t}(u)$ policy impulse response $\mathcal{D}^{i}_{t}(s)$ primitive-noise control
-$(\bar{H}^{X,i},H^{X,i})$ physical adjoints $(\bar{H}^{k,i},H^{k,i})$ belief adjoints
-$\mathcal{V}^{i}_{t}=\bar{\mathcal{V}}^{i}_{t}+\int_{0}^{t}\mathcal{V}^{i}_{t}(r)\,dW_{r}$ information wedge
-4 Bilateral belief manipulation in a two-player game
-To make the general theory concrete, we specialize to the simplest game that exhibits bilateral belief manipulation: two symmetric players with scalar state, opposing targets, and equal signal precision.
-This strips away all structure except the core strategic friction.
-Neither player observes the state directly; each receives a private diffusion signal of precision $p>0$.
-Because each player’s control enters the state dynamics, it reshapes the opponent’s signal and thereby the opponent’s posterior.
-The information wedge $\mathcal{V}^{i}_{t}$ of Theorem 3.2 prices this externality exactly; the mean wedge $\bar{\mathcal{V}}^{i}_{t}$ is the component we visualize below.
-4.1 Primitives, objectives, and symmetry
-Fix a horizon $[0,T]$. The state $X_{t}\in\mathbb{R}$ satisfies
-$dX_{t}=\big(D_{t}^{1}+D_{t}^{2}\big)\,dt+dW_{t}^{0},\qquad X_{0}=x_{0},$ (4.1)
-where $W^{0}$ is a standard Brownian motion. Player $i\in\{1,2\}$ does not observe $X$ directly; instead they observe the diffusion channel
-$dY_{t}^{1}=\sqrt{p}X_{t}\,dt+dW_{t}^{1},\qquad dY_{t}^{2}=\sqrt{p}X_{t}\,dt+dW_{t}^{2},$ (4.2)
-with $(W^{0},W^{1},W^{2})$ mutually independent and $p>0$ representing root precision. Player $i$ chooses $D^{i}$ adapted to $\mathcal{F}_{t}^{i}=\sigma(Y_{s}^{i}:s\leq t)$.
-Running costs are symmetric tracking losses with quadratic effort penalty $r>0$:
-$J^{1}(D^{1};D^{2})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}-1)^{2}+r\,(D_{t}^{1})^{2}\Big)\,dt,\quad J^{2}(D^{2};D^{1})=\mathbb{E}\!\int_{0}^{T}\Big((X_{t}+1)^{2}+r\,(D_{t}^{2})^{2}\Big)\,dt.$ (4.3)
-There is no terminal cost in this worked example.
-The model is invariant under the sign/label symmetry
-$(X,\;Y^{1},\;Y^{2},\;D^{1},\;D^{2})\ \longmapsto\ (-X,\;Y^{2},\;Y^{1},\;-D^{2},\;-D^{1}).$
-[…]
-Figure 2: Information-to-action feedback in the worked example. Actions move $X_{t}$ and shift opponents’ beliefs, so equilibrium mean actions depend on precision $p$.
-4.2 Numerical equilibrium and kernel visualization
-We now present the numerical equilibrium obtained from the Picard iteration described above.
-All figures below are generated from the converged kernel profile $D^{1}$, its primitive-noise representation $\mathcal{D}^{1}$, the state kernel $X$, and the filtering kernel $F^{1}$.
-Computation.
-Equilibrium is computed by iterating the best-response operator on the deterministic impulse-response system, discretized on a uniform grid of $N=40$ points; no Monte Carlo simulation is needed.
-At each iterate the forward environment $(X,\widetilde{X}^{i},F^{i})$ is recomputed from the closure equations, the backward adjoint $H^{X,i}$ is solved by explicit backward Euler, and the updated policy is blended into the current profile.
-Parameters: $T=1$, $x_{0}=0$, $r=0.1$, $G^{XX,i}(T)=0$.
-The outer residual reaches $10^{-5}$ in $19$ iterations( ${\approx}\,10$ ms in C++).
-At longer horizons the binding constraint is representational (the filtering kernel becomes more singular) rather than economic.
-Convergence of the fixed point.
-The outer Picard residual decays geometrically, confirming convergence of the fixed-point scheme in this symmetric benchmark; we use the iteration as a constructive computation rather than an existence proof.
-Impulse-response viewpoint.
-The key distinction is between $D^{1}_{t}(s)$ (response to a shock as inferred through the filter).
-(a) Primitive-noise state kernel $X_{t}(s)$.
-(b) Equilibrium noise-state feedback kernel $D^{1}_{t}(s)$.
-Figure 3: Equilibrium kernel time-slices, shown channel-by-channel for $(W^{0},W^{1},W^{2})$ with curves colored by evaluation time $t$.
-(a) Both players jointly stabilize older fundamental shocks (decay in the $W^{0}$ panel); observation-noise channels show smaller, opposed effects mediated by the belief loop.
-(b) The system-noise channel dominates; observation-noise channels exhibit curvature reflecting drift-based inference.
-4.3 Mean actions, separation failure, and bilateral belief manipulation
-Since the stochastic integral in $D_{t}^{i}=\bar{D}^{i}_{t}+\int_{0}^{t}D^{i}_{t}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ has zero unconditional mean, $\mathbb{E}[D_{t}^{i}]=\bar{D}^{i}_{t}$ isolates the deterministic tug-of-war over the target $\pm 1$, while the kernel term captures impulse responses to posterior shocks.
-Under perfect information, the mean policy does not depend on signal precision—a certainty-equivalence property.
-Decentralized information breaks this: changing precision changes the speed at which opponents update, giving each player a reason to act on posteriors rather than just on the physical state.
-Figure 4: Equilibrium mean control $\bar{D}^{1}_{t}$ as a function of signal precision $p$, with the perfect-information benchmark (dashed).
-Low precision makes opponents’ posteriors sluggish, amplifying the incentive to manipulate beliefs; as $p\to\infty$ the mean policy converges to the perfect-information limit.
-The information wedge as a dynamic rent density.
-In the general theory (Theorem 3.2), the backward equation for $\bar{H}^{X}$ contains opponent-filter coupling terms that vanish under perfect information or in the single-player case.
-Here these terms aggregate into
-$\bar{\mathcal{V}}^{1}_{t}=P^{2}(t)\int_{0}^{t}\widetilde{X}^{2}_{t}(z)\cdot\bar{H}^{2}_{t}(z)\,dz,$
-which enters the mean adjoint equation as $\tfrac{d}{dt}\bar{H}^{X}_{t}=-(\bar{X}_{t}-1)-\bar{\mathcal{V}}^{1}_{t}$.
-The unresolved kernel $\widetilde{X}^{2}_{t}(z)$ measures how much of the $z$-th shock remains unresolved by player 2 at time $t$; the belief-adjoint $\bar{H}^{2}_{t}(z)$ measures the shadow value to player 1 of that shift.
-Their inner product $\bar{\mathcal{V}}^{1}_{t}$ prices the mean channel of bilateral belief manipulation at time $t$; Figures 56 visualize this object across precision asymmetries.
-Figure 5: Asymmetric equilibrium with $p_{1}=3$ fixed and $p_{2}$ varying.
-Left: mean controls $\bar{D}^{1}_{t}$ (solid) and $\bar{D}^{2}_{t}$ (dashed) diverge as the precision gap widens.
-Center: the mean state path $\bar{X}_{t}$ tilts toward the better-informed player’s target.
-Right: aggregate mean effort $|\bar{D}^{1}|+|\bar{D}^{2}|$ exceeds the perfect-information benchmark (gray dashed), with the excess growing in the precision asymmetry.
-Figure 6: Mean information wedges $\bar{\mathcal{V}}^{1}_{t}$ and $\bar{\mathcal{V}}^{2}_{t}$ as $p_{2}$ varies ($p_{1}=3$ fixed).
-Both wedges are hump-shaped in $t$, confirming that belief manipulation is most valuable at intermediate dates.
-Increasing the opponent’s precision amplifies the wedge: a better-informed opponent reacts more sharply to drift, raising the marginal value of manipulating their posterior.
-The wedges vanish at the terminal date, consistent with the zero terminal condition on the belief adjoint.
-Figure 7: Equilibrium costs under private signals (dashed) and pooled signals (solid) with $p_{1}=3$ fixed and $p_{2}$ varying.
-Left: competitive targets ($\theta_{1}=1$, $\theta_{2}=-1$).
-Right: cooperative targets ($\theta_{1}=\theta_{2}=0$).
-Pooling is Pareto-improving in both cases, but the gain is an order of magnitude larger under competition, where it eliminates the bilateral belief-manipulation externality priced by $\mathcal{V}^{i}_{t}$.
-Blue: player 1; red: player 2.
-Information pooling as a Pigouvian intervention.
-Figure 7 compares equilibrium costs under private signals with those under a pooled signal of precision $P^{1}+P^{2}$, for both competitive ($\theta_{1}=1$, $\theta_{2}=-1$) and cooperative ($\theta_{1}=\theta_{2}=0$) targets.
-Pooling is Pareto-improving in both cases, but the gains differ by an order of magnitude: roughly 0.020.020.050.05 under cooperation versus 0.30.30.50.5 under competition.
-Since the estimation benefit of pooling enters identically regardless of the target structure, the difference isolates the strategic externality channel priced by $\mathcal{V}^{i}_{t}$: nearly all of the welfare gain from mandatory disclosure comes from eliminating bilateral belief manipulation, not from improving state estimation.
-5 Market microstructure as a special case
-The two-player example operates within the quadratic-cost framework of Section 2.
-We now show that the noise-state technique embeds the Kyle–Back model with no additional machinery, despite traders maximizing expected profits against the market maker’s price rather than minimizing a quadratic objective.
-The embedding works because $\lambda(t)>0$ provides the strict convexity needed for the best-response closure, the filtering closure applies verbatim to the stacked observation channel, and the backward system of Appendix D yields the optimality condition.
-We derive the equilibrium characterization for a single informed trader and identify the two channels through which belief manipulation operates; the extension to multiple asymmetrically informed traders follows the same backward system and is a natural application of the general theory.
+$H^{X,i}_{t}$ physical costate (linear process) $H^{k,i}_{t}(u)$ belief adjoints (linear process)
+$\mathcal{V}^{i}_{t}$ information wedge (linear process)
+5 Trading under asymmetric informedness
+The noise-state technique extends to strategic trading with asymmetric private learning, a setting where guess-and-verify breaks down once traders differ in precision or learn continuously from private sources.
+The price each trader faces decomposes into an exogenous component and a history-linear function of their own demand, with Kyle’s lambda playing the role of $G^{DD,i}$; the quadratic structure is distributed across time, so the first-order condition yields an integral equation whose time derivative pins down the optimal trading rate.
 5.1 Model and embedding
-Adjoin a fundamental-value Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$ and projector $\Pi^{V}$.
-The fundamental is $V_{t}:=\sigma_{V}W_{t}^{V}$.
-All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dS_{t}^{i}=\sqrt{p_{i}}\,V_{t}\,dt+dW_{t}^{i}$.
-The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]=\sigma_{V}\int_{0}^{t}E^{V}d_{u}\widehat{W}_{t}^{0}(u)$.
-Trader $i$ maximizes $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$.
-By symmetry $W\mapsto-W$, all mean terms vanish.
-Embedding.
-Rescaling the order-flow channel to unit noise, trader $i$’s stacked observation is
-$d\!\begin{pmatrix}Z_{t}/\sigma_{0}\\ S_{t}^{i}\end{pmatrix}=\underbrace{\begin{pmatrix}\sigma_{0}^{-1}\\[2.0pt] \sqrt{p_{i}}\end{pmatrix}}_{\sqrt{P^{i}}}\!\!\begin{pmatrix}\text{demand drift}\\ \sigma_{V}E^{V}W_{t}\end{pmatrix}\!dt\;+\;\underbrace{\begin{pmatrix}E^{0}\\ E^{i}\end{pmatrix}}_{\mathcal{E}^{i}}\!dW_{t},$ (5.1)
-which is the observation equation of Section 2 with $\sqrt{P^{i}}=(\sigma_{0}^{-1},\sqrt{p_{i}})^{\top}$ and $E^{i}\to\mathcal{E}^{i}$; the market maker has $\sqrt{P^{0}}=\sigma_{0}^{-1}$, $\mathcal{E}^{0}=E^{0}$.
-Kyle’s lambda.
+There are $n$ informed traders and a competitive market maker.
+A fundamental $V_{t}:=\sigma_{V}W_{t}^{V}$ is driven by a Brownian motion $W^{V}$ independent of $(W^{0},\dots,W^{n})$, with selector $E^{V}$.
+All players observe aggregate order flow $dZ_{t}=\sum_{j}D_{t}^{j}\,dt+\sigma_{0}\,dW_{t}^{0}$; trader $i$ additionally observes a private signal $dY_{t}^{i}=\Gamma^{i}(t)\,V_{t}\,dt+dW_{t}^{i}$.
+The market maker observes only $Z$ and posts the competitive price $P_{t}:=\mathbb{E}[V_{T}\mid\mathcal{F}_{t}^{0}]$.
+Trader $i$ maximizes expected profit $\pi^{i}:=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$.
+By the symmetry $W\mapsto-W$, all mean terms vanish.
+Stacking trader $i$’s two observation channels into a single equation recovers the observation model of Section 3 (Appendix E), so the filtering machinery of Appendix A applies verbatim.
+Price decomposition.
+Fix opponents’ noise-state linear strategy maps. The market maker’s filtering gains are deterministic, so the price decomposes as
+$P_{t}\;=\;P_{t}^{-i}\;+\;\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds,$ (5.1)
+where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ off equilibrium (the price that would obtain if trader $i$ submitted zero demand) and $\ell(t,s)$ is a deterministic kernel encoding how each unit of past demand at time $s$ has moved the current price by time $t$.
+The profit becomes
+$\pi^{i}=\mathbb{E}\!\int_{0}^{T}D_{t}^{i}\underbrace{(V_{T}-P_{t}^{-i})}_{\text{exogenous mispricing}}\,dt\;-\;\mathbb{E}\!\int_{0}^{T}D_{t}^{i}\underbrace{\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds}_{\text{endogenous price impact}}\,dt.$
+The first term is linear in $D^{i}$; the second is a quadratic form in the history of $D^{i}$ through the deterministic kernel $\ell$: $(V_{T}-P_{t}^{-i})$ plays the role of $G^{X,i}$, and the integrated price-impact term plays the role of $D^{i,\top}G^{DD,i}D^{i}$, distributed across time rather than concentrated at each instant.
+Kyle’s lambda and strict concavity.
 Define the aggregate primitive-noise demand kernel $\mathcal{D}_{\Sigma,t}(s):=\sum_{j}\mathcal{D}^{j}(t,s)$. Since $E^{V}E^{0,\top}=0$, all price information comes through drift-based inference:
-$dP_{t}=\lambda(t)\,dI_{t}^{0},\qquad\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du,$ (5.2)
-where $\widetilde{C}^{0}$ is the market maker’s unresolved drift kernel under the projection decomposition (A.12).
-Remark 5.1 (Quadratic structure of the profit).
-Fix opponents’ strategy maps.
-Since the market maker’s filtering gains are deterministic, the price decomposes as $P_{t}=P_{t}^{-i}+\int_{0}^{t}\ell(t,s)\,D_{s}^{i}\,ds$, where $P_{t}^{-i}$ is a linear process independent of $D^{i}$ and $\ell(t,s)$ is a deterministic kernel.
-Symmetrically, the future execution cost of a unit impulse at time $t$ is $\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau$, where $\delta P_{\tau}$ propagates forward through the deviation system with deterministic coefficients.
-The profit $\pi^{i}=\mathbb{E}[\int_{0}^{T}D_{t}^{i}(V_{T}-P_{t})\,dt]$ is therefore concave quadratic in $D^{i}$, when $\lambda(t)>0$ provides strict concavity.
-5.2 Equilibrium characterization
-The optimality condition from a unit demand impulse at $t$ equates the conditional mispricing to the integrated impact cost (Appendix D):
-$\underbrace{\sigma_{V}E^{V}\,\partial_{u}\bigl(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0}\bigr)(u)}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\int_{t}^{T}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau}_{H^{i,\mathrm{imp}}_{t}(u):\;\text{impact cost density}},$ (5.3)
-where $\delta P_{\tau}$ is the price perturbation for an impulse at $t$. Differentiating both sides in $t$ and solving for $D^{i}_{t}(u)$ yields a closed backward ODE:
-$\lambda(t)\,D^{i}_{t}(u)=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (5.4)
-with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
+$dP_{t}=\lambda(t)\,dI_{t}^{0}.$ (5.2)
 Assumption 5.1 (Persistent price informativeness).
 $\lambda(t)>0$ for all $t\in[0,T]$.
-This requires that aggregate order flow remains informative about the fundamental at every date, the economically relevant case since temporary uninformativeness would present an arbitrage opportunity.
-The condition is self-reinforcing: positive $\lambda$ induces finite trading rates, which maintain the drift informativeness that sustains $\lambda>0$.
-It plays the role of $G^{DD,i}\succ 0$ (Lemma 3.4), ensuring strict convexity of the best-response problem on the domain of the fixed-point operator.
-The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix D.
-5.3 Two channels of belief manipulation
-Decompose $\delta P_{\tau}$ into $\delta P_{\tau}^{\mathrm{dir}}$ (computed with $D^{k}\equiv 0$, i.e.
-the market maker’s response alone) and $\delta P_{\tau}^{\mathrm{ind}}:=\delta P_{\tau}-\delta P_{\tau}^{\mathrm{dir}}$ (opponent-mediated).
-Since $\delta P_{\tau}^{\mathrm{dir}}=\lambda(t)$ for all $\tau\geq t$, the impact adjoint splits into:
-  1. (i)
-    $\mathcal{V}^{i,\mathrm{price}}(t,u)=\lambda(t)\int_{t}^{T}D^{i}_{\tau}(u)\,d\tau$: instantaneous impact times the integrated future policy kernel.
-This is the standard Kyle channel—the trader internalizes how each demand impulse moves the price.
-  2. (ii)
-    $\mathcal{V}^{i,\mathrm{opp}}(t,u)=\int_{t}^{T}\delta P_{\tau}^{\mathrm{ind}}\,D^{i}_{\tau}(u)\,d\tau$: opponent-mediated impact, arising because opponents also observe order flow and adjust their trading, which feeds back into prices.
-Back’s model.
-When $n=1$ and $p_{1}=0$ (Back’s model), the opponent channel vanishes ($\delta P_{\tau}^{\mathrm{ind}}\equiv 0$) and (5.3) reduces to Back’s integral equation for the optimal trading rate.
-With $n\geq 2$ and $p_{i}>0$ both channels are active: each trader tempers demand not only to limit direct price impact but to prevent opponents from inferring and front-running through shared order flow.
-Guess-and-verify cannot close this system; the noise-state technique handles it with no additional machinery.
+This is the economically relevant case: if aggregate order flow were temporarily uninformative ($\lambda=0$), traders could trade at zero price impact—an arbitrage opportunity that would restore informativeness.
+The condition plays the role of $G^{DD,i}(t)\succ 0$ in the LQG framework (Lemma 4.4), ensuring strict concavity of the profit in $D^{i}$ and hence necessity and sufficiency of the first-order condition, just through the temporally distributed quadratic form rather than pointwise at each $t$.
+5.2 Best-response characterization
+Fix player $i$ and suppose every opponent $k\neq i$ uses a noise-state linear strategy as in Section 4.
+Because the quadratic structure is cross-temporal, the first-order condition does not directly yield $D^{i}_{t}(u)$; instead, the optimality condition equates the conditional mispricing to the integrated future impact cost (Appendix E):
+$\underbrace{\,\widehat{V}^{i}_{t}-P_{t}}_{H^{i,\mathrm{val}}_{t}(u):\;\text{information rent density}}\;=\;\underbrace{\mathbb{E}\left[\int_{t}^{T}l(\tau,t)\,D^{i}_{\tau}(u)\,d\tau|\mathcal{F}^{i}_{t}\right]}_{H^{i,\mathrm{imp}}_{t}:\;\text{impact cost density}},$ (5.3)
+where $\delta P_{\tau}$ is the price perturbation propagated by the deviation system (E.1)–(E.3).
+The left side is trader $i$’s informational advantage over the market maker; the right side is the present value of price distortion from a marginal unit of demand.
+Since the integral runs over $[t,T]$, it does not isolate $D^{i}_{t}(u)$—only the entire future trading plan weighted by price impact.
+Differentiating both sides in $t$, the Leibniz boundary term produces $\lambda(t)D^{i}_{t}(u)$—where $\lambda(t)$ enters as the endogenous control cost and $D^{i}_{t}(u)$ first appears explicitly.
+Solving:
+$\lambda(t)\,D^{i}_{t}=-\frac{d}{dt}H^{i,\mathrm{val}}_{t}+\mathbb{E}\left[\int_{t}^{T}\frac{\partial l}{\partial t}(\tau,t)\,D^{i}_{\tau}\,d\tau|\mathcal{F}^{i}_{t}\right],$ (5.4)
+with terminal condition $D^{i}_{T}(u)=\lambda(T)^{-1}H^{i,\mathrm{val}}_{T}(u)$.
+The derivative $\frac{d}{dt}H^{i,\mathrm{val}}$ and the deviation system determining $\delta P_{\tau}$ are given in Appendix E.
 6 Conclusion
@@
 The information wedge $\mathcal{V}^{i}_{t}$ prices bilateral belief manipulation and vanishes when signals are exogenous.
-The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs, identifying a second channel through opponent inference via shared order flow that applies to asymmetric precisions and continuous private learning.
+The two-player example shows that nearly all welfare gains from disclosure come from eliminating this channel; the Kyle–Back embedding shows the technique extends beyond quadratic costs to strategic trading with endogenous price impact, asymmetric precisions, and continuous private learning.
 Further directions.
 The main open question is whether the deterministic impulse-response class is without loss of generality.
-A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class and which iterated best responses are expected to contract.
-Beyond this, stochastic coefficients (Remark C.2), graph-structured interactions, and global uniqueness via the dissipative structure of $F^{i}$ are natural extensions.
+A natural program proceeds in three steps: (i) a Blackwell reduction showing that under quadratic costs the best response depends on the posterior $\pi_{t}^{i}:=\mathcal{L}(W_{[0,t]}\mid\mathcal{F}_{t}^{i})$ only through its mean and covariance; (ii) a belief-of-belief collapse in which player $i$’s estimate of $\pi_{t}^{k}$ factors through a deterministic operator—the posterior analogue of the kernel maps $(F^{i},\widetilde{X}^{i})$; (iii) a contraction argument on the variance functional $V(\mathbf{D}):=\sum_{i}\mathbb{E}\int\!\!\int\|D^{i}(t,s,\pi_{t}^{i})-\mathbb{E}[D^{i}(t,s,\pi_{t}^{i})]\|^{2}\,ds\,dt$, which vanishes exactly in the deterministic impulse-response class.
+The precision optimization (Appendix C.2) reveals that multi-player rational inattention couples the filtering and control problems that separate in single-agent settings: the marginal value of precision depends on the equilibrium through the state kernel.
+Extending the framework to rational inattention requires care: varying the gain $\Gamma^{i}$ directly changes the filtration in a non-monotone way, breaking the envelope argument.
+The difficulty disappears when the control variable is the portfolio of independent sources monitored (a Brownian sheet formulation), since each additional source strictly enlarges the filtration.
+The total precision decomposes as $P^{i}=P^{i}_{0}+\tau^{i}$, the marginal value of a source is the value of optimally exploiting its independent increment, and the envelope theorem applies cleanly (Appendix C).
+Overlapping source portfolios endogenize cross-player noise correlations, amplifying the belief-manipulation channel priced by $\mathcal{V}^{i}_{t}$.
+An alternative to the mean-field limit is games on vertex-transitive graphs, where graph symmetry permits a representative-agent reduction while preserving the information wedge.
+The natural conjecture is that equilibrium impulse-response maps inherit the graph symmetry, reducing the fixed point to a single node coupled to its neighborhood—interpolating between the mean-field limit (complete graph, vanishing edge weights, $\bar{\mathcal{V}}^{i}_{t}=0$) and the finite-$n$ case.
+Stochastic coefficients (Remark D.2) and global uniqueness are natural next steps; the dissipative structure of Lemma D.3 suggests sufficient regularity to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here.
 Appendix A Filtering
 Roadmap.
-Throughout, formal calculations identify the kernel equations for $F^{i}$; we then define $F^{i}$ as the unique solution of those equations, with uniqueness following from uniqueness of $L^{2}$-projection.
+We identify the kernel equations for $F^{i}$ via formal calculations, then define $F^{i}$ as their unique solution (uniqueness follows from $L^{2}$-projection).
 The pipeline proceeds in three steps.
-A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations, using an ansatz with undetermined coefficients matched against exponential test martingales.
-A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain via Itô isometry, establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$, and recording the contraction identity $\int_{0}^{t}X_{t}(s)\,\widetilde{X}^{i}_{t}(s)^{\top}\,ds=\Sigma^{XX,i}(t)$.
-A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
+Subsection A.1 derives the dynamics of $t\mapsto\widehat{W}_{t}^{i}(u)$ from a characterizing identity for conditional expectations.
+Subsection A.2 specializes the abstract dynamics to the physical observation model, identifying $\widetilde{X}^{i}_{t}(u)$ as the filtering gain and establishing the algebraic identity $\widetilde{X}^{i}=X(I-\Pi^{i})-X*F^{i}$.
+Subsection A.3 defines $F^{i}$ as the primitive filtering object through its forward evolution equation, derives the explicit three-term formula, and verifies uniqueness.
 A.1 Dynamics of the noise-state
@@
 We now specialize to the observation equation of Section 3,
-$dY_{t}^{i}=\sqrt{P^{i}(t)}\,X_{t}\,dt+E^{i}\,dW_{t},$
-which corresponds to setting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ in the general framework of Proposition A.2.
+$dY_{t}^{i}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},$
+which corresponds to setting $C(t,s)=\Gamma^{i}(t)\,X_{t}(s)$ in the general framework of Proposition A.2.
 Innovation process.
 Define the innovation
-$dI_{t}^{i}:=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt,$
+$dI_{t}^{i}:=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt,$
 where $\widehat{X}_{t}^{i}:=\mathbb{E}[X_{t}\mid\mathcal{F}_{t}^{i}]$.
@@
 Innovation form of the dynamics.
-Substituting $C(t,s)=\sqrt{P^{i}(t)}\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\sqrt{P^{i}(t)}\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
-$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\sqrt{P^{i}(t)}\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
+Substituting $C(t,s)=\Gamma^{i}(t)\,X_{t}(s)$ into Proposition A.2 and using $dI_{t}^{i}=dY_{t}^{i}-\Gamma^{i}(t)\,\widehat{X}_{t}^{i}\,dt$, the dynamics (A.2) become
+$d_{t}\widehat{W}_{t}^{i}(u)=\Big(\int_{0}^{t}\partial_{s}\mathsf{C}_{t}^{i}(u,s)\,X_{t}(s)^{\top}\,ds\Big)\Gamma^{i}(t)\,dI_{t}^{i}\,+\;E^{i,\top}\mathbf{1}_{\{u\geq t\}}\,dI_{t}^{i}.$ (A.9)
 Identifying $\widetilde{X}^{i}$ as the filtering gain.
 For $u<t$, the indicator term in (A.9) is locally constant in $u$, so the mixed $(t,u)$-update takes the form
-$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
-To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$.
-The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\sqrt{P^{i}(t)}(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$.
-Thus $\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
+$d_{t}\!\bigl(d_{u}\widehat{W}_{t}^{i}(u)\bigr)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,dI_{t}^{i}\,du,\qquad 0\leq u<t\leq T.$ (A.10)
+To verify the gain coefficient, recall that $X_{t}-\widehat{X}_{t}^{i}=\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$ and that the innovation (A.17) has martingale component $E^{i}\,dW_{t}$ and absolutely continuous component $\Gamma^{i}(t)\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}\,dt$.
+The gain in (A.10) equals the conditional cross-covariance density $\operatorname{Cov}(dW_{u},\,\Gamma^{i}(t)(X_{t}{-}\widehat{X}_{t}^{i})\,dt\mid\mathcal{F}_{t}^{i})/dt$; by Itô isometry applied to $\int_{0}^{t}\widetilde{X}^{i}_{t}(s)\,dW_{s}$, this picks out $\widetilde{X}^{i}_{t}(u)$.
+Thus $\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)$ is the Kalman gain for updating the $u$-th noise increment: it maps observation innovations at time $t$ into revisions of the noise estimate at the earlier time $u$.
 Remark A.1 (Conditional cross-covariance interpretation).
@@
 In the single-player case with no endogenous drift, this reduces to the classical relation between the Kalman gain and the error covariance.
-In the multi-player setting the identity persists because the forward kernel environment is deterministic.
+In the multi-player setting the identity persists because opponents’ strategy maps have deterministic coefficients.
 A.3 The filtering kernel $F^{i}$
@@
 $\partial_{t}F_{t}^{i}(u,s)$ $=\widetilde{X}^{i}_{t}(u)^{\top}P^{i}(t)\,\widetilde{X}^{i}_{t}(s),\qquad u,s<t,$ (A.14)
-$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u).$ (A.15)
+$F_{t}^{i}(u,t)$ $=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\Gamma^{i}(t)\,\widetilde{X}^{i}_{t}(u).$ (A.15)
 Since (A.12) expresses $\widetilde{X}^{i}$ as a deterministic linear functional of $F^{i}$, the system (A.14)–(A.15) is a closed forward ODE for $F^{i}$ (quadratic through the substitution).
@@
 Integrating (A.14) from $\max(u,s)$ to $t$ and adding the boundary contributions gives
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.16)
 The three terms are: (i) the unresolved response at observation time $s$, projected through the observation channel; (ii) the drift component of the innovation at time $u$, revealing information about earlier increments at $s\leq u$; (iii) accumulated indirect inference from observations at times $r\in[\max(u,s),\,t]$.
@@
 Recall the innovation form of the estimated-noise dynamics (equation (A.9)): for $u\leq t$,
-$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}\Bigg)\,du,$
+$d_{u}\widehat{W}_{t}^{i}(u)=E^{i,\top}\,dI_{u}^{i}+\Bigg(\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}\Bigg)\,du,$
 where the first term is the direct observation at time $u$ and the second is the accumulated drift-based revision over $(u,t]$.
@@
 The physical-measure decomposition of the innovation is
-$dI_{t}^{i}=E^{i}\,dW_{t}\;+\;\sqrt{P^{i}(t)}\,(X_{t}-\widehat{X}_{t}^{i})\,dt.$
+$dI_{t}^{i}=E^{i}\,dW_{t}\;+\;\Gamma^{i}(t)\,(X_{t}-\widehat{X}_{t}^{i})\,dt.$
 The first piece is the martingale component; the second is absolutely continuous.
 Substituting the linear forms and using $\widetilde{X}^{i}_{t}(r):=X_{t}(r)-\widehat{X}^{i}_{t}(r)$ (the definition of $\widetilde{X}^{i}$ and (A.13)):
-$dI_{t}^{i}=E^{i}\,dW_{t}+\sqrt{P^{i}(t)}\int_{0}^{t}\widetilde{X}^{i}_{t}(r)\,dW_{r}\,dt.$ (A.17)
+$dI_{t}^{i}=E^{i}\,dW_{t}+\Gamma^{i}(t)\int_{0}^{t}\widetilde{X}^{i}_{t}(r)\,dW_{r}\,dt.$ (A.17)
 Expanding the direct term.
 Substituting (A.17) at $t=u$ into $E^{i,\top}\,dI_{u}^{i}$:
-$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
+$E^{i,\top}\,dI_{u}^{i}$ $=\underbrace{E^{i,\top}E^{i}\,dW_{u}}_{=\;\Pi^{i}\,dW_{u}}\;+\;E^{i,\top}\Gamma^{i}(u)\int_{0}^{u}\widetilde{X}^{i}_{u}(r)\,dW_{r}\,du.$ (A.18)
 The first piece contributes the $\Pi^{i}\,dW_{u}$ in the density representation.
-The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
+The second is absolutely continuous in $u$ with primitive-noise density $E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)$ at source-noise index $s\leq u$.
 Expanding the integral term.
-Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$:
-$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
+Substituting (A.17) into $\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}$:
+$\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,dI_{s}^{i}$ $=\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\,dW_{s}}_{\text{(a) martingale}}$
 $\quad+\;\underbrace{\int_{u}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\int_{0}^{s}\widetilde{X}^{i}_{s}(r)\,dW_{r}\,ds}_{\text{(b) abs.\ continuous}}.$
-Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
+Term (a) contributes $\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}$ at source-noise index $s\in(u,t]$. Applying stochastic Fubini to term (b) and reading off the coefficient of $dW_{r}$:
 $\text{(b)}=\int_{0}^{t}\Bigg(\int_{\max(u,r)}^{t}\widetilde{X}^{i}_{s}(u)^{\top}P^{i}(s)\,\widetilde{X}^{i}_{s}(r)\,ds\Bigg)\,dW_{r}.$
@@
 Collecting the coefficient of $dW_{s}\,du$ from all three contributions and noting that causality ($\widetilde{X}^{i}_{s}(u)=0$ for $u>s$) automatically zeroes the irrelevant indicators:
-$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
-The three terms admit a direct interpretation:
-  • $\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}$: the unresolved response at observation time $s$, projected through the observation channel.
-Nonzero only when $u\leq s$ (causal in the observation direction).
-  • $E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)$: the drift component of the innovation at time $u$, which reveals information about earlier noise increments at index $s\leq u$.
-Nonzero only when $s\leq u$ (causal in the noise direction).
-  • $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$: accumulated indirect inference—the joint information about the $u$-th and $s$-th increments extracted from observations at times $r\in[\max(u,s),\,t]$.
-Remark A.7 (Self-adjointness and uniqueness).
-Under $(u,s)\mapsto(s,u)$ with transpose, the first two terms exchange and the integral is invariant (since $P^{i}(r)$ is symmetric).
-Hence $F_{t}^{i}(u,s)=F_{t}^{i}(s,u)^{\top}$, confirming that the induced map $(\mathcal{P}_{t}^{i}f)(s):=\Pi^{i}f(s)+\int_{0}^{t}F_{t}^{i}(r,s)^{\top}f(r)\,dr$ is self-adjoint, as required by its interpretation as an orthogonal projection.
-Uniqueness follows from Itô isometry: if $\tilde{F}$ also satisfies the density representation, then $\int_{0}^{t}\|F_{t}^{i}(u,s)-\tilde{F}_{t}^{i}(u,s)\|^{2}\,ds=0$ a.e.
+$F_{t}^{i}(u,s)=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\;+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr.$ (A.19)
+The three terms have the same interpretation as in Proposition A.3.
 Remark A.7 (Boundary condition and evolution).
 Evaluating (A.19) at $s=t$ (using $\widetilde{X}^{i}_{u}(t)=0$ for $u<t$):
-$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\sqrt{P^{i}(t)}\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\sqrt{P^{i}(t)}\,\widetilde{X}^{i}_{t}(u),$
+$F_{t}^{i}(u,t)=\widetilde{X}^{i}_{t}(u)^{\top}\Gamma^{i}(t)\,E^{i},\qquad F_{t}^{i}(t,u)=E^{i,\top}\Gamma^{i}(t)\,\widetilde{X}^{i}_{t}(u),$
 which are each other’s transposes. Differentiating (A.19) in $t$ at fixed $u,s<t$ gives the evolution
@@
 which is manifestly self-adjoint under $(u,s)\mapsto(s,u)^{\top}$.
-Remark A.9 (Computational simplification).
-The formula (A.19) is explicit in the gain kernel $\widetilde{X}^{i}$: once $\widetilde{X}^{i}$ is computed from (A.12), the filtering kernel $F^{i}$ requires no further fixed-point iteration.
-In the numerical scheme of Section 4, the integral $\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr$ is accumulated forward alongside the $\widetilde{X}^{i}$-computation, and the two boundary terms are set at each new observation time.
 A.3.2 Formal characterization
@@
 A deterministic kernel $F^{i}$ is said to be admissible if it is causal ($F_{t}^{i}(u,s)=0$ when $\max\{u,s\}>t$) and, together with the induced kernel
-$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\sqrt{P^{i}(s)}\,E^{i}\;+\;E^{i,\top}\sqrt{P^{i}(u)}\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
+$\begin{split}F_{t}^{i}(u,s)&=\widetilde{X}^{i}_{s}(u)^{\top}\Gamma^{i}(s)\,E^{i}\;+\;E^{i,\top}\Gamma^{i}(u)\,\widetilde{X}^{i}_{u}(s)\\ &\quad+\;\int_{\max(u,s)}^{t}\widetilde{X}^{i}_{r}(u)^{\top}P^{i}(r)\,\widetilde{X}^{i}_{r}(s)\,dr,\qquad u,s\leq t.\end{split}$ (A.20)
 We define $F^{i}$ to be the unique causal solution of (A.12)–(A.20).
@@
 By the innovations exogeneity established in Corollary 4.5, the best-response problem for player $i$ (opponents’ maps fixed) is an optimization over deterministic impulse-response maps weighting the exogenous innovation $I^{i}$.
-[^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [16, Theorem 7.12]; Lemma 3.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 3.5.
+[^2] To avoid circularity: opponents’ deterministic kernels make $I^{i}$ a standard Brownian motion independent of $D^{i}$ [19, Theorem 7.12]; Lemma 4.4 then gives strict convexity over the full $L^{2}$ class; the FOC derived here is therefore necessary and sufficient, yielding the deterministic-kernel best response of Corollary 4.5.
 B.1 The best-response problem in innovations coordinates
@@
 The spike perturbation $D_{s}^{i,\rho,\varepsilon}:=D_{s}^{i}+\rho\,v\,\mathbf{1}_{[t,t+\varepsilon]}(s)$ induces the normalized first variation $\delta X_{s}:=\lim_{\varepsilon\downarrow 0}\frac{1}{\varepsilon}\frac{\partial}{\partial\rho}X_{s}^{\rho,\varepsilon}\big|_{\rho=0}$ with $\delta X_{t}=v$.
-Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients—not a linearization but the full response.
+Because opponents apply fixed linear maps with deterministic kernels, the map $v\mapsto(\delta X_{s},\{\delta w_{s}^{k}(\cdot)\}_{k\neq i})$ is exactly linear in $v$ with deterministic coefficients.
 The deviator’s own innovation is unaffected ($\delta I_{s}^{i}=0$), since $\delta X_{s}\in\mathcal{F}_{t}^{i}$ implies $\delta\widehat{X}_{s}^{i}=\delta X_{s}$.
@@
 with $\Phi^{XX}(t,t)=I$ and $\Phi^{Xk}(t,t,u)=0$.
+Remark B.1 (Fixed coefficients off equilibrium).
+[…]
 B.4 First variation of costs and the adjoint kernels
@@
 As $\theta_{i}\to 0$, $\widehat{W}^{\,i,\theta}_{t}\to\widehat{W}_{t}^{i}$ and (B.18) recovers (B.8).
-The first variation $\delta J^{i}_{\theta_{i},t}(v)$ is computed by differentiating $J^{i}_{\theta_{i}}$ in the spike amplitude (Section B.2), which introduces the exponential tilt $Z^{i}:=e^{\theta_{i}\mathcal{C}^{i}}/\mathbb{E}[e^{\theta_{i}\mathcal{C}^{i}}]$.
-Setting $\delta J^{i}_{\theta_{i},t}(v)=0$ for all $\mathcal{F}_{t}^{i}$-measurable $v$ gives the tilted stationarity condition $G^{DD,i}(t)D^{i}_{t}+\mathbb{E}^{Q^{i}}[M^{i}_{t}\mid\mathcal{F}_{t}^{i}]=0$, where $\mathbb{E}^{Q^{i}}[\cdot\mid\mathcal{F}_{t}^{i}]:=\mathbb{E}[Z^{i}(\cdot)\mid\mathcal{F}_{t}^{i}]/\mathbb{E}[Z^{i}\mid\mathcal{F}_{t}^{i}]$ is the conditional expectation under the cost-tilted measure.
-Since $\mathcal{C}^{i}$ is quadratic in $W$ under the linear profile, $Q^{i}$ is a Gaussian measure, and computing the tilted conditional mean of the stochastic costate $M^{i}_{t}$ yields (B.18).
-B.7 Information cost: proof of Proposition 3.6
-Full- and partial-information controls.
-At the equilibrium profile $D^{*}$, the stochastic costate is
-$M^{i}_{t}:=\bar{H}^{X,i}_{t}+\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}W_{u},$
-where $(\bar{H}^{X,i},H^{X,i})$ solve the backward system (B.11)–(B.14) evaluated at $D^{*}$. The full-information and equilibrium controls are
-$D^{i,\mathrm{full}}_{t}$ $=-(G^{DD,i}(t))^{-1}M^{i}_{t},$ (B.20)
-$D^{i}_{t}$ $=-(G^{DD,i}(t))^{-1}\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}].$ (B.21)
-Both are evaluated at the same adjoint kernels: opponents’ maps are fixed and linear with deterministic kernels, so the backward system is identical regardless of player $i$’s information.
-The control gap is
-$\delta D^{i}_{t}:=D^{i}_{t}-D^{i,\mathrm{full}}_{t}=(G^{DD,i}(t))^{-1}\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u),$
-where $\widetilde{W}^{i}_{t}(u):=W_{u}-\widehat{W}^{i}_{t}(u)$.
-Completion of the square.
-Since $D^{i,\mathrm{full}}$ minimizes the strictly convex quadratic $D^{i}\mapsto J^{i}(D^{i};D^{*,-i})$ over $L^{2}(\mathcal{F})$ (Lemma 3.4), and $\delta D^{i}\in L^{2}(\mathcal{F})$, the first variation vanishes and
-$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}(\delta D^{i}_{t})^{\top}G^{DD,i}(t)\,\delta D^{i}_{t}\,\mathrm{d}t\right].$ (B.22)
-To verify that only $G^{DD,i}$ appears: since $\delta X$ is affine in $\delta D^{i}$ (the dynamics are linear), the full cost perturbation contains terms of order zero, one, and two in $\delta D^{i}$.
-The zero-order term vanishes by definition.
-All first-order terms—including those from $\delta X^{\top}G^{i}_{XX}X^{\mathrm{full}}$ and $(G^{X,i})^{\top}\delta X$—vanish by the stationarity condition at $D^{i,\mathrm{full}}$.
-The second-order terms comprise the direct control cost $(\delta D^{i})^{\top}G^{DD,i}\,\delta D^{i}$ and the state cost $\delta X^{\top}G^{i}_{XX}\,\delta X$ (plus terminal).
-The standard LQG completion-of-the-square identifies these: using $G^{DD,i}(t)\,D^{i,\mathrm{full}}_{t}+M^{i}_{t}=0$ and the definition of $M^{i}_{t}$ through the transition kernel $\Phi_{XX}$ (which incorporates opponents’ linear feedback), the state-cost second-order terms cancel against cross-terms from the adjoint, leaving (B.22).
-Kernel formula for $\Sigma^{M,i}_{t}$.
-Substituting the control gap into (B.22):
-$\Delta J^{i}=\mathbb{E}\!\left[\int_{0}^{T}\left(\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widetilde{W}^{i}_{t}(u)\right)^{\top}(G^{DD,i}(t))^{-1}\left(\int_{0}^{t}H^{X,i}_{t}(v)\,\mathrm{d}\widetilde{W}^{i}_{t}(v)\right)\mathrm{d}t\right].$ (B.23)
-Since $\widetilde{W}^{i}_{t}$ is Gaussian with deterministic covariance $C^{i}_{t}(u,v)$, the expectation evaluates to $\mathrm{tr}[(G^{DD,i}(t))^{-1}\,\Sigma^{M,i}_{t}]$ at each $t$, giving (3.7).
-It remains to verify the kernel representation of $\Sigma^{M,i}_{t}$.
-The conditional mean of $M^{i}_{t}$ given $\mathcal{F}^{i}_{t}$ is
-$\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\bar{H}^{X,i}_{t}+\int_{0}^{t}\widehat{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
-where the resolved adjoint kernel is
-$\widehat{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)\,\Pi^{i}+\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u.$ (B.24)
-This is obtained by substituting the noise-state decomposition $\mathrm{d}_{u}\widehat{W}^{i}_{t}(u)=\Pi^{i}\,\mathrm{d}W_{u}+\bigl(\int_{0}^{t}F^{i}_{t}(u,s)\,\mathrm{d}W_{s}\bigr)\,\mathrm{d}u$ (Theorem 3.1(i)) into $\int_{0}^{t}H^{X,i}_{t}(u)\,\mathrm{d}\widehat{W}^{i}_{t}(u)$ and applying stochastic Fubini.
-Then
-$M^{i}_{t}-\mathbb{E}[M^{i}_{t}\mid\mathcal{F}^{i}_{t}]=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\mathrm{d}W_{s},$
-where the unresolved adjoint kernel $\widetilde{H}^{X,i}_{t}(s):=H^{X,i}_{t}(s)-\widehat{H}^{X,i}_{t}(s)=H^{X,i}_{t}(s)(I-\Pi^{i})-\int_{s}^{t}H^{X,i}_{t}(u)\,F^{i}_{t}(u,s)\,\mathrm{d}u$ is the analogue of $\widetilde{X}^{i}$ for the backward system.
-By Itô isometry,
-$\Sigma^{M,i}_{t}=\int_{0}^{t}\widetilde{H}^{X,i}_{t}(s)\,\widetilde{H}^{X,i}_{t}(s)^{\top}\,\mathrm{d}s.$
-Separation from control.
-The covariance $C^{i}_{t}(u,v)$ of $\widetilde{W}^{i}_{t}$ is deterministic (Gaussianity of the conditional law) and depends on the signal structure $(P^{i},X)$ but not on the realized signal path or on $D^{i}$.
-The adjoint $H^{X,i}$ depends on the equilibrium profile $D^{*,-i}$ through the forward environment, not on $D^{i}$ itself.
-Both $\widetilde{H}^{i}_{X}$ and $\Sigma^{M,i}_{t}$ are therefore deterministic functionals of $(P^{i},D^{*,-i})$, completing the proof.
-∎
-Appendix C Well-posedness of the deterministic impulse-response fixed point
+Appendix C Endogenous Signal Precision
+This appendix identifies the obstruction to precision optimization in the multi-player setting.
+C.1 Why precision variation is hard
+In single-agent LQG, the separation principle makes precision optimization straightforward: the policy gain $S(t)$ does not depend on signal precision, so the value of information can be computed independently of the control problem.
+With multiple agents and endogenous signals, separation fails (Proposition 4.6): the policy kernel $D^{i}_{t}(u)$ depends on precision through the information wedge, so the marginal value of precision depends on the equilibrium.
+A deeper difficulty concerns the observation model itself. The observation equation (3.4),
+$dY^{i}_{t}=\Gamma^{i}(t)\,X_{t}\,dt+E^{i}\,dW_{t},$
+treats each player’s information channel as exogenously given.
+This is adequate when the analyst asks: “holding the information structure fixed, what is the equilibrium?” It does not extend to a player choosing which sources to monitor, because rescaling the gain $\Gamma^{i}\to\Gamma^{i}+\delta\Gamma^{i}$ changes the filtration in a way that is not monotone.
+Concretely, the filtrations generated by $\Gamma^{i}X\,dt+dW^{i}$ and $(\Gamma^{i}+\delta\Gamma^{i})X\,dt+dW^{i}$ are generically incomparable: neither contains the other.
+A player operating at the perturbed precision cannot reconstruct the noise-state path $\widehat{W}^{i}_{t}(\cdot)$ they would have maintained under the original gain—they have different information, not more information.
+This breaks the envelope argument that makes single-agent rational inattention tractable: the player at $P^{i}+\delta P^{i}$ cannot implement the control they would have chosen at $P^{i}$.
+The incomparability disappears when precision is increased by adding an independent source (the Brownian-sheet formulation of Section 3.2).
+Monitoring an additional channel strictly enlarges the filtration, and the marginal value of the source is the value of optimally exploiting its independent increment.
+The full rational-inattention game—in which source portfolios are chosen endogenously—requires this formulation and is left to future work.
+C.2 Two channels from a precision perturbation
+Despite the incomparability, one can compute the first-order effect of a gain perturbation on equilibrium objects by linearizing the forward–backward system.
+The calculation reveals two channels through which $\delta P^{i}$ affects cost.
+Fix an equilibrium profile and spike-perturb player $i$’s gain at time $t_{0}$.
+By the envelope theorem, $\delta D^{i}$ has no first-order cost effect (the FOC is satisfied).
+But the primitive-noise control $\mathcal{D}^{i}_{\tau}(r)=D^{i}_{\tau}(r)\Pi^{i}+\int_{0}^{\tau}D^{i}_{\tau}(u)\,F^{i}_{\tau}(u,r)\,du$ responds at first order through $\delta F^{i}$: the filtering kernel shifts, changing which part of the fixed policy kernel is resolved by the player’s own information.
+The variation of the unresolved adjoint decomposes as
+[…]
+where $\mathcal{G}^{i}$ is the cost propagator and $\widetilde{\delta\mathcal{D}}^{i}$ is the unresolved component of the control variation induced by $\delta F^{i}$.
+Channel (a) resolves more of the fixed adjoint—the standard value-of-information effect present in single-agent problems.
+Channel (b) arises because the same $\delta F^{i}$ shifts the projection boundary, pushing part of $\mathcal{D}^{i}_{\tau}$ outside the player’s current resolution.
+This is the equilibrium channel: it operates through the information wedge, and vanishes precisely when $\mathcal{V}^{i}_{t}=0$ (exogenous signals, Corollary 4.3).
+Multi-player rational inattention therefore couples the filtering and control problems that separate in single-agent settings: the marginal value of precision depends on the equilibrium through the wedge, and optimizing attention requires solving for the equilibrium first.
+Appendix D Well-posedness of the deterministic impulse-response fixed point
 This appendix formulates the equilibrium as a dynamic programming problem on the impulse-response state space, records the a priori bounds that follow from the projection structure of the filtering map, and establishes well-posedness: contraction for short horizons, existence for arbitrary horizons.
@@
 $C$ denotes a generic constant depending on $(n,d,L,\lambda)$; $C(M)$ when it also depends on a policy-norm bound $M$.
-C.1 Standing assumptions
-Assumption C.1.
+D.1 Standing assumptions
+Assumption D.1.
 There exist $L>0$ and $\lambda>0$ such that, for all $t\in[0,T]$ and all $i$, $\|A(t)\|+\|\Sigma(t)\|+\|P^{i}(t)\|+\|G^{XX,i}(t)\|+\|\bar{G}^{X,i}_{t}\|+\|G^{X,i}(t,\cdot)\|_{L^{\infty}}+\|G^{XX,i}(T)\|+\|\bar{G}^{X,i}_{T}\|+\|G^{X,i}(T,\cdot)\|_{L^{\infty}}\leq L$ and $G^{DD,i}(t)\succeq\lambda\,I_{d}$.
-C.2 The kernel state space and its dynamics
+D.2 The kernel state space and its dynamics
 At time $t$ the kernel state is
-$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (C.1)
+$\mathcal{Z}_{t}:=\bigl(\bar{X}_{t},\;X_{t}(\cdot),\;\{F_{t}^{j}\}_{j=1}^{n}\bigr)\;\in\;\mathbb{R}^{d}\times L^{2}([0,t];\mathbb{R}^{d\times(n+1)d})\times\prod_{j=1}^{n}L^{2}_{\mathrm{sym}}([0,t]^{2};\mathbb{R}^{(n+1)d\times(n+1)d}).$ (D.1)
 The unresolved kernel $\widetilde{X}^{j}$ is not a state variable; it is determined algebraically from $(\mathcal{Z}_{t})$ by
-$\widetilde{X}^{j}_{t}(u)=X_{t}(u)(I-\Pi^{j})-\int_{0}^{t}X_{t}(z)\,F_{t}^{j}(z,u)\,dz.$ (C.2)
+$\widetilde{X}^{j}_{t}(u)=X_{t}(u)(I-\Pi^{j})-\int_{0}^{t}X_{t}(z)\,F_{t}^{j}(z,u)\,dz.$ (D.2)
 Under a noise-state linear profile $\mathbf{D}=\{(\bar{D}^{j},D^{j})\}_{j=1}^{n}$, the kernel state evolves by
-$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (C.3)
-$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (C.4)
-$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (C.5)
-with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\sqrt{P^{j}(t)}\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
-The policy $\mathbf{D}$ enters (C.3)–(C.4) but does not enter (C.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (C.2).
-C.3 The Hamilton–Jacobi–Bellman equation
-Let $V^{i}(t,\mathcal{Z})$ denote the equilibrium continuation cost for player $i$ from kernel state $\mathcal{Z}$ at time $t$.
-Since the kernel-state dynamics are deterministic, the HJB is a first-order terminal-value problem:
-$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (C.6)
-where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing, $\ell^{i}$ is the running cost, and $g^{i}$ is the terminal cost, both evaluated at the kernel state.
-Player $i$’s control $(\bar{D}^{i}_{t},D^{i}_{t}(\cdot))$ enters $\dot{\bar{X}}$ and $\dot{X}$ linearly but does not enter $\dot{F}^{j}$; the Hamiltonian minimization over $D^{i}$ therefore involves only $D_{\bar{X}}V^{i}$, $D_{X}V^{i}$, and the quadratic effort cost, giving the first-order condition
-$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u).$ (C.7)
-Along the equilibrium path, $D_{\bar{X}}V^{i}=\bar{H}^{X,i}_{t}$ and $D_{X}V^{i}(\cdot)=H^{X,i}_{t}(\cdot)$, recovering Corollary 3.5.
-The information wedge as a shadow price.
-The $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term in (C.6) captures the indirect effect of the kernel state’s evolution on player $i$’s continuation cost through opponents’ filters.
-Along the equilibrium path, differentiating through the chain $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies
-$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle,$ (C.8)
-i.e.
-the mean information wedge of Theorem 3.2 is the shadow price of the opponent’s filter evolution, contracted against the sensitivity of that evolution to the state kernel.
+$\dot{\bar{X}}(t)$ $=A(t)\bar{X}_{t}+\textstyle\sum_{k}\bar{D}^{k}_{t},$ (D.3)
+$\dot{X}_{t}(s)$ $=A(t)X_{t}(s)+\textstyle\sum_{k}\mathcal{D}^{k}_{t}(s),\qquad X(s,s)=\Sigma(s),$ (D.4)
+$\partial_{t}F_{t}^{j}(u,s)$ $=\widetilde{X}^{j}_{t}(u)^{\top}P^{j}(t)\,\widetilde{X}^{j}_{t}(s),\qquad F_{0}^{j}=0,$ (D.5)
+with boundary $F_{t}^{j}(u,t)=\widetilde{X}^{j}_{t}(u)^{\top}\Gamma^{j}(t)\,E^{j}$, and primitive-noise control $\mathcal{D}^{k}_{t}(s):=D^{k}_{t}(s)\Pi^{k}+\int_{0}^{t}D^{k}_{t}(u)\,F_{t}^{k}(u,s)\,du$.
+The policy $\mathbf{D}$ enters (D.3)–(D.4) but does not enter (D.5) directly: $F^{j}$ depends on $\mathbf{D}$ only through $X$ via (D.2).
+D.3 The Hamilton–Jacobi–Bellman equation
+Since the kernel-state dynamics (D.3)–(D.5) are deterministic, the equilibrium continuation cost $V^{i}(t,\mathcal{Z})$ satisfies a first-order terminal-value PDE (no diffusion term):
+$\partial_{t}V^{i}+D_{\bar{X}}V^{i}\cdot\dot{\bar{X}}+\langle D_{X}V^{i},\,\dot{X}\rangle+\sum_{j=1}^{n}\langle D_{F^{j}}V^{i},\,\dot{F}^{j}\rangle+\ell^{i}(t,\mathcal{Z},\mathbf{D})=0,\qquad V^{i}(T,\cdot)=g^{i}(\cdot),$ (D.6)
+where $\langle\cdot,\cdot\rangle$ denotes the $L^{2}$ pairing. Player $i$’s control enters $\dot{\bar{X}}$ and $\dot{X}$ but not $\dot{F}^{j}$, so the Hamiltonian minimization yields
+$\bar{D}^{i}_{t}=-(G^{DD,i}(t))^{-1}D_{\bar{X}}V^{i},\qquad D^{i}_{t}(u)=-(G^{DD,i}(t))^{-1}D_{X}V^{i}(u),$ (D.7)
+recovering Corollary 4.5 along the equilibrium path.
+The remaining $\sum_{j}\langle D_{F^{j}}V^{i},\dot{F}^{j}\rangle$ term captures the indirect cost of opponents’ filter evolution; differentiating through $D^{i}\to X\to\widetilde{X}^{k}\to F^{k}$ identifies the mean information wedge as a shadow price:
+$\bar{\mathcal{V}}^{i}_{t}=\sum_{k\neq i}\int_{0}^{t}P^{k}(t)\,\widetilde{X}^{k}_{t}(z)\,\bar{H}^{k,i}_{t}(z)\,dz=\sum_{k\neq i}\langle D_{F^{k}}V^{i},\,\partial_{X}\dot{F}^{k}\rangle.$ (D.8)
 The kernel wedge $\mathcal{V}^{i}_{t}(r)$ admits the same interpretation with $\bar{H}^{k,i}$ replaced by $H^{k,i}(\cdot,\cdot,r)$.
-Endogenous precision.
-If $P^{i}(\cdot)$ is player $i$’s control subject to attention cost $\int_{0}^{T}c^{i}(t,P^{i}(t))\,dt$, then $P^{i}$ enters (C.6) only through $\dot{F}^{i}$ via (C.5), and the FOC
-$\bigl\langle D_{F^{i}}V^{i},\;\widetilde{X}^{i}_{t}(\cdot)^{\top}(\cdot)\,\widetilde{X}^{i}_{t}(\cdot)\bigr\rangle=\frac{\partial c^{i}}{\partial P^{i}}(t,P^{i}(t))$ (C.9)
-decouples from the FOC for $D^{i}$: the left side is the marginal value of accelerating the filter, which vanishes as $\mathcal{P}_{t}^{i}\to I$ (Lemma C.3), guaranteeing an interior solution when $\partial c^{i}/\partial P^{i}>0$.
-C.4 A priori bounds from projection structure
+D.4 A priori bounds from projection structure
 The filtering kernel $F_{t}^{j}$ parametrizes a portion of the conditional projection $(\mathcal{P}_{t}^{j}f)(u):=\Pi^{j}f(u)+\int_{0}^{t}F_{t}^{j}(u,s)\,f(s)\,ds$, an orthogonal projection on $L^{2}([0,t];\mathbb{R}^{(n+1)d})$ (Remark A.4).
-Lemma C.2 (Projection bounds).
+Lemma D.2 (Projection bounds).
 For each $j$ and $t\geq 0$: (i) $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$; (ii) both $\mathcal{P}_{t}^{j}$ and $I-\mathcal{P}_{t}^{j}$ are $L^{2}$-contractions; (iii) $\|\mathcal{F}_{t}^{j}\|_{\mathrm{op}}\leq 2$ where $\mathcal{F}_{t}^{j}:=\mathcal{P}_{t}^{j}-\Pi^{j}$.
@@
 Standard properties of orthogonal projections; (iii) is the triangle inequality. ∎
-Lemma C.3 (Dissipation).
+Lemma D.3 (Dissipation).
 $\{\mathcal{P}_{t}^{j}\}_{t\geq 0}$ is non-decreasing in the positive semidefinite order, and $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}([0,t])}\leq\|X(t,\cdot)\|_{L^{2}([0,t])}$ for all $t$.
 Proof.
-Equation (C.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing.
-Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (C.2), Lemma C.2(ii) gives the bound.
+Equation (D.5) adds a positive semidefinite increment to $F_{t}^{j}$, so $\mathcal{P}_{t}^{j}$ is non-decreasing.
+Since $\widetilde{X}^{j}(t,\cdot)=(I-\mathcal{P}_{t}^{j})X(t,\cdot)$ by (D.2), Lemma D.2(ii) gives the bound.
 ∎
 These bounds hold uniformly in the precision paths, the policy profile, and the horizon $T$.
-The quadratic right-hand side of (C.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$.
-This prevents finite-time blowup despite the quadratic nonlinearity and is the primitive noise-linear manifestation of Kalman variance collapse.
-C.5 Well-posedness
-Definition C.1.
+The quadratic right-hand side of (D.5) is self-limiting: as $\mathcal{P}_{t}^{j}$ grows toward $I$, the complementary projection contracts, driving $\widetilde{X}^{j}\to 0$ and decelerating $F^{j}$.
+This prevents finite-time blowup despite the quadratic nonlinearity—the noise-state analogue of Kalman variance collapse.
+D.5 Well-posedness
+Definition D.1.
 $\mathbb{B}_{T}^{M}:=\bigl\{\mathbf{D}=\{(\bar{D}^{i},D^{i})\}_{i=1}^{n}:\;\bar{D}^{i}\in C([0,T];\mathbb{R}^{d}),\;D^{i}\in L^{\infty}(\Delta_{T}),\;\|\mathbf{D}\|\leq M\bigr\}$, where $\|\mathbf{D}\|:=\max_{i}(\|\bar{D}^{i}\|_{C}+\|D^{i}\|_{L^{\infty}})$.
-Definition C.2 (Best-response operator).
+Definition D.2 (Best-response operator).
 $\mathcal{T}(\mathbf{D}):=\bigl\{(-(G^{DD,i})^{-1}\bar{H}^{X,i,\mathbf{D}},\;-(G^{DD,i})^{-1}H^{X,i,\mathbf{D}})\bigr\}_{i=1}^{n}$, where the adjoints solve the backward system of Theorem 4.2 at the forward environment induced by $\mathbf{D}$.
 Fixed points are Nash equilibria (Corollary 4.5, Remark 4.3).
-The following estimates are proved by Grönwall arguments on the forward system (C.3)–(C.5) (using the a priori bounds of Lemmas C.2C.3 to control the quadratic nonlinearity in (C.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
-Proposition C.4 (Bounds and Lipschitz continuity).
-Under Assumption C.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
+The following estimates are proved by Grönwall arguments on the forward system (D.3)–(D.5) (using the a priori bounds of Lemmas D.2D.3 to control the quadratic nonlinearity in (D.5)) and by backward Grönwall on the linear adjoint system (B.11)–(B.14) (using Cauchy–Schwarz on the belief-adjoint coupling integrals $\int_{0}^{t}P^{k}\widetilde{X}^{k}\,\bar{H}^{k,i}\,dz$).
+Proposition D.4 (Bounds and Lipschitz continuity).
+Under Assumption D.1, for $\mathbf{D}\in\mathbb{B}_{T}^{M}$:
 (i) The forward system has a unique global solution with $\|\bar{X}\|_{C}+\|X\|_{L^{\infty}}\leq K_{\mathrm{fwd}}$, $\|\widetilde{X}^{j}(t,\cdot)\|_{L^{2}}\leq K_{\mathrm{fwd}}\sqrt{T}$, $\|\mathcal{P}_{t}^{j}\|_{\mathrm{op}}=1$.
@@
 For the forward system: $(\bar{X},X)$ solve linear equations with bounded coefficients once $\mathcal{D}^{k}$ is controlled.
-Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma C.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
-For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (C.2).
+Since $\mathcal{D}^{k}$ involves $F^{k}$ through a linear integral, and $\|\mathcal{F}_{t}^{k}\|_{\mathrm{op}}\leq 2$ (Lemma D.2), the coupled system for $\phi(t):=\|X\|_{L^{\infty}(\Delta_{t})}$ closes as $\phi(t)\leq C(M)+C(M)\int_{0}^{t}\phi(r)\,dr$.
+For the Lipschitz bound: the variation $\delta\mathcal{D}^{k}$ decomposes into a direct term bounded by $C(M)\|\delta\mathbf{D}\|$ and an indirect term through $\delta F^{k}$, which couples to $\delta\widetilde{X}^{k}$ and back to $\delta X$ via (D.2).
 A Grönwall argument on the aggregate $\Psi(t):=\|\delta X\|_{L^{\infty}(\Delta_{t})}+\max_{j}\sup_{r\leq t}\|\delta\widetilde{X}^{j}(r,\cdot)\|_{L^{2}}$ gives $\Psi(T)\leq C(M)Te^{C(M)T}\|\delta\mathbf{D}\|$; the factor of $T$ arises because the policy enters through integrals of length $\leq T$.
 The backward estimates follow from linearity of the adjoint system, Cauchy–Schwarz on coupling terms, and backward Grönwall. ∎
-Theorem C.5 (Short-horizon equilibrium).
-Under Assumption C.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$.
+Theorem D.5 (Short-horizon equilibrium).
+Under Assumption D.1, there exists $T^{*}>0$ depending only on $(n,d,L,\lambda)$ such that for $\in(0,T^{*}], 𝒯\mathcal{T}$ is a contraction on $\mathbb{B}_{T}^{M_{0}}$ for a suitable $M_{0}$.
 In particular, there exists a unique Nash equilibrium in noise-state linear strategies with bounded kernels, Picard iterates converge geometrically, and the equilibrium is Nash over the full admissible $L^{2}$ class.
 Proof.
-By Proposition C.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$.
+By Proposition D.4(iii), $\mathrm{Lip}(\mathcal{T})\leq\lambda^{-1}C(M_{0})Te^{C(M_{0})T}$.
 Self-mapping ($\mathcal{T}:\mathbb{B}_{T}^{M_{0}}\to\mathbb{B}_{T}^{M_{0}}$) holds because at $M=0$ the belief-adjoint decouples ($D^{k}\equiv 0$ kills the forcing in (B.12)), giving finite $K_{\mathrm{bwd}}(0,T)$; the intermediate value theorem yields $M_{0}$.
@@
 ∎
-Theorem C.6 (Global existence).
-Under Assumption C.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
+Theorem D.6 (Global existence).
+Under Assumption D.1, for every $T>0$ there exists a Nash equilibrium over the full admissible $L^{2}$ class.
 Proof sketch.
@@
 $\tau$-continuity of $\mathcal{T}$: on the bounded set $\mathbb{B}_{T}^{M_{0}}$, $L^{2}$-convergence of policy kernels propagates through the forward linear equations (via Cauchy–Schwarz) and the linear backward system.
-$\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition C.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([1, Theorem 1.3.2]).
+$\tau$-precompactness of the image: the uniform time-derivative bound on $\bar{H}^{X,i}$ (Proposition D.4(ii)) gives equicontinuity of $\bar{D}^{i,\mathrm{new}}$ in $C([0,T])$; the same bound gives pointwise-in-$u$ equicontinuity of $D^{i,\mathrm{new}}(t,u)$ in $t$, which with the uniform $L^{\infty}$ bound yields $L^{2}$-equicontinuity of $t\mapsto D^{i,\mathrm{new}}(t,\cdot)$ and hence sequential compactness in $L^{2}(\Delta_{T})$ by the vector-valued Arzelà–Ascoli theorem ([2, Theorem 1.3.2]).
 ∎
-Remark C.1 (Uniqueness beyond the contraction regime).
-Theorem C.6 gives existence but not uniqueness for $T>T^{*}$.
-The dissipation of Lemma C.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here.
+Remark D.1 (Uniqueness beyond the contraction regime).
+Theorem D.6 gives existence but not uniqueness for $T>T^{*}$.
+The dissipation of Lemma D.3 suggests enough structure to preclude multiplicity, but a proof requires monotonicity estimates beyond the Grönwall bounds used here.
 In Section 2, Picard iteration converges to the same limit from multiple initial conditions for all parameter values tested.
-Remark C.2 (Stochastic coefficients).
-If $A(t,\omega)$, $P^{i}(t,\omega)$, or the cost matrices depend on a public factor process $\xi_{t}$ (e.g.
-a finite-state Markov chain representing regime switches or stochastic automation events), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$.
-The projection bounds of Lemmas C.2C.3 are algebraic properties of orthogonal projections and hold pathwise, so the forward well-posedness extends to each sample path between jumps of $\xi$.
+Remark D.2 (Stochastic coefficients).
+If primitives depend on a public factor process $\xi_{t}$ (e.g.
+a finite-state Markov chain representing regime switches), the kernel state augments to $(\mathcal{Z}_{t},\xi_{t})$.
+The projection bounds of Lemmas D.2D.3 are algebraic and hold pathwise, so forward well-posedness extends between jumps of $\xi$.
 Each jump triggers a regime transition in which filtering gains, strategies, and the information wedge shift discontinuously, while $\mathcal{Z}$ is continuous across jumps.
-The backward ODE is replaced by a backward stochastic differential equation in the adjoint kernels, driven by jumps of $\xi$; the deterministic Grönwall arguments extend to pathwise bounds conditional on the regime path.
-Appendix D Kyle–Back derivations
-This appendix derives the deviation propagation system, the first-order condition (5.3), and the closed backward ODE (5.4) stated in Section 5.
-D.1 Drift kernels
-The market maker’s drift kernel is $C^{0}(t,s):=\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)$; for trader $k$:
-$C^{k}(t,s):=\begin{pmatrix}\sigma_{0}^{-1}\mathcal{D}_{\Sigma,t}(s)\\[2.0pt] \sqrt{p_{k}}\,\sigma_{V}\,E^{V}\end{pmatrix}\!,\qquad\Pi_{\mathcal{E}}^{k}:=\Pi^{0}+\Pi^{k}.$ (D.1)
-The unresolved kernels $\widetilde{C}^{0}$ and $\widetilde{C}^{k}$ are defined by the projection decomposition (A.12) with $C^{j}$ replacing $\sqrt{P^{j}}X$ and $\Pi_{\mathcal{E}}^{j}$ replacing $\Pi^{j}$.
-D.2 Deviation propagation
-A spike perturbation of trader $i$’s demand at time $t$ enters aggregate order flow, triggering a coupled forward response.
-For $s>t$, let $\delta D_{\Sigma,s}$ be the aggregate demand perturbation and $\delta w_{s}^{j}(u)$ the noise-state density perturbation for observer $j$.
-The deviation system (cf.
-Section B.2) is:
-$\delta D_{\Sigma,s}$ $=\sum_{k\neq i}\int_{0}^{s}D^{k}_{s}(u)\,\delta w_{s}^{k}(u)\,du,$ (D.2)
-$\partial_{s}\delta w_{s}^{0}(u)$ $=\widetilde{C}^{0}(s,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,s}-\int_{0}^{s}C^{0}(s,r)\,\delta w_{s}^{0}(r)\,dr\biggr),$ (D.3)
-$\partial_{s}\delta w_{s}^{k}(u)$ $=\widetilde{C}^{k}(s,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,s}\\ 0\end{pmatrix}-\int_{0}^{s}C^{k}(s,r)\,\delta w_{s}^{k}(r)\,dr\biggr),\quad k\neq i,$ (D.4)
-$\delta P_{s}$ $=\sigma_{V}\int_{0}^{s}E^{V}\delta w_{s}^{0}(u)\,du,$ (D.5)
-with $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, $\delta I_{s}^{i}=0$.
-The demand perturbation enters each observer’s stacked observation through the order-flow row only.
-Equation (D.2) closes the loop: opponents’ noise-state shifts feed through their policy kernels back into aggregate demand.
-Since $E^{V}E^{0,\top}=0$, $\delta P_{t}=\lambda(t)$; when $n=1$, $\delta D_{\Sigma,s}=0$ for $s>t$ and $\delta P_{s}=\lambda(t)$ for all $s\geq t$.
-D.3 First-order condition
-The optimality condition from a unit demand impulse at $t$ is
-$\mathbb{E}\!\left[V_{T}-P_{t}-\int_{t}^{T}D_{\tau}^{i}\,\delta P_{\tau}\,d\tau\;\Big|\;\mathcal{F}_{t}^{i}\right]=0.$ (D.6)
-The mispricing is $\mathbb{E}[V_{T}-P_{t}\mid\mathcal{F}_{t}^{i}]=\sigma_{V}\int_{0}^{t}E^{V}\,d_{u}(\widehat{W}_{t}^{i}-\widehat{W}_{t}^{0})(u)$ (since $\mathcal{F}_{t}^{0}\subset\mathcal{F}_{t}^{i}$).
-The conditional future demand in noise-state linear form is $\mathbb{E}[D_{\tau}^{i}\mid\mathcal{F}_{t}^{i}]=\int_{0}^{t}D^{i}_{\tau}(u)\,d_{u}\widehat{W}_{t}^{i}(u)$ for $\tau>t$ (the noise-state is a martingale, so only $u\leq t$ contributes).
-Both terms are linear in $\widehat{W}_{t}^{i}$, so matching noise-state densities gives (5.3).
-$D^{i}_{t}(u)$ does not appear directly: the integral over $[t,T]$ does not isolate $\tau=t$.
-To pin down $D^{i}_{t}(u)$, we differentiate both sides in $t$.
-D.4 Leibniz differentiation and backward ODE
-Impact adjoint.
-Leibniz on $H^{i,\mathrm{imp}}$ gives
-$\frac{d}{dt}H^{i,\mathrm{imp}}_{t}(u)=-\lambda(t)\,D^{i}_{t}(u)+\int_{t}^{T}\partial_{t}\delta P_{\tau}\,D^{i}_{\tau}(u)\,d\tau,$ (D.7)
-where the boundary term $\delta P_{t}\cdot D^{i}_{t}(u)=-\lambda(t)D^{i}_{t}(u)$ is instantaneous self-impact—this is where $\lambda(t)$ enters as the endogenous $G^{DD,i}$ and where $D^{i}_{t}(u)$ first appears explicitly.
-Value adjoint.
-In primitive-noise coordinates (Theorem 3.1(i)), $H^{i,\mathrm{val}}_{t}(u)$ has a $t$-independent term (the direct observation gap $\Pi_{\mathcal{E}}^{i}-\Pi^{0}$) and an integral $\int_{0}^{t}(F_{t}^{i}-F_{t}^{0})(v,u)\,dv$.
-Differentiating by Leibniz and the kernel evolution $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$:
-$\frac{d}{dt}H^{i,\mathrm{val}}_{t}(u)=\sigma_{V}E^{V}\biggl[\bigl(F_{t}^{i}-F_{t}^{0}\bigr)(t,u)+\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,v)^{\top}\widetilde{C}^{i}(t,u)-\widetilde{C}^{0}(t,v)^{\top}\widetilde{C}^{0}(t,u)\bigr)\,dv\biggr].$ (D.8)
-Closed backward ODE.
-Equating $\frac{d}{dt}H^{i,\mathrm{val}}=\frac{d}{dt}H^{i,\mathrm{imp}}$ and solving for $D^{i}_{t}(u)$ yields (5.4).
-The derivative $\partial_{t}\delta P_{\tau}$ is obtained by differentiating (D.2)–(D.5) in the impulse time, following the same logic as Section B.2 with $\sqrt{P^{k}}X\to C^{k}$, $\widetilde{X}^{k}\to\widetilde{C}^{k}$.
-Convexity (from $\lambda(t)>0$) ensures sufficiency.
+The backward ODE becomes a BSDE driven by jumps of $\xi$; the Grönwall arguments extend to pathwise bounds conditional on the regime path.
+Appendix E Kyle–Back derivations
+This appendix derives the price-impact kernel $l(\tau,t)$, the first-order condition (5.3), and the backward ODE (5.4), with all opponents’ noise-state linear strategy maps held fixed.
+E.1 The price-impact kernel and Kyle’s lambda
+A unit spike perturbation of trader $i$’s demand at time $t$ triggers a coupled forward response: the market maker revises the price, each opponent revises their estimate, adjusts their trading, and the adjustments feed back into prices.
+Let $\delta D_{\Sigma,\tau}$ denote the aggregate demand perturbation and $\delta w_{\tau}^{j}(u)$ the noise-state density perturbation for observer $j$.
+The deviation system is:
+$\delta D_{\Sigma,\tau}$ $=\sum_{k\neq i}\int_{0}^{\tau}D^{k}_{\tau}(u)\,\delta w_{\tau}^{k}(u)\,du,$ (E.1)
+$\partial_{\tau}\delta w_{\tau}^{0}(u)$ $=\widetilde{C}^{0}(\tau,u)^{\top}\biggl(\sigma_{0}^{-1}\delta D_{\Sigma,\tau}-\int_{0}^{\tau}C^{0}(\tau,r)\,\delta w_{\tau}^{0}(r)\,dr\biggr),$ (E.2)
+$\partial_{\tau}\delta w_{\tau}^{k}(u)$ $=\widetilde{C}^{k}(\tau,u)^{\top}\biggl(\begin{pmatrix}\sigma_{0}^{-1}\delta D_{\Sigma,\tau}\\ 0\end{pmatrix}-\int_{0}^{\tau}C^{k}(\tau,r)\,\delta w_{\tau}^{k}(r)\,dr\biggr),\quad k\neq i,$ (E.3)
+with initial conditions $\delta w_{t}^{j}\equiv 0$, $\delta D_{\Sigma,t}=1$, and $\delta I_{\tau}^{i}=0$ (the deviator’s own innovation is unaffected).
+Equation (E.1) closes the loop: each opponent’s noise-state shift propagates through their policy kernel back into aggregate demand, which in turn shifts the market maker’s and other opponents’ noise-states.
+Price-impact kernel.
+Define the price-impact kernel as the price response to a unit demand impulse at time $t$:
+$l(\tau,t):=\sigma_{V}\int_{0}^{\tau}E^{V}\delta w_{\tau}^{0}(u)\,du,\qquad\tau\geq t.$ (E.4)
+This is the object $\ell(t,s)$ of the price decomposition (5.1): the market maker’s filtering gains are deterministic, so $l(\tau,t)$ is deterministic.
+Kyle’s lambda.
+Since $E^{V}E^{0,\top}=0$, the market maker learns about the fundamental only through drift-based inference, and the instantaneous price impact is
+$l(t,t)=\lambda(t):=\sigma_{V}\!\int_{0}^{t}E^{V}\widetilde{C}^{0}(t,u)^{\top}du.$ (E.5)
+E.2 Leibniz differentiation and the backward ODE
+Differentiating (5.3) in $t$, Leibniz on the impact side produces $\lambda(t)\,D^{i}_{t}$ plus $\mathbb{E}[\int_{t}^{T}\frac{\partial l}{\partial t}(\tau,t)\,D^{i}_{\tau}\,d\tau\mid\mathcal{F}_{t}^{i}]$.
+The value side requires more care.
+Value side.
+The conditional mispricing $H^{i,\mathrm{val}}_{t}=\widehat{V}_{t}^{i}-P_{t}=\sigma_{V}E^{V}\bigl(\widehat{W}^{i}_{t}(t)-\widehat{W}^{0}_{t}(t)\bigr)$ can be written as
+$\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\bigl(F^{i}_{t}(u,s)-F^{0}_{t}(u,s)\bigr)\,du\biggr]dW_{s}\,.$
+The integrand kernel is deterministic and evolves through $\partial_{t}F_{t}^{j}(v,u)=\widetilde{C}^{j}(t,v)^{\top}\widetilde{C}^{j}(t,u)$. Differentiating by Leibniz:
+$dH^{i,\mathrm{val}}_{t}=\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\bigl(\widetilde{C}^{i}(t,u)^{\top}\widetilde{C}^{i}(t,s)-\widetilde{C}^{0}(t,u)^{\top}\widetilde{C}^{0}(t,s)\bigr)\,du\biggr]dW_{s}\,dt\;+\;dM_{t}\,$, (E.6)
+where $dM_{t}$ collects the martingale increment from the boundary at $s=t$.
+The remaining boundary term (from the upper limit of the inner integral) vanishes because $E^{V}E^{i,\top}=E^{V}E^{0,\top}=0$: neither player’s observation noise correlates with the fundamental.
+To condition on $\mathcal{F}_{t}^{i}$, note that $\widetilde{C}^{i}(t,\cdot)$ is by definition the component of the drift kernel that player $i$ has not resolved—it is orthogonal to $\mathcal{F}_{t}^{i}$.
+The $\widetilde{C}^{i}$ cross terms therefore vanish under conditioning, and $dW_{s}$ projects to $d\widehat{W}^{i}_{t}(s)$, giving
+$\mathbb{E}\!\left[\frac{dH^{i,\mathrm{val}}_{t}}{dt}\;\Big|\;\mathcal{F}_{t}^{i}\right]=-\sigma_{V}E^{V}\int_{0}^{t}\biggl[\int_{0}^{t}\widetilde{C}^{0}(t,u)^{\top}\widetilde{C}^{0}(t,s)\,du\biggr]d\widehat{W}^{i}_{t}(s)\,.$